### v5

# Are Biological Systems More Intelligent Than Artificial Intelligence?

Abstract
Are biological self-organising systems more ‘intelligent’ than artificial intelligence (AI)? If so, why? I explore this through a mathematical lens which frames intelligence in terms of adaptability. I model systems as stacks of abstraction layers (Stack Theory) and compare them by how they delegate agentic control down their stacks, illustrating with examples of computational, biological, human military, governmental and economic systems. Contemporary AI rests on a static, human-engineered stack in which lower layers are static during deployment. Put provocatively, static stacks resemble inflexible bureaucracies, adapting only top-down. Biological stacks are more ‘intelligent’ because they delegate adaptation. Formally, I prove a theorem (The Law of the Stack) showing adaptability in higher layers requires sufficient adaptability in lower layers. Generalising bio-electric explanations of cancer as isolation from collective informational structures, I explore how cancer-like failures occur in non-biological systems when delegation is inadequate. This helps explain how to build more robust systems, by delegating control like the military doctrine of mission command. It also provides a design perspective on hybrid agents (e.g. organoids, systems involving both humans and AI): hybrid creation is a boundary-condition design problem in which human-imposed constraints prune low-level policy spaces to yield desired collective behaviour while preserving collective identity.
1 Introduction
Biological systems self-organise; their parts and the environment interact to produce spatiotemporal order [ashby1947, seeley2002, rosas2018, bell2007, kelso1997, Friston2010, tognoli2014]. They pursue homeostatic and reproductive goals. Some proactively learn [kukushkin2024], storing information about the past in order to anticipate future needs. To make accurate predictions in complex environments, a system must learn cause and effect [pearl2018, merker2005, barron2016]. I consider a system’s ‘intelligence’ to be a measure of how efficiently it does this [bennett2024a, chollet2019, bennettmaruyama2022b, bennett2023c] [^1] Throughout, “intelligence” is used in an efficiency-of-adaptation sense (sample and energy efficiency). For a comparison with other common definitions, see Electronic Supplementary Materials (ESM) Note S1.. In this view intelligence is a long-term adaptation hard-wired into the organism, that enables short-term adaptation during its lifetime [bennett2025thesis].
Artificial intelligence (AI) is the study of adaptive machines, and machine learning (ML) is the study of machines that learn [russell2005ai, bishop2006, sutton2018, goertzel2014]. Contemporary approaches to machine learning are often inefficient [bennettmaruyama2022a]. They can require vast quantities of data and energy to learn tasks a human can pick up after one or two examples. Even organoids [^2] By organoid I mean a collective of cells grown in a lab, placed in a computational interface and life support system, and then ‘trained’ to communicate with conventional computing technology. appear more ‘intelligent’ in this sense [khajehnejad2024]. For example, an organoid outperformed conventional reinforcement learning systems at the game Pong, adapting with greater sample efficiency. That organoid was a highly tuned hybrid system whose behaviour was shaped by human-imposed boundary conditions. This is similar to how an acellular slime mold can be made to ‘solve’ shortest-path problems [nakagaki2000]. These sorts of results are highly dependent upon the human imposed boundary conditions, but they raise an interesting question. Are biological self-organising systems more intelligent than artificial intelligence? If so, why?
Some say the key difference between biological systems and contemporary artificial intelligence is that the former are active inference systems, while the latter are typically not [parr2024, Friston2010, friston2013]. It is true that active inference can improve efficiency [ororbia2022, ororbia2023, ororbia2023b]. Active inference has been applied in the context of the ARC-AGI benchmark [chollet2019], which to some extent tests sample efficiency. Active inference systems performed well (for example the 2024 ARC-AGI submission by Jack Cole), but still underperform humans. Hence there may be more to the picture [^3] From the perspective of this paper, Active Inference/FEP is primarily a theory of how an agent updates beliefs and selects actions (e.g. via variational free energy minimisation), whereas this paper is more about where and at what level of abstraction adaptation is realised in a system’s stack. Hence I treat the approaches as potentially complementary. An extended comparison and related work is provided in ESM Note S6..
Others seem to suggest biological systems are not more intelligent [^4] In the sense of sample and energy efficiency of adaptation., and the difference is primarily one of inductive bias [chollet2019]. This inductive bias is hard-coded foreknowledge of the environment, which makes it easier to learn and adapt to that environment. To whatever extent this is true, it substitutes the question ‘what is intelligence?’ for ‘what is the inductive bias?’.
To reiterate, I want to know why biological self-organisation seems more adaptive [^5] Sample and energy efficient. than contemporary AI. Is there something else biological systems do that allows them to be more ‘intelligent’? Could there be more to the picture than active inference and an inductive bias?
I suggest there is. Computers are organised into layers of abstraction, each layer being the behaviour of the layer below. For example, software runs on hardware. If software is making decisions, then it controls the hardware top-down. In dominant contemporary AI deployments, adaptation is concentrated at a comparatively high level of abstraction (e.g. model parameters and software-level state), while the lower computational stack (instruction set, microarchitecture, hardware) is treated as fixed during deployment. [^6] This is not a claim about a fundamental limitation of computational systems: reconfigurable or self-modifying substrates (e.g. FPGAs, adaptive firmware, agentic materials, or bio-hybrid substrates) can in principle shift adaptation downward. The claim here is about the dominant architectures and workflows in contemporary AI. In contrast, biological self-organisation uses the multiscale competency architecture (MCA). Adaptation takes place at every scale, from cells to organs to organisms [fields2020]. It is ‘scale-free’. The same dynamics are observed at different scales. Control is delegated, in the sense that individual cells pursue goals [^7] As opposed to in a computer where the parts are usually designed to passively await instruction.. Cells are a “soft” material that has agency [man2019, ball2023]. A cell has a perceptual field. A cell can co-operate with other cells in this perceptual field. This allows them to form multicellular networks that have a larger perceptual field than any individual cell, to maintain coherent higher level “identity” like that of an organism [levin2024]. Each cell has its own goals and is intelligent in the sense of being able to reach those goals by various means [james1890, ciaunica2023]. However each cell can also be part of a collective of cells that in turn pursues its own collective goals. Smaller collectives can form larger collectives, and the pattern repeats at larger scales. Collectives are nested [ciaunica2021, ciaunica2023b]. Each “thinks” in the enactive sense, interacting with its environment [thompson2007, varela2016, rolla2021]. Biological collectives can exhibit high level goal directed behaviour far beyond the capability of the parts of which they’re formed [reid2011].
The somewhat provocative claim this paper will make is that biological self-organising systems are more intelligent [^8] In the specific efficiency-based sense defined above. than contemporary, predominantly fixed-stack implementations of AI [^9] This is not a claim about all possible future AI. One aim of the paper is to identify design principles for building more adaptable hybrid and cyberphysical systems by shifting where adaptation occurs in the stack.. This is because biological self-organisation delegates adaptation to lower levels of abstraction [^10] For readers who want the generalisation of “levels of abstraction” beyond computing, see ESM Supplementary Note S2.. Formally, previous work showed adaptability in embodied systems can be measured by the ‘weakness’ of the constraints they embody [bennett2025d, bennett2024a]. Building on this, I prove a theorem (The Law of the Stack) that the ability to embody weak constraints at higher levels of abstraction depends on the ability to embody weak constraints at lower levels. This result shows how we can build more robust and efficient hybrid and cyberphysical systems, laying a foundation for future empirical research. Its applications extend beyond artificial intelligence, for example to human organisational and economic systems.
2 Abstraction, Distribution and Delegation
Conventional computing systems are organised into stacks of abstraction layers. More generally, I will characterise systems along three interacting dimensions: abstraction (how layers relate), distribution (how work is spread within a layer), and delegation (the level of abstraction at which adaptation or control resides). How efficiently a system adapts depends on the interaction between these dimensions.
In proposing Stack Theory [^11] Initially proposed in my PhD thesis [bennett2025thesis]. [bennett2025thesis], I merely extend this idea to systems in general. I’ll begin illustrating what I mean by abstraction, delegation and distribution; dimensions along which one may plot any system. For intuition, I will give examples of limit cases for computational, biological, human organisational and economic systems.
2.1 Abstraction
In a computer, a predictive model typically relies on a machine learning library, which runs on CUDA [^12] This is the software library that allows for the large scale utilisation of GPUs (a type of hardware originally developed for games), for machine learning and artificial intelligence., which runs on a GPU [^13] A graphical processing unit (GPU) is a type of hardware that allows for massively parallel vector processing. It was developed for the purpose of graphics processing in videogames, and later repurposed for machine learning..
GPUGPUCUDACUDAlibrarylibrarymodelmodel
These are all examples of abstraction layers. A programming language is a sort of abstraction layer. For example Python is an abstraction layer over C, in the sense that it is interpreted by a program written in the less abstract C programming language. C in turn depends on a program written in Assembly language, which runs on machine code which is interpreted by hardware. Each layer of language is written in the language of the layer below. This structure allows programmers to work more efficiently, within the confines of a standardised and simplified representation of the system. However, these abstraction layers are incomplete, simplified pictures of the system. They “abstract away” unnecessary details. Most importantly, they are embodied in the layer below. Each language is a subset of the behaviour that the layer below can exhibit. For example, a subset of the possible behaviour a given computer can exhibit would correspond to running Python code, and a broader subset of that computer’s behaviour would necessarily correspond to running C code. Hence each layer in this stack is a subset of the behaviours of the layers below [^14] A higher-level abstraction is not uniquely determined by its lower-level implementation: the same abstract interface can be realised by multiple implementations. The claim here is that higher-level behaviour must be realised by lower-level behaviour, and that different implementations can differ in resource cost (time, energy, information), which matters for adaptability.. This reductionist framing does not deny that useful abstractions can be robust to lower-level variation; invariants can be treated as higher-level task constraints. Even the physical hardware itself can be framed as an abstraction layer over the environment, being a subset of the behaviour of that environment in which the hardware exists [bennett2025thesis].
Think of a stack of abstraction layers as a directed graph. Highly abstract behaviours are a consequence of ‘weaker’, more generalisable goal directed behaviour at lower levels of abstraction. Framed this way, everything is a stack of abstraction layers [^15] This is not an analogical or metaphorical application of computational frames to others systems, as in Bratton’s re-framing of geopolitics [bratton2016]. Rather, this is a more general claim about how all systems work.. For example a collective of cells can enact an organ, and a system of organs can enact an organism [levin2024] which enacts even more abstract behaviours like language or tools. Each of these is an ‘abstraction layer’ in a biological stack [bennett2024a].
cellscellsorgansorgansorganismorganism
2.2 Distribution
In the case of biological systems lower levels of abstraction are formed of smaller or simpler parts [delgado1997], which is to say the work is distributed among parts at a level of abstraction. Because of this, looking at a lower level of abstraction is like ‘zooming in’ to look at a smaller scale. Conversely, rather than looking at low levels of abstraction like cells I can shift the level of abstraction to higher in the stack, and look at individual humans and how they form collectives at higher levels of abstraction. For example a population of soldiers can be a squad, and if squads co-operate they can be a platoon. Each layer above is the behaviour of a layer below, and the parts get smaller when I go down.
soldierssoldierssquadssquadsplatoonplatoon
In each layer work is distributed among members of the population. If level of abstraction is one dimension along which a system may be plotted, then distribution of work is another. In a computer I can distribute work by splitting it up into sub-tasks and allocating it to many computers.
However, unlike biological systems a lower level of abstraction in a computer is not necessarily made up of smaller parts that form a collective. Looking at a lower level of abstraction is not like ‘zooming in’. I can run software on a solitary computer, or a cluster of computers. For example, Python code is run by an interpreter, which is a C program that in turn is built on Assembly, which is built on machine code that has a fixed interpretation physically hard-wired in the instruction set architecture of the CPU. The scale of the smallest part at each layer of abstraction is the same. Think of one Python program running on one interpreter, on one computer. This illustrates how level of abstraction can vary without affecting the scale or distribution.
2.3 Delegated Control
The aforementioned stack on which Python relies is not a collective. It is not made up of parts capable of functioning independently. They do not have agency. This means control of the system resides at another level of abstraction. If I am programming a computer, then control resides at the level of abstraction at which I am programming. If I am programming in Python, then adaptation occurs at the Python level of abstraction [^16] This is a statement about where learning/adaptation is typically implemented in conventional software stacks. Lower layers can of course carry mutable state (data, caches, firmware state), but are usually not designed to be autonomously goal-directed in the way MCA components are..
Running a program in a distributed system does not always affect this. Distributing work is not the same thing as delegating control or agency. Work can be distributed among many GPUs, but if I remove a GPU from the computer it will not then pursue goals independently. It will sit, inert until once again integrated into a larger system. It requires no incentives. Control is imposed top-down, from higher to lower levels of abstraction. The parts at lower levels have no agency.
Conversely in a biological system each cell is independently capable of goal directed behaviour [ball2023], assuming it remains alive. Control in a complex biological system begins bottom-up, before top-down. Control then propagates in both directions [^17] In the sense that members of a collective are constrained by the collective, for example members of a human society are incentivised to obey laws. This is a form of top-down control. Bottom-up then becomes just whatever the individual humans do within those constraints, even though the system as a whole might have emerged entirely bottom-up from anarchy to begin with.. Each part of a collective acts on its own, and its actions imply a ‘policy’. Each individual policy determines how that part interacts with the other parts, or more to the point their policies. These interactions cause the system as a whole to act, which implies a collective policy. If collective and individual policies align in service of homeostasis, then the system can persist. It has a collective identity or goal. That identity can constrain parts of the collective. For example, I might say the level of abstraction is cells, and then describe an organ as a collective policy of those cells. Likewise, I might look at the level of organs, and describe a human as a collective policy of those organs. Or I could describe the human as the collective policy of cells, and see the organs as a sort of middle management layer.
A good example to understand the limit cases for delegation is an economy. In an absolutely centrally planned economy, control is concentrated at the highest level of abstraction. One might think of the people in the economy as parts at lower level of abstraction. They do work but have no control. Those people could adapt, but central planner’s edicts constrain them. The opposite extreme would be to have control delegated entirely to the lowest level of abstraction, in anarchy unconstrained by any laws.
The level of abstraction is often related to the granularity at which the system is represented. For example if the level of abstraction at which I view an economic system is humans, then I don’t worry about how work is distributed or control is delegated within those humans. I just concern myself with what systems can be formed of humans, like economies or armies (see fig 1). Of course, there are limits on what sort of physical forms can store or process information in service of a goal [sole2024], and this may limit how much I can distribute work or delegate control. Delegating to a lower level of abstraction often requires distributing work among parts at smaller scales [^18] For example, delegating control to the level of individual humans would mean those humans control their own work potential, but not that of others, or else we’d just be delegating to collectives within the human population.. However as the static abstraction layers in a computer illustrate, a lower level of abstraction is not always more granular, and thus work is not always more distributed (see fig 2).
Figure 1: Control or agency and thus adaptation in a given system can take place at a particular level of abstraction. For example, I can program in Python, or I can go down a level and program in C. If I program in C it doesn’t matter whether I program one computer or fifty. Either way my control is at the same level of abstraction (figure from [bennett2025thesis]).
Figure 2: Work, resources and so on can be distributed across a system. Distribution is distinct from delegation of control. For example, both food stamps and universal basic income might distribute the same amount of resources to the same people within an economic system, but the former more tightly constrains the recipients in how they use those resources (figure from [bennett2025thesis]).
2.4 Reasons to Suspect Delegation Facilitates Adaptation
I will now give a brief explanation of why I believe delegation of control to lower levels of abstraction is relevant to adaptation.
Throughout this subsection I use examples (e.g. distributed drone swarms, mission command, and economies) as intuition-building thought experiments. They are not intended as comparative empirical evidence for the formal claim, which is established later in Theorem 5.1. All else being equal the lower the level of abstraction at which the system can be controlled, the more it can be specialised. For example a program written in C can be far more efficient than a program written in Python [oden2020]. This is because the Python program has to be interpreted by the Python interpreter, which is a program written in C. By programming in C I simply cut out an abstraction layer. This gives the programmer more leeway to optimise the system, or make mistakes. Likewise another way to make computers more efficient is to use a FPGA to write a program at the hardware level, instead of compiling that program into machine code and running it on a one-size-fits-all instruction set architecture like x86 [nguyen2020].
With each successive abstraction layer introduced, one is further removed from the ground truth to which one must adapt. For example software is nothing more than a state of hardware [^19] That software is a state of hardware has been addressed at length elsewhere [bennett2024a, bennett2025thesis], beyond the scope of this paper.. It is a label I ascribe to the behaviour of hardware. It is a useful simplification that lets us consider what I want hardware to do in the abstract, but working at such a high level of abstraction can cause problems. Proposals for software ‘intelligence’ have been undermined by the fact that any claims regarding the behaviour of software are contingent upon hardware. The software doesn’t interact with the environment, the hardware does. This is because the ability to adapt is related to perceived complexity, which is entirely determined by the abstraction layer through which an agent interacts with its environment [bennett2024b]. Consequently an ‘objective’ measure of artificial intelligence [leike2015, bennett2024a, bennett2025b] must be ‘enactive’ [thompson2007] in that it hinges not just software but hardware and the local environment as well. To understand the difference between biological systems and artificial intelligence, I must account for the level of abstraction at which an adaptive process takes place.
In distributed computational systems, like a drone swarm or network, delegating control can facilitate what is called ‘graceful degradation’ [kuniavsky2010]. For example, to control a drone swarm I could single out one drone to be the ‘leader’, and have it run a program that co-ordinates the other drones. However that would make the swarm vulnerable. If the leader drone were damaged, the swarm would de-cohere. It is a single point of failure. In contrast if I can get the swarm to co-ordinate through entirely decentralised control, delegating to the level of individual drones, the swarm will remain coherent regardless of how many drones are disabled [^20] This sort of decentralised swarm control can also be observed in nature, such as with flocks of starlings which fly in complex formations [cavagna2010].. In other words, the swarm ‘adapts’ to adverse conditions.
The advantages of delegating control are visible in human organisations as well. In military doctrine, “mission command” delegates decision-making to units closest to the situation, granting latitude to deviate from detailed orders in service of higher-level intent [Shamir01102010]. Conceptually, each layer of management is an abstraction layer: when control is delegated, feedback loops shorten because fewer layers sit between the controlling entity and local conditions. This supports the broader hypothesis that delegation is a key ingredient of biological self-organisation that is largely absent from mainstream contemporary artificial intelligence. See ESM Note S7 for a concrete historical illustration (Nordbat 2).
3 What is Adaptation?
I will discuss what it really means to learn and adapt, in terms of cause and effect.
causeeffect
To make accurate inferences a system must be able to tell the difference between what it has caused, and what it has merely observed [merker2005, pearl2018, dawid2002]. It is often difficult to determine cause. ‘Confounding’ occurs when an event appears to be the cause of the effect we’re trying to predict, when in truth it is merely correlated. This can lead one to draw erroneous conclusions.
not the causeeffectcause
By actively intervening in the environment, a system can experiment to establish cause [^21] Assuming such experimentation is enough to mitigate ambiguity and eliminate unknowns, and that a heuristic such as weakness is applied to ensure the system converges on variable definitions which represent cause accurately [bennett2023c].. Then, it can construct a causal graph. There are two ways to do that. One option would be to assume the nodes (variables, actions and other abstract objects and properties), and then learn the causal relations between them. This might work well in theory [richens2024robust], but it requires foreknowledge of objects and properties between which causal relations can be established. The alternative is to assume the causal relations (the edges in the graph), and then learn the variables. This needs only attraction and repulsion from physical states [bennett2023c, bennettwelshciaunica2024, bennett2025thesis], otherwise known as valence. I take this latter approach.
Learning nodes to fit a graph is like figuring out which computer program [^22] Computation in this paper is about constraints, like how slime mold can be made to “compute” by imposing the right constraints on them[nakagaki2000]. I’m not assuming Turing machines. was used to generate a sequence. Assume I have fed a set of inputs into a computer. For each input the computer responded with a particular output. The computer runs a ‘policy’ program which maps inputs to outputs [^23] I call such programs policies, but this should not be confused with the more specific definition policy in reinforcement learning [sutton2018].. That policy is the cause of the input-output history observed so far. I want to know the policy that caused historical outputs so that I can use it to predict future outputs. The problem is that more than one policy could have produced the history we’ve observed, and each of those policies produce different outputs when I give the computer inputs it hasn’t encountered before. From a naive Bayesian perspective, all of them are equally likely to have caused the history I saw. If I wish to know which policy actually caused the outputs, I need a way to whittle down the remaining possibilities.
There are two ways to compare policies that look the same with respect to historical data, but differ with respect to possible future data [bennett2024b]. One is to compare their form, and choose policies which are simpler in the sense of being more compressible. However, in an interactive setting simplicity is an unreliable indicator of plausibility [leike2015, bennett2024b]. The alternative is to compare policies by their function, in terms of how they behave given inputs for which I do not have historical data. Adaptive systems must satisfy functional constraints [sole2022]. I can compare systems by these constraints. Each policy implies a constraint on possible outputs given a set of possible inputs. The weaker that constraint is, the more variety of future behaviour it permits. Put another way, if two policies satisfy a goal, then the weaker policy satisfies it in more possible worlds [^24] To the extent that such possible worlds can be differentiated from one another within the subjective confines of an abstraction layer.. Previous research showed that if two policies are identical with respect to historical data, then the policy most likely to have generated historical data is the weaker [^25] Formal definitions and proofs for weakness and w-maxing are in [bennett2024appendices, bennett2025thesis]; an extended intuition is in ESM Supplementary Note S3. of the two [bennett2023c, bennett2025d, bennett2025e]. The policy most likely to have caused behaviour is the weakest [bennett2023c], and in experiments a system which preferred weaker policies outperformed one that preferred the simpler policies by $110-500\%$ [bennett2023b]. This has been used to establish an upper bound on embodied ‘intelligence’ [bennett2024a]. I will use the term ‘w-maxing’ to mean a system optimises for weak policies, and ‘simp-maxing’ to mean a system optimises for simpler policies [^26] Basically minimum description length policies [rissanen1978]., to conform to the terminology used in previous work [bennett2025b, bennett2025thesis] [^27] For formal definitions of learning and efficiency used in this framework, see [bennett2024appendices] and the thesis [bennett2025thesis].. Put another way, if two hypotheses appear equally good with respect to past data, the w-maxing system will use policy weakness as a heuristic to decide which policy to use, whereas a simp-maxing system will use simplicity.
An abstraction layer can be understood as a border between form and function. For example a program takes a form when it is written in a language, but when it is run on a computer it has a particular function or effect upon the world [^28] For example, hardware like an Apple M2 system on a chip interprets software to physical effect, based on the ARM instruction set architecture.. Simp-maxing is a matter of form, while w-maxing is a matter of function [bennett2025b]. When we account for interpretation between form and function it becomes clear that software is a physical thing, like a tool. A tool can serve more than one purpose. For example, a hammer is both a weapon and a paperweight. This is what “weakness” measures. A tool that implies a weaker constraint can complete more tasks [bennettmaruyama2022a] than another tool [^29] To be exact, a weaker policy completes a superset of tasks, rather than just more tasks. This is because policies are compared in the context of an ostensive definition.. By preferring weaker policies, one avoids unnecessary commitments, meaning the weakest policies complete a ‘wider range of tasks’ than others [^30] See ESM Supplementary Note S3 for extended discussion of Bennett’s Razor, weakness, and related connections (e.g. maximum entropy).. To complete a ‘wider range of tasks’ in the embodied sense, a tool must be both more sample and more energy efficient, and so the weakest policy is the most efficient in both senses. Put another way, the same physical matter can simultaneously contribute to many different tasks or computations superimposed over each other at different scales. A ‘weaker’ policy serves more masters.
ML systems do not typically optimise for either weaker constraints on function (w-maxing), or simplicity of form (simp-maxing). This may explain some of their relative inefficiency [bennett2022a, bennett2023b]. Even if they were w-maxing, the capacity for w-maxing is severely limited by the abstraction layer. An abstraction layer can be a “causal isolator” [friston2023] between a software “mind” at higher levels of abstraction, and its environment at lower levels. Embodiment by another name. Different abstraction layers permit different upper bounds on policy weakness, in the same way that two bodies might differ in how able they are to achieve a given goal. Different embodiments give different upper bounds on ‘intelligence’ [bennett2024a].
That biological self-organising systems adapt efficiently means they are able to embody weaker policies than today’s ML systems. This capacity hinges on the abstraction layer [bennett2024a]. An abstraction layer is like a window through which only part of the world can be seen. It conveys an inductive bias that helps or hinders adaptation. When I optimise the system at a lower level of abstraction, this is like saying I can “move” this window to convey a different inductive bias. The bigger the window, the “weaker” the policy it is possible to learn. What this means is that intelligence is primarily a matter of the lens through which we interact with a problem. Hardware, not just software. In enactivist terms, an intelligent choice of abstraction layer is one that reduces the “big world” of all information to a “small world” of relevant information [vervaeke2012, vervaeke2013a, vervaeke2013b]. In a typical computer, lower levels of abstraction tend to be static. The window cannot move. I will show a static abstraction layer limits adaptability at higher levels of abstraction by limiting how weak the policies embodied by the system can be. Adaptability at high levels requires adaptability at low levels, and delegation of control facilitates this.
4 Formal Framework and Definitions
The formal definitions below are required for the core claim and for Theorem 5.1. The surrounding narrative provides intuition, but the theorem and its assumptions depend on these definitions. However this paper is based on a formal mathematical framework and it does contain a very simple proof supporting the claim. The formal details are also useful to connect this work here to related mathematical and experimental work [bennett2025thesis, bennett2023b, bennett2023c, bennett2024a, bennett2024b, bennettwelshciaunica2024, bennett2025e].
Stack Theory describes an adaptive system or computation by the binary constraints or boundary conditions that define it. It formalises embodiment as a choice of abstraction layer: a finite vocabulary of possible constraints (programs) representing the possible configurations of a body embedded in its environment [bennett2025thesis], which are used to express “statements” constraining the body. Policies and tasks are then defined using statements in this embodied language. Delegation, distribution and abstraction are captured by the task hierarchy. A longer conceptual motivation (including the environment/states/programs framing and links to enactivist and pancomputational interpretations) is provided in the Electronic Supplementary Materials (ESM) Note S5 and is not required to follow the theorem below.
For example, if a human raises their arm, then they are expressing a statement in their embodied language. That statement constrains what the body does next. The vocabulary $\mathfrak{v}$ of programs express this physical constraint. The inputs, outputs and policies of a system are all just subsets of its vocabulary. A $\mathfrak{v}$-task $\langle I,O\rangle$ is a set of inputs and corresponding outputs. If the $\mathfrak{v}$-task describes a ‘fit’ organism, then the outputs it contains are ‘correct’ according to natural selection. A system [ciaunica2023b] that conformed to the free energy principle (FEP) would be a $\mathfrak{v}$-task for which $\langle I,O\rangle$ minimised surprisal. In other words, a $\mathfrak{v}$-task can describe an organism as behaviour that serves homeostatic and reproductive goals.
For readers who wish to check the formal dependencies (see also [bennett2025thesis, bennett2024appendices]), the key definitions used in the theorem are summarised below.
  1. 1.
    A set $\Phi$ is assumed, whose elements are called states. A declarative program is a set $f$ of states, and $f$ is true relative to $\phi\in\Phi$ iff $\phi\in f$. $P=2^{\Phi}$ is the set of all programs. Embodiment is formalised as an abstraction layer, for which I single out a finite set $\mathfrak{v}\subseteq P$ called a vocabulary. A choice of vocabulary $\mathfrak{v}$ implies a formal language
    $L_{\mathfrak{v}}=\{l\subseteq\mathfrak{v}:\bigcap l\neq\emptyset\}.$
  2. 2.
    Members of $L_{\mathfrak{v}}$ are called statements, and a statement $l$ is considered true relative to a state $\phi$ iff $\phi\in\bigcap l$. A completion of a statement $l_{1}$ is any statement $l_{2}$ s.t. $l_{1}\subseteq l_{2}$.
    The extension of a statement $l$ is the set $E_{l}$ of all its completions, s.t. $x\in L_{\mathfrak{v}}$ has
    $E_{x}=\{y\in L_{\mathfrak{v}}:x\subseteq y\}.$
    The extension of a set $X\subseteq L_{\mathfrak{v}}$ of statements is
    $E_{X}=\bigcup\limits_{x\in X}E_{x}.$
    The weakness of a policy $x$ is the cardinality of its extension $\lvert E_{x}\rvert$.
  3. 3.
    A $\mathfrak{v}$-task $\alpha$ is a pair $\langle I_{\alpha},O_{\alpha}\rangle$ where $I_{\alpha}\subset L_{\mathfrak{v}}$ are the inputs, $E_{I_{\alpha}}$ is the set of outputs available given $I_{\alpha}$, and $O_{\alpha}\subset E_{I_{\alpha}}$ are the correct outputs. All tasks form a hierarchy. If $\alpha\sqsubset\omega$ then $\omega$ is a parent of $\alpha$, meaning $I_{\alpha}\subset I_{\omega}$ and $O_{\alpha}\subseteq O_{\omega}$. A parent is higher level than its children. A parent task is delegated to child tasks, and as each task is embodied this describes control being delegated within a system.
  4. 4.
    A policy is a statement $\pi$ and a policy is correct for $\alpha$ iff
    $\pi\in\Pi_{\alpha}=\{\pi\in L_{\mathfrak{v}}:E_{I_{\alpha}}\cap E_{\pi}=O_{\alpha}\}.$
    Given input $i\in I_{\alpha}$ and a policy $\pi$, to infer is to choose $o\in E_{i}\cap E_{\pi}$. A task is complete if $o\in O_{\alpha}$. Given a task definition $\alpha$ s.t. $\alpha\sqsubset\omega$, to learn is to choose a policy $\pi\in\Pi_{\alpha}$, and $\omega$ has been learned from $\alpha$ if one chooses $\pi\in\Pi_{\omega}$.
  5. 5.
    If $\alpha_{1}\sqsubset\alpha_{2}\sqsubset\omega$ then $\alpha_{1}$ contains fewer examples of $\omega$ than $\alpha_{2}$. System $\mathfrak{a}$ is more intelligent than $\mathfrak{b}$ if it learns $\omega$ from fewer examples. For example $\mathfrak{a}$ is more intelligent than $\mathfrak{b}$ if $\mathfrak{a}$ learns $\omega$ from either $\alpha_{1}$ or $\alpha_{2}$, while $\mathfrak{b}$ learns $\omega$ from $\alpha_{2}$ but not from $\alpha_{1}$.
  6. 6.
    A causal identity for $\mathfrak{b}$ is a policy classifying the interventions undertaken by $\mathfrak{b}$.
  7. 7.
    $\Gamma_{P}$ contains every task in every vocabulary. For every $\rho\in\Gamma_{P}$ there exists an uninstantiated-task function $\lambda_{\rho}:2^{P}\rightarrow\Gamma_{P}.$ $\lambda_{\rho}$ takes a vocabulary $\mathfrak{v}^{\prime}\in 2^{P}$ and returns the highest-level child of $\rho$ that is a $\mathfrak{v}^{\prime}$-task.
  8. 8.
[…]
I set out to compare computers with biological systems. Since both are stacks of abstraction layers, I define multilayer architecture (MLA) below to describe them in comparable terms.
  • The stack of layers of increasing abstraction is a sequence of uninstantiated tasks $\langle\lambda^{0},\lambda^{1},\dots,\lambda^{n}\rangle$ s.t. $\lambda^{i+1}\sqsubset\lambda^{i}$. Lower in the stack is a lower level of abstraction.
  • $\mathfrak{f}:2^{P}\times L_{\mathfrak{v}}\rightarrow 2^{P}$ is a function that takes a vocabulary $\mathfrak{v}$ and a policy $\pi\in L_{\mathfrak{v}}$, and returns a vocabulary
    $\mathfrak{v^{\prime}}=\{f\in P:\exists o\in E_{\pi}(\bigcap o=f)\}.$
  • The state of an MLA is a sequence of policies $\langle\pi^{0},\pi^{1},\dots,\pi^{n}\rangle$ and a sequence of vocabularies $\langle\mathfrak{v}^{0},\mathfrak{v}^{1},\dots,\mathfrak{v}^{n}\rangle$ such that
    $\mathfrak{v}^{i+1}=\mathfrak{f}(\mathfrak{v}^{i},\pi^{i})\quad\text{and}\quad\pi^{i}\in\Pi_{\lambda^{i}(\mathfrak{v}^{i})}.$
I say the MLA is over-constrained when there exists $i<n$ such that $\Pi_{\lambda^{i}(\mathfrak{v}^{i})}=\emptyset$, and multilayer-causal-learning (MCL) is when the MLA is not over-constrained.
5 Delegated Control and Collective Identity
The vocabulary $\mathfrak{v}$ of an abstraction layer can describe a distributed system. For example, it could describe an organ $\omega$ formed of two cells $\alpha_{1}$ and $\alpha_{2}$. The behaviour of two cells as individuals is formalised as two $\mathfrak{v}$-tasks $\alpha_{1}$ and $\alpha_{2}$. The collective behaviour of the two cells is a third $\mathfrak{v}$-task $\omega$ s.t. $\alpha_{1}\sqsubset\omega$ and $\alpha_{2}\sqsubset\omega$. Intuitively, this just means the two cells are parts of a larger whole. If there exists a correct policy for $\omega$, then the cells can behave in service of a common goal and thus have a common identity (for example if both cells use policies in $\Pi_{\omega}=\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$).
$\alpha_{1}$$\alpha_{2}$$\omega$
The probability that the parts of a system embody a collective policy is higher if they embody weaker policies. This is because weaker policies are more likely to generalise between $\mathfrak{v}$-tasks [bennett2025e].
Consider the aforementioned $\alpha_{1}$ and $\alpha_{2}$. A sufficiently weak policy at the level of individual cells is necessary for the collective to have an identity, because only the weakest policies in $\Pi_{\alpha_{1}}$ and $\Pi_{\alpha_{2}}$ will also be in $\Pi_{\omega}=\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$. A collective of cells remains cells regardless of whether I deem it to be an organ. The parts are still there. In terms of the stack, the cells $\alpha_{1}$ and $\alpha_{2}$ and the collective $\omega$ share a vocabulary. Just as the outputs of $\alpha_{1}$ can affect the inputs of $\alpha_{2}$ and vice versa, the outputs of $\omega$ determine the inputs of $\alpha_{1}$ and $\alpha_{2}$. The difference is a matter of scale. Just as w-maxing facilitates causal learning at the scale of an individual [bennett2023c], in the context of an MCA it will facilitate learning at multiple scales, because the system as a whole is just another $\mathfrak{v}$-task.
Mechanistically, “negotiation” between levels is represented as constraint compatibility between policy sets. Individual components select policies that remain correct for their local tasks while being consistent with the collective task, and the collective task in turn constrains components by shaping the inputs they experience. When the intersection of viable policies becomes empty, no collective policy exists and collective identity collapses (a failure mode analysed in Section 6). In hybrids, the designer mediates this negotiation by adjusting boundary conditions (which reshape the component policy sets) rather than by directly specifying a single low-level policy.
5.1 The Stack
From the enactive perspective, a human is something cells do, rather than are. An organ boils down to a collection of cells. The cells are an abstraction layer in which the collective policy of the organ is expressed, like software. However, software is just a state of hardware [bennett2024a]. Put another way, ‘organ’ is a label I ascribe to the the behaviour of some cells, and ‘human’ is a label I ascribe to the behaviour of some organs. ‘x86’ is a label I ascribe to a product of human behaviour. It is a static abstraction layer an engineer has embodied in silicon and then left alone; ‘fire and forget’. Humans act as the lower layer from which x86 emerges, but once it is embodied in silicon we disconnect it from our stack. We causally ‘intervene’ in the stack. To understand the effect of this intervention, I need to abstract away the lower levels. I must “abstract away” cells and focus on the consequences of their collective behaviour.
The cells are tasks. For tasks at higher levels of abstraction, capacity for w-maxing depends on lower levels. A policy $\pi\in\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$ implies an abstraction layer in the sense that for every $o\in E_{\pi}$ there exists $f\in P$ s.t. $\bigcap o=f$ [^31] I can substitute $o$ for a declarative program $f$ that has the same extension, because that means $f$ and $o$ are true of the same states, having the same ‘truth conditions’ so to speak.. Hence I can transform policies to abstraction layers, using a function $\mathfrak{f}$ that takes the extension $E_{\pi}$ of a collective’s policy $\pi$, and returns a new vocabulary $\mathfrak{v^{\prime}}=\{f\in P:\exists o\in E_{\pi}(\bigcap o=f)\}$. I can then use a sequence of uninstantiated tasks $\langle\lambda^{0},\lambda^{1}...\lambda^{n}\rangle$ s.t $\lambda^{i+1}\sqsubset\lambda^{i}$ to represent selection pressures that define ‘correctness’ at every level. At each successive level, ‘correct’ gets more specific. Because everything that happens in a higher level is in some sense correct according to lower levels, each higher level must more tightly constrain what is considered to be correct behaviour if that higher level is to exist in a meaningful sense.
$\lambda^{i-1}$$\lambda^{i}$$\lambda^{i+1}$
To construct layers I would start with a vocabulary $\mathfrak{v}^{i}$ at $i=0$, for example implied by a collective of cells or a lower level system. If this system exhibits goal directed behaviour, then it is given by the $\mathfrak{v}^{i}$-task $\lambda^{i}(\mathfrak{v}^{i})$ and a policy $\pi^{i}\in\Pi_{\lambda(\mathfrak{v}^{i})}$. $\pi^{i}$ implies behaviour. I translate policy into an abstraction layer using the aforementioned $\mathfrak{f}$. $\mathfrak{f}$ is applied to $E_{\pi^{i}}$ to obtain a vocabulary $\mathfrak{v}^{i+1}$, and then $\lambda^{i+1}$ is applied to $\mathfrak{v}^{i+1}$ to obtain a task $\lambda^{i+1}(\mathfrak{v^{i+1}})$. If $\Pi_{\lambda^{i+1}(\mathfrak{v^{i+1}})}\neq\emptyset$, then there exists a common purpose or “self” at this level of abstraction and I select a another policy $\pi^{i+1}\in\Pi_{\lambda^{i+1}(\mathfrak{v^{i+1}})}$.
In purely narrative terms, the implication of all this is that adaptation at higher levels depends on adaptation at lower levels. Below I state The Law of the Stack with a brief proof sketch; a full proof is provided in the Electronic Supplementary Materials.
Theorem 5.1 (The Law of the Stack).
Let $\langle\lambda_{0},\lambda_{1},\dots,\lambda_{n}\rangle$ be an MLA with abstractor $\mathfrak{f}$ and state $\langle(\pi_{0},\dots,\pi_{n}),(v_{0},\dots,v_{n})\rangle$ such that $v_{i+1}=\mathfrak{f}(v_{i},\pi_{i})$ and $\pi_{i}\in\Pi_{\lambda_{i}(v_{i})}$ for all $i<n$. Fix any $i<n$, and write
$v_{i+1}=\mathfrak{f}(v_{i},\pi_{i}),\qquad\gamma_{i+1}=\lambda_{i+1}(v_{i+1}).$
Then the achievable utility at layer i+1i\!+\!1 is upper-bounded by the weakness of the policy at layer $i$ in the following sense:
$\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\;\leq\;2^{|E_{\pi_{i}}|}.$
Equivalently,
$|E_{\pi_{i}}|\;\geq\;\log_{2}\!\big(\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\big).$
In particular, increasing $\epsilon(\lambda_{i+1}(v_{i+1}))$ requires (at least) increasing $|E_{\pi_{i}}|$, i.e. weakening the policy at the lower layer.
Proof sketch.
Utility $\epsilon(\gamma_{i+1})$ is attained by some correct policy at layer i+1i\!+\!1, whose extension is a subset of $L_{v_{i+1}}$, so $\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\leq|L_{v_{i+1}}|$. Because $v_{i+1}$ is finite, $|L_{v_{i+1}}|\leq 2^{|v_{i+1}|}$. By construction of the abstractor, $v_{i+1}$ is the image of $E_{\pi_{i}}$ under the map $o\mapsto\bigcap o$, hence $|v_{i+1}|\leq|E_{\pi_{i}}|$. Combining yields $\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\leq 2^{|E_{\pi_{i}}|}$. ∎
Intervening in the stack limits the extent to which control can be delegated, and thus limits capacity for w-maxing or MCL. It makes lower layers of abstraction static in the sense that they no longer change to facilitate adaptation at higher levels. By Theorem 5.1, any increase in achievable utility at layer i+1i{+}1 requires (at least) an increase in the extension size of the realised policy at layer $i$, i.e. support for weaker constraints below.
5.2 Human Abstractions
Humans are not just part of MCAs. A human is an MCA. Human abstractions are emergent behaviours of that MCA.
cellscellshumanshumansabstractionsabstractions
When we humans create a computer, we create a machine that embodies a subset of our more abstract behaviours. An engineer hard-wires those behaviours into the instruction set architecture (ISA), and we all then ascribe meaning to that behaviour. Two results can look the same at the higher level (e.g. the number two, spoken by a human or displayed on a screen), but be fundamentally different at the lower level (e.g. resources consumed, data format etc). The results are only the same if one ignores the lower levels of abstraction.
Software might be agentic, but computer hardware is passive. Control is top-down. Once I have built an ISA, it exists in the environment independently of the human MCA from which it emerged. It is static. It cannot adapt, which is why there is demand for field programmable gate array (FPGA) [^32] An FPGA is like a CPU that can be reconfigured for specialised applications..
humanshumansISAISAabstractionsabstractions
If an abstraction layer is a window onto the world, then utility is how “good” the view is for a task. A static abstraction layer is a window that cannot move, while an MCA can change policy at lower levels, to move the window into an ideal location. When I create an ISA I insert a static abstraction layer at an intermediate level, so the vocabulary at the highest levels is limited to what I can do with those abstractions I have embodied in the ISA. In comparison, an MCA can learn policies at lower levels, to change the vocabulary of higher levels and facilitate multilayer causal learning. This shows how adaptability at higher levels demands adaptability at lower levels. I am not suggesting lower levels only optimise for utility at higher levels. After all, an increase in utility may increase the computational cost of searching for correct policies. Furthermore, a spatially extended environment constrains the size of vocabularies at all levels [bennett2024b], because bounded systems can contain only finite information [bekenstein1981]. I mean lower levels learn an inductive bias for higher levels, balancing utility with other factors in service of goals at every scale.
6 Cancer-Like Failure Modes
I have explored some advantages of bio-mimetic distribution and delegation. Are there disadvantages? Yes. A state analogous to cancer exists for distributed systems with insufficient delegation in adverse conditions. Some have argued cancer is a loss of collective identity [levin2021, davies2011]. In a biological system, when cells are isolated from the informational structure of their collective, they lose their high level identity. When that happens, they independently pursue their own lower level goals. They revert to primitive transcriptional behaviour, while the collective of which they were formerly part continues on without them. They escape from immune control mechanisms by becoming less differentiated, and closer to embryonic states.
In the context of my formalism, a state analogous to cancer occurs when a system is ‘over-constrained’ so that no collective policy exists, while the capacity for independent action remains. For example, assume a collective represented by a task. Adverse environmental conditions are represented as a task for which fewer correct policies exist than less adverse environmental conditions. Likewise excessive top down control can limit adaptability, and thus the number of correct policies which exist. If the constraints are too tight, then no correct policies will exist and the collective’s identity will be lost.
The set of correct policies shrinks as task difficulty increases, meaning $\mathfrak{v}$-task $\alpha$ is easier than $\omega$ if $\mid\Pi_{\alpha}\mid>\mid\Pi_{\omega}\mid$. If the policy at a lower level i1i-1 changes such that the task at level $i$ becomes ‘harder’333333i.e. by reducing $\mid\Pi^{i}\mid$ at the higher level $i$., then this could result in there being no correct policy at level $i$. If there are no longer any correct policies at level $i$, then the only way the system can continue to function is for part of the collective to ‘splinter’ or break away from the group, because that will loosen the constraint on correct policies for both the breakaway part and the collective it leaves behind. Say I have a collective of cells $\lambda^{i}(\mathfrak{v}^{i})$ that includes a cell $\alpha\sqsubset\lambda^{i}(\mathfrak{v}^{i})$ s.t. $\mathfrak{f}(\mathfrak{v}^{i-1}_{\mathfrak{a}},{\pi^{i-1}_{\mathfrak{a}}})=\mathfrak{v}^{i}_{\mathfrak{a}}$. Now assume the policy $\pi^{i-1}_{\mathfrak{a}}$ changes to $\pi^{i-1}_{\mathfrak{b}}$ s.t. $\Pi_{\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})}=\emptyset$. A subset of the collective may continue functioning as before with a shared policy, but doing so requires discarding a part of the collective which cannot adopt that policy. This is akin to ‘selective forgetting’ [bennett2023c] of possible false negatives or positives in order to interpolate otherwise noisy data, but instead of forgetting data inconsistent with an otherwise good policy, I would discard a child task $\alpha\sqsubset\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})$ of $\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})$ when $\Pi_{\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})}=\emptyset$, so that I obtain a task $\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})\sqsubset\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})$ s.t. $\alpha\not\sqsubset\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})$ and $\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})$ has correct policies, meaning $\Pi_{\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})}\neq\emptyset$. $\alpha$ is now isolated from the informational structure of the collective and so may have an extremely broad set of ‘correct’ policies to choose from.
This is analogous to cancer. $\alpha$ is isolated and can now pursue goals which do not align with those of the collective. This suggests systems should be minimally constrained so as to ensure graceful degradation or ‘sloppy fitness’ [agosta2010], maintaining core functionality in as wide a variety of unanticipated circumstances as possible. The more difficult the environment or the more top-down the system is controlled, the more tightly constrained behaviour becomes until the system becomes over-constrained and splinters. Co-operation becomes impossible and the parts of a system start to compete. Conversely, if control of a system is delegated to the lowest levels, then adverse conditions are less likely to eliminate all correct policies. Put another way, there will be less incentive for the parts of a system to splinter, as that would not greatly increase the number of correct policies available to them. To summarise; a state analogous to cancer occurs in a distributed system when there is an over-reliance on top-down causation to make the parts of a system conform to a collective identity they are ill suited to, under the circumstances.
7 Hybrid Agency as Boundary-Condition Design
Consider the aforementioned example of the organoid adapted to play the game Pong with notable sample efficiency relative to conventional reinforcement-learning baselines [khajehnejad2024]. Hybrid systems like organoids are not fixed interpreters of software-level policy, but collectives whose lower-level parts retain bottom-up agency. In such systems, hybrid creation is the act of designing an interface and environment that reshapes the feasible policy space of the biological components. Rather than “programming” the parts directly, the designer engineers boundary conditions so that the remaining viable collective policies implement a target input–output relation at the interface.
The designer selects boundary conditions that reshape each component’s viable policy set, hoping that the intersection admits a collective policy that realises the desired input–output mapping at the interface.
According to The Law of the Stack, achieving higher utility at an interface layer requires sufficiently weak policies at lower layers. Therefore, boundary conditions should be as weak (least exclusionary) as possible while still making the target behaviour achievable and preserving a non-empty intersection of component policy sets (collective identity). Overly tight constraints can collapse that intersection and force failure, while overly loose constraints may fail to stabilise any desired collective policy. A schematic formalisation of this boundary-condition design view is provided in the ESM.
8 Conclusions and Future Research
In this paper I used Stack Theory to describe MCAs as stacks of abstraction layers. In this view policies are embodied at each layer, approximating causality in service of adaptation, and cause is best approximated by weaker policies (w-maxing) [bennett2023c]. Formally, I prove a theorem (The Law of the Stack 5.1) showing w-maxing in higher layers requires sufficient w-maxing in lower layers. One reason artificial intelligence tends to be less adaptable is because it rests upon an inflexible stack of human abstractions. This claim is directed at contemporary, predominantly fixed-stack AI architectures. It is not a fundamental limitation on future computational or hybrid systems.
My claim is that, to build more robust systems, control should be delegated to the greatest extent possible. Obviously this must be balanced against desired functionality. Policies must be aligned. The system must still discriminate between correct and not. In the earlier sections I discussed the doctrine of mission command. What I am suggesting is that the concept of mission command should be applied to systems in general. At least, to the extent that adaptability is desired. Put another way, any constraints which are not strictly necessary to maintain the integrity of the system should be eliminated [^34] What counts as “strictly necessary” is relative to design goals and value judgements; see ESM Supplementary Note S4.. Humans tend to prefer ‘additive’ solutions [adams2021], by which I mean if a problem can be solved by adding a rule or subtracting a rule, we humans tend to add a rule. To effectively delegate control, we should subtract rules instead.
There are also software-first approaches that approximate aspects of this delegation picture, including federated learning [nori2025federated], constructivist AGI programmes [thorisson2012, goertzel2014], and active inference [Friston2010, friston2013]. An expanded discussion (including AERA, Hyperon/ActPC-Chem, and links to scale-free active inference) is provided in the ESM (Supplementary Note S6).
Future research might consider how adaptive hardware can be designed, to delegate control. Multilayer-causal-learning (MCL) provides a formal basis for ‘soft robotics’ [man2019]. Soft robotics is a proposed class of robots made of ‘soft’ materials that adapt. I have formalised adaptive materials as adaptation lower in the stack, providing a formal foundation for soft robotics. Self-assembling nanoparticle networks [borghi2024_brainlikehw] are a step in this direction. They can be used to form reconfigurable logic gates [paroli2023_receptron], facilitating adaptation at very low levels of abstraction. Taken to its logical conclusion, what might be required is a sort of artificial life, with small parts interacting form complex systems from the small scales and low levels of abstraction up [suzuki2006, ikegami2008, ikegami2007simulating].
As for the inductive bias mentioned in the introduction [chollet2019], it boils down to the choice of abstraction layer [bennett2024a]. I have described a means of learning a suitable abstraction layer. In theory, a multilayer-causal-learning system should be sufficient to solve ARC-AGI benchmark problems in terms of both sample and energy efficiency. Future research should test this empirically through (i) agent-based simulations of collectives with tunable delegation and abrupt environment shifts, measuring sample/energy efficiency and robustness; (ii) comparative benchmarks in which adaptation is confined to high-level software state versus systems with reconfigurable substrates (e.g. adaptive firmware or FPGAs); and (iii) hybrid experiments in which boundary conditions on biological components are systematically varied to quantify trade-offs. A more detailed set of candidate evaluation designs is provided in ESM Note S8.
Beyond hardware, cyberphysical systems in general can be evaluated in terms of abstraction, distribution and delegation. I might consider how states analogous to cancer could arise in human organisational structures, or how control can be delegated throughout an organisation to facilitate robustness and efficiency. For example, this could be applied to facilitate transformational bottom-up rather than incremental top-down adaptation to environmental changes.
Free-market trade is a great example of how delegation can facilitate adaptability in a large collective. As Hayek pointed out [hayek1945use], not all of the information is available to a central planner and so control is best delegated to the individual. The accumulation of resources within my formalism is implicit, rather than explicit. I have been concerned only with whether a system satisfies whatever arbitrary notion of correct I choose (i.e. does it achieve homeostatic and reproductive goals in whatever amount of time care about?), rather than how it evolves over time as a dynamical system. In other words, I have concerned myself only with whether a system arrives on time at the destination I want, not the journey it took to get there. Future research might formalise the accumulation of resources explicitly. One way to describe an evolving sequence of tasks over time could be something like a stack. To speculate about what my results imply for delegated systems like an economy, under less than ideal conditions resources may become concentrated in the hands of a few. The resulting system then delegates work, but not control. Those parts deprived of control may then not be able to construct a policy that aligns with the collective, and so they may splinter off and become ‘cancerous’. To prevent this, mechanisms can be put in place to preserve decentralisation without the need for intervention from outside the system (e.g. government intervening to break up a monopoly). Each system could be located along axes of redistribution and delegation, like an alternative political compass. Just enough redistribution of control and the system remains delegated. However, too much redistribution could have a similar effect, preventing parts from adapting to local conditions. An example of a mechanism that can effectively preserve delegation of control is adaptive rubber banding [mi2022]; a game mechanic (e.g. in Mario Kart) where success accrues penalties as well as resources, keeping the game competitive.
Another possibility is AI regulation. For example the UK government recently initiated “Safeguarded AI” [dalrymple2024b, dalrymple2024a], a project that aims to construct an AGI [goertzel2014] of level ASL-4 [anthropic2023] in a cyberphysical system, to safeguard against rogue AI. The idea of a “rogue AI” from within a cyberphysical system is analogous to cancer, adopting an “unsafe” policy at odds with that of the collective. AI regulations may over-constrain systems and trigger the very cancerous state they seek to prevent. To reiterate the earlier point about subtractive solutions [adams2021], any guardrails which are not strictly necessary must be eliminated to make room for those that are.
In summary, I have described systems in terms of delegation and abstraction. I proved that adaptability at higher levels of abstraction is constrained by adaptability at lower levels (The Law of the Stack). I have argued that one reason biological systems are more ‘intelligent’, meaning sample and energy efficient, is because they delegate adaptation down the stack. This Stack Theoretic perspective lays a foundation for future empirical research, showing how we might build more robust systems of any kind.
9 Electronic Supplementary Materials
9.1 Mathematical Results
The Law of the Stack
Theorem 9.1 (The Law of the Stack).
Assume the formal setting defined in the main text: finite vocabulary $v\subseteq P$, language $L_{v}=\{\,\ell\subseteq v:\bigcap\ell\neq\emptyset\,\}$, extensions $E_{\pi}=\{\,y\in L_{v}:\pi\subseteq y\,\}$, $v$-tasks $\gamma=\langle I_{\gamma},O_{\gamma}\rangle$, correct policies $\Pi_{\gamma}=\{\pi\in L_{v}:E_{I_{\gamma}}\cap E_{\pi}=O_{\gamma}\}$, and utility $\epsilon(\gamma)=\max_{\pi\in\Pi_{\gamma}}(|E_{\pi}|-|O_{\gamma}|)$. Let $\langle\lambda_{0},\dots,\lambda_{n}\rangle$ be an MLA with abstractor $\mathfrak{f}$ and state $\langle(\pi_{0},\dots,\pi_{n}),(v_{0},\dots,v_{n})\rangle$ such that $v_{i+1}=\mathfrak{f}(v_{i},\pi_{i})$ and $\pi_{i}\in\Pi_{\lambda_{i}(v_{i})}$ for all $i<n$.
Fix any $i<n$, and define $\gamma_{i+1}=\lambda_{i+1}(v_{i+1})$ where $v_{i+1}=\mathfrak{f}(v_{i},\pi_{i})$. Then
$\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\;\leq\;2^{|E_{\pi_{i}}|}.$
Equivalently,
$|E_{\pi_{i}}|\;\geq\;\log_{2}\!\big(\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\big).$
Thus, higher utility at layer i+1i\!+\!1 requires (at least) a weaker policy at layer $i$ in the sense of larger extension cardinality.
Proof.
We proceed by a sequence of elementary bounds that make the layer-to-layer dependency explicit.
Lemma 1 (Languages are bounded by the power set).
For any finite vocabulary $v$, the language $L_{v}$ satisfies $|L_{v}|\leq 2^{|v|}$.
Proof. By definition, $L_{v}\subseteq\mathcal{P}(v)$ because every $\ell\in L_{v}$ is a subset of $v$. Since $|\mathcal{P}(v)|=2^{|v|}$ for finite $v$, we have $|L_{v}|\leq 2^{|v|}$. $\square$
Lemma 2 (Extensions live inside the language).
If $\pi\in L_{v}$, then $E_{\pi}\subseteq L_{v}$, hence $|E_{\pi}|\leq|L_{v}|$.
Proof. By definition $E_{\pi}=\{y\in L_{v}:\pi\subseteq y\}$, so $E_{\pi}\subseteq L_{v}$. Taking cardinalities yields $|E_{\pi}|\leq|L_{v}|$. $\square$
Lemma 3 (Utility implies a lower bound on language size).
Let $\gamma$ be a $v$-task. Then
$\epsilon(\gamma)+|O_{\gamma}|\leq|L_{v}|.$
Proof. Because $v$ is finite, $L_{v}$ is finite, hence $\Pi_{\gamma}\subseteq L_{v}$ is finite, so the maximum defining $\epsilon(\gamma)$ is attained by some $\pi^{\star}\in\Pi_{\gamma}$. By definition,
$\epsilon(\gamma)=|E_{\pi^{\star}}|-|O_{\gamma}|\quad\Rightarrow\quad\epsilon(\gamma)+|O_{\gamma}|=|E_{\pi^{\star}}|.$
By Lemma 2, $|E_{\pi^{\star}}|\leq|L_{v}|$. Combining yields the claim. $\square$
Lemma 4 (The abstractor output vocabulary is an image of an extension).
Let $T:L_{v_{i}}\to P$ be the map $T(o)=\bigcap o$. Under the MLA definition,
$v_{i+1}=\mathfrak{f}(v_{i},\pi_{i})=\{\,T(o):o\in E_{\pi_{i}}\,\},$
so $v_{i+1}$ is the image of $E_{\pi_{i}}$ under $T$. Consequently,
$|v_{i+1}|\leq|E_{\pi_{i}}|.$
Proof. This is just re-writing the definition of $\mathfrak{f}$ in functional form, and then applying the general fact that for finite sets $A$ and any function $g$, $|g(A)|\leq|A|$. $\square$
Conclusion.
Apply Lemma 3 to $\gamma_{i+1}$ in vocabulary $v_{i+1}$ to obtain
$\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\leq|L_{v_{i+1}}|.$
Apply Lemma 1 to bound $|L_{v_{i+1}}|\leq 2^{|v_{i+1}|}$, giving
$\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\leq 2^{|v_{i+1}|}.$
Finally apply Lemma 4: $|v_{i+1}|\leq|E_{\pi_{i}}|$. Since $2^{x}$ is monotone increasing,
$2^{|v_{i+1}|}\leq 2^{|E_{\pi_{i}}|}.$
Combining the last two displays yields
$\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\leq 2^{|E_{\pi_{i}}|},$
and the logarithmic form follows by applying $\log_{2}$ (also monotone). ∎
Remark. In the thesis version of this result, the intuition is phrased as: larger vocabularies at layer i+1i{+}1 permit weaker (more general) policies, hence more adaptation at higher layers. The proof here makes the dependency explicit by bounding the achievable utility at layer i+1i{+}1 as a monotone function of the extension size $|E_{\pi_{i}}|$ at layer $i$, without requiring any additional monotonicity assumptions about $\epsilon(\lambda(\cdot))$ under set inclusion.
Hybrid case study formalisation (Pong organoid)
Let $C$ be a set of cells. For each cell $i\in C$, let $\Pi_{i}$ denote the set of viable policies for that cell under some interface/environmental conditions. The set of viable collective policies for the organoid is $\Pi_{C}=\bigcap_{i\in C}\Pi_{i}$. A conventional computer interface (electrodes, stimulus generator, game process) imposes boundary conditions $c$ that restrict which policies remain viable; write this restriction schematically as an operator $R_{c}$ such that $\Pi_{i}^{(c)}=R_{c}(\Pi_{i})$ and $\Pi_{C}^{(c)}=\bigcap_{i\in C}\Pi_{i}^{(c)}$. The design objective is to choose $c$ so that (i) $\Pi_{C}^{(c)}\neq\emptyset$ and (ii) the remaining collective policies implement a target output set $O_{\text{pong}}$ at the interface. In this view, building the hybrid is a constraint-design problem: impose the weakest (least exclusionary) boundary conditions that make $O_{\text{pong}}$ achievable while preserving non-empty intersection (collective identity).
9.2 Supplementary Notes and Extended Discussion
S1. Working definition of intelligence
There are many definitions of intelligence, and the merits of each are beyond the scope of this paper. Mine is based on [bennett2025thesis], which frames intelligence as the ability to complete a wide range of tasks, which is sample and energy efficiency in adaptation. It is similar to Wang’s definition of intelligence as adaptation with limited resources [wang2019], and Hutter’s earlier definition of intelligence as ‘the ability to satisfy goals in a wide range of environments’. Mainstream AI often operationalizes ‘intelligence’ as general-purpose task performance. By that standard, one may argue that current large language models are ‘more intelligent’ than many biological systems. This paper instead treats intelligence as the cost of adaptation (sample and energy efficiency), and the results should be read accordingly.
S2. Levels of abstraction beyond computing
The notion of levels of abstraction is more typically used in pure computer science settings, but can be applied to any system [bennett2025thesis]. A higher level of abstraction is the behaviour of the layers below. For example, an organ like a heart is a behaviour of cells, human is a behaviour of some organs, and language is a behaviour of humans.
S3. Weakness and w-maxing (extended intuition)
Weakness is formally defined in the appendix. Related work proved w-maxing to be optimal regardless of the abstraction layer [bennett2025e, bennett2023b, bennett2024a, bennett2025thesis]. In narrative terms, assume two policies $x$ and $y$ both constrain the system to examples of “correct” behaviour given past inputs, but differ in weakness. If $x$ is weaker than $y$, that means it commits to less, but achieves the same end. For example, if I assume the lawyer was paid at least $\mathdollar 5$, I am more weakly constrained than if I assume that lawyer was paid at least $\mathdollar 6$.
As an epistemological razor, this means “explanations should be no more specific than necessary” or alternatively “rules should constrain no more than necessary” (see Bennett’s Razor in [bennett2023b]). This is a bit like the principle of maximum entropy [jaynes1957a, jaynes1957b], difference being that w-maxing also accounts for the abstraction layer, which is important because abstraction can render complexity subjective [bennett2024b].
For a rigorous definition of learning and efficiency, see definitions in the online repository for the thesis [bennett2024appendices]. It includes proofs showing w-maxing is both necessary and sufficient to maximise the probability of ‘generalising’ from one task to another, and thus adaptability.
S4. Normativity and “strict necessity” in design
As Hume argued, a description of what ought to be cannot be derived from only descriptions of what is. “Strictly necessary” is a value judgement, or an implicit choice of what ought to be. In the thesis [bennett2025thesis] on which this is based this ought from which other descriptions of ought follow is derived from the state transitions of the environment. These transitions preserve some aspects of that environment, while destroying others. Here however I am more concerned with the design of systems which serve a human purpose. Strictly necessary is a matter of desired functionality.
S5. Conceptual motivation for the formal setting
Stack Theory frames the description of an objective reality in terms of choices of subjective lens (an abstraction layer) [bennett2025thesis]. The full justification for these definitions is out of scope and addressed at length elsewhere [bennett2024a, bennett2025thesis, bennett2025e]. However I will provide a brief overview. To make claims that hold regardless of the abstraction layer, we must rely only on what holds for every abstraction layer. From our subjective perspective we can’t know what is at the bottom of our stack (it may be infinite). I call this bottom layer the environment. Whatever the environment is, there must be change or difference for there to be anything at higher levels of abstraction. Hence there must be a set of states where each state is one point of difference [^35] Nothing is assumed about what these states might ontologically be or epistemologically contain.. Because these states must be mutually exclusive, this frames time as difference. Each possible sequence of states is a timeline or possible world. Truth is defined ontologically, relative to states. A set of states is effectively a set of truth conditions, so a possible fact is called a declarative program, and a program is true given a state $\phi$ if it contains that state. Every aspect of the environment is just a set of programs. Everything that exists is a set of facts, and each timeline preserves different sets of facts. A timeline is like an opinion; as time progresses some aspects of the environment are preserved and others are destroyed. Each timeline preserves different aspects, expressing a different “ought” which related work uses to explain the origins of goal directed behaviour [bennett2025thesis]. If the set of states is infinite then there are infinite possible worlds. Assume I have a robotic system. I get an abstraction layer for the system by choosing a finite aspect $\mathfrak{v}$ of programs called a vocabulary. Everything the abstraction layer can express is a subset of its vocabulary. For example if a system has an arm it can raise or lower, then there are programs in $\mathfrak{v}$ which correspond to actuation of the arm in some form, though it need not be neat symbolic representations like ’raise arm’ or ’lower arm’. Since states are mutually exclusive, $\mathfrak{v}$ gives us an embodied formal language. I do not mean like the words humans use to communicate. I mean a language of embodiment, where every statement in that language constrains rather than describes the system, like the aforementioned boundary conditions on slime mold [nakagaki2000] [^36] Using this frame cognition can be formalised in terms that align with both enactivist [thompson2007] and pancomputationalist [piccinini2021] perspectives. Hence, cognition within Stack Theory is also called Pancomputational Enactivism[bennett2024a].
S6. Extended related work on delegated software architectures and Active Inference/FEP
There are examples of systems that take something like the present approach, albeit within the confines of software abstraction layers.
Federated machine learning delegates control to some extent (e.g. shifting data ownership and parts of optimisation to edge devices), which can improve privacy and efficiency [nori2025federated]. Constructivist theories of artificial general intelligence emphasise multilayered self-organisation at the level of software while acknowledging at least some of the constraints imposed by hardware [thorisson2012, goertzel2014]. The Auto-catalytic Endogenous Reflective Architecture (AERA) is a goal-driven ampliative reasoning system [nivel2013] based on constructivist principles that relies on causal knowledge representation to achieve autonomy [thorisson2018cumlearncause], “growing” knowledge in a collective meta-model through cumulative continuous learning [thorisson2019cumlearning], and explicitly targeting adaptability [thorisson2014autonomousNatCom, thorisson2020, eberding2024, sheikhlar2024] [^37] See also http://www.openaera.org.. ActPC-Chem [goertzel2024] is a recent theory behind the distributed Hyperon architecture [goertzel2023] that was influenced by theoretical and methodological insights from AERA.
Active Inference/FEP in relation to MCL.
Active Inference (and the Free Energy Principle, FEP) provides a normative account of perception and action in which an agent maintains a generative model and selects actions to minimise variational free energy (and, by extension, expected free energy) [Friston2010, friston2013, parr2024]. The present paper addresses a different axis: it states a structural constraint on stacks of abstraction layers, namely that what can be achieved at layer i+1i{+}1 (utility $\epsilon(\gamma_{i+1})$) is bounded by the extension (weakness) of the realised policy at layer $i$ (Theorem 5.1). In this sense, MCL is about where adaptation is implemented (and how much latitude lower layers retain), rather than about the particular inferential objective used within a layer. An Active Inference agent implemented on a predominantly fixed computational substrate can still be bottlenecked by static lower layers; conversely, a delegated multiscale substrate can in principle host a range of learning rules, including Active Inference.
Potential complementarity and open questions.
Several lines of work suggest points of contact, without implying formal equivalence. ActPC-Chem frames delegation in explicitly Active Inference terms [goertzel2024, Friston2010, friston2013]. “Scale-free” active inference proposes that similar dynamical principles repeat across scales [friston2023b, friston2024a], which is conceptually close to the MCA picture and to the idea that policy alignment must be maintained across layers to preserve collective identity. Delegated [ororbia2022, ororbia2023, ororbia2023b] or federated [friston2024b] Active Inference approaches may therefore be promising candidates for building efficient hybrid systems at scale. Separately, work on “mortal computation” seeks to approach thermodynamic limits (e.g. the Landauer bound) by reducing memory–processor separation and associated bottlenecks [ororbia2024mortalcomputationfoundationbiomimetic, aurell2012, zou2021], which resonates with the present paper’s emphasis on efficiency and with the intuition that pushing adaptation downward can reduce overhead. However, the theories still differ in their primitives (e.g. tasks/weakness versus generative models/free energy), and it remains an open question whether weakness (w-maxing) can be recovered as, or justified by, a principle within the Active Inference apparatus. Establishing such a mapping is an important direction for future work.
S7. Mission command example (Nordbat 2)
A well documented example of the effectiveness of mission command in facilitating adaptation is Nordbat 2, a joint peacekeeping force deployed to Bosnia in the early 90s. Mission command allowed individual units to deviate from convention, adapt to local conditions and adopt a more ‘trigger-happy’ disposition that proved to be substantially more effective than the more rigid, rule based approach taken by the Dutch forces operating in the same region [ingesson2016].
S8. From analogy to evidence: candidate empirical and simulation tests
The main text uses analogies (mission command, economic delegation, and distributed swarms) as intuition-building thought experiments rather than as evidence. To turn the delegation hypothesis into a research programme, one can state testable predictions and compare systems under controlled variation of delegation depth.
1. Agent-based simulations with tunable delegation.
Construct an agent-based model of a collective where (i) individuals have local perception and local policy spaces, (ii) a collective objective is defined at a higher level, and (iii) a “delegation parameter” controls how much decision-making is centralised versus left to individuals (e.g. by constraining local policy spaces or routing decisions through a leader). Expose the collective to non-stationary environments (distribution shifts, component failures) and measure: sample efficiency (data required to regain acceptable performance), robustness (graceful degradation under failures), and identity preservation (whether the collective policy remains coherent rather than splintering into incompatible sub-collectives). Theorem 5.1 predicts that as delegation is pushed upward (lower layers become static or over-constrained), achievable utility at higher layers should become bounded by the reduced weakness of the lower-layer realised policies.
2. Comparative benchmarking across stack depth.
Compare learning systems that differ only in where adaptation is permitted. For example, keep a fixed high-level learning rule but vary whether lower layers are static (fixed hardware / fixed microcode) versus adaptive (e.g. reconfigurable hardware, adaptive firmware, or reconfigurable substrates). Measure sample and energy efficiency on tasks with distribution shift and compositional generalisation (e.g. ARC-style tasks). This isolates the “where in the stack” hypothesis from the choice of algorithm at the top.
3. Hybrid boundary-condition sweeps.
For hybrid agents such as simulated systems of human and AI agents, where the goal is to “align” the AI agents with goal goals, systematically vary the boundary conditions (stimulation regimes, feedback mappings, resource constraints) and quantify the trade-off between task performance and preservation of collective identity (non-empty intersections of viable policy sets, as in the organoid formalisation above). The design expectation is that weaker (less exclusionary) boundary conditions should preserve a larger viable policy space while still enabling the desired interface-level behaviour.
4. Interaction with Active Inference/FEP.
Implement Active Inference-style control at one or more layers, but run the same comparison as above: (i) active inference on a fixed substrate versus (ii) active inference on a substrate that permits deeper delegation (adaptation in lower layers). This helps disentangle improvements due to the inferential objective from improvements due to stack-level delegation.
These designs are intended as starting points rather than definitive protocols; the unifying aim is to replace analogy with controlled variation of delegation, and to measure adaptability in the efficiency-of-adaptation sense used throughout this paper.

### v6

@@
 Abstract
-Are biological self-organising systems more ‘intelligent’ than artificial intelligence (AI)?
+Are biological self-organising systems more “intelligent” than artificial intelligence (AI)?
 If so, why?
-I explore this through a mathematical lens which frames intelligence in terms of adaptability.
-I model systems as stacks of abstraction layers (Stack Theory) and compare them by how they delegate agentic control down their stacks, illustrating with examples of computational, biological, human military, governmental and economic systems.
-Contemporary AI rests on a static, human-engineered stack in which lower layers are static during deployment.
-Put provocatively, static stacks resemble inflexible bureaucracies, adapting only top-down.
-Biological stacks are more ‘intelligent’ because they delegate adaptation.
-Formally, I prove a theorem (The Law of the Stack) showing adaptability in higher layers requires sufficient adaptability in lower layers.
-Generalising bio-electric explanations of cancer as isolation from collective informational structures, I explore how cancer-like failures occur in non-biological systems when delegation is inadequate.
-This helps explain how to build more robust systems, by delegating control like the military doctrine of mission command.
-It also provides a design perspective on hybrid agents (e.g.
-organoids, systems involving both humans and AI): hybrid creation is a boundary-condition design problem in which human-imposed constraints prune low-level policy spaces to yield desired collective behaviour while preserving collective identity.
+I address this question using a mathematical framework that defines intelligence in terms of adaptability.
+Systems are modelled as stacks of abstraction layers (Stack Theory) and compared by how effectively they delegate agentic control down their stacks.
+I illustrate this using computational, biological, military, governmental, and economic systems.
+Contemporary AI typically relies on static, human-engineered stacks whose lower layers are fixed during deployment.
+Put provocatively, such systems resemble inflexible bureaucracies that adapt only top-down.
+Biological systems are more intelligent because they delegate adaptation.
+Formally, I prove a theorem (The Law of the Stack) showing that adaptability at higher layers is bottlenecked by adaptability at lower layers.
+I further show that, under standard viability assumptions, maximising adaptability is equivalent to minimising variational free energy, implying that delegation is necessary for free-energy minimisation.
+Generalising bioelectric accounts of cancer as isolation from collective informational structures, I analyse cancer-like failure modes in non-biological systems when delegation is inadequate.
+This yields design principles for building robust systems via delegated control, and reframes hybrid agents (e.g.
+organoids or human–AI systems) as weak boundary-condition design problems in which constraints shape low-level policy spaces while preserving collective identity.
 1 Introduction
-Biological systems self-organise; their parts and the environment interact to produce spatiotemporal order [ashby1947, seeley2002, rosas2018, bell2007, kelso1997, Friston2010, tognoli2014].
-They pursue homeostatic and reproductive goals.
-Some proactively learn [kukushkin2024], storing information about the past in order to anticipate future needs.
-To make accurate predictions in complex environments, a system must learn cause and effect [pearl2018, merker2005, barron2016].
-I consider a system’s ‘intelligence’ to be a measure of how efficiently it does this [bennett2024a, chollet2019, bennettmaruyama2022b, bennett2023c] [^1] Throughout, “intelligence” is used in an efficiency-of-adaptation sense (sample and energy efficiency).
-For a comparison with other common definitions, see Electronic Supplementary Materials (ESM) Note S1..
-In this view intelligence is a long-term adaptation hard-wired into the organism, that enables short-term adaptation during its lifetime [bennett2025thesis].
-Artificial intelligence (AI) is the study of adaptive machines, and machine learning (ML) is the study of machines that learn [russell2005ai, bishop2006, sutton2018, goertzel2014].
-Contemporary approaches to machine learning are often inefficient [bennettmaruyama2022a].
-They can require vast quantities of data and energy to learn tasks a human can pick up after one or two examples.
-Even organoids [^2] By organoid I mean a collective of cells grown in a lab, placed in a computational interface and life support system, and then ‘trained’ to communicate with conventional computing technology.
-appear more ‘intelligent’ in this sense [khajehnejad2024].
-For example, an organoid outperformed conventional reinforcement learning systems at the game Pong, adapting with greater sample efficiency.
-That organoid was a highly tuned hybrid system whose behaviour was shaped by human-imposed boundary conditions.
-This is similar to how an acellular slime mold can be made to ‘solve’ shortest-path problems [nakagaki2000].
-These sorts of results are highly dependent upon the human imposed boundary conditions, but they raise an interesting question.
-Are biological self-organising systems more intelligent than artificial intelligence?
-If so, why?
-Some say the key difference between biological systems and contemporary artificial intelligence is that the former are active inference systems, while the latter are typically not [parr2024, Friston2010, friston2013].
-It is true that active inference can improve efficiency [ororbia2022, ororbia2023, ororbia2023b].
-Active inference has been applied in the context of the ARC-AGI benchmark [chollet2019], which to some extent tests sample efficiency.
-Active inference systems performed well (for example the 2024 ARC-AGI submission by Jack Cole), but still underperform humans.
-Hence there may be more to the picture [^3] From the perspective of this paper, Active Inference/FEP is primarily a theory of how an agent updates beliefs and selects actions (e.g.
-via variational free energy minimisation), whereas this paper is more about where and at what level of abstraction adaptation is realised in a system’s stack.
-Hence I treat the approaches as potentially complementary.
-An extended comparison and related work is provided in ESM Note S6..
-Others seem to suggest biological systems are not more intelligent [^4] In the sense of sample and energy efficiency of adaptation., and the difference is primarily one of inductive bias [chollet2019].
-This inductive bias is hard-coded foreknowledge of the environment, which makes it easier to learn and adapt to that environment.
-To whatever extent this is true, it substitutes the question ‘what is intelligence?’ for ‘what is the inductive bias?’.
-To reiterate, I want to know why biological self-organisation seems more adaptive [^5] Sample and energy efficient.
-than contemporary AI.
-Is there something else biological systems do that allows them to be more ‘intelligent’?
-Could there be more to the picture than active inference and an inductive bias?
-I suggest there is.
-Computers are organised into layers of abstraction, each layer being the behaviour of the layer below.
-For example, software runs on hardware.
-If software is making decisions, then it controls the hardware top-down.
-In dominant contemporary AI deployments, adaptation is concentrated at a comparatively high level of abstraction (e.g.
-model parameters and software-level state), while the lower computational stack (instruction set, microarchitecture, hardware) is treated as fixed during deployment.
-[^6] This is not a claim about a fundamental limitation of computational systems: reconfigurable or self-modifying substrates (e.g.
-FPGAs, adaptive firmware, agentic materials, or bio-hybrid substrates) can in principle shift adaptation downward.
-The claim here is about the dominant architectures and workflows in contemporary AI.
-In contrast, biological self-organisation uses the multiscale competency architecture (MCA).
-Adaptation takes place at every scale, from cells to organs to organisms [fields2020].
-It is ‘scale-free’.
-The same dynamics are observed at different scales.
-Control is delegated, in the sense that individual cells pursue goals [^7] As opposed to in a computer where the parts are usually designed to passively await instruction..
-Cells are a “soft” material that has agency [man2019, ball2023].
-A cell has a perceptual field.
-A cell can co-operate with other cells in this perceptual field.
-This allows them to form multicellular networks that have a larger perceptual field than any individual cell, to maintain coherent higher level “identity” like that of an organism [levin2024].
-Each cell has its own goals and is intelligent in the sense of being able to reach those goals by various means [james1890, ciaunica2023].
-However each cell can also be part of a collective of cells that in turn pursues its own collective goals.
-Smaller collectives can form larger collectives, and the pattern repeats at larger scales.
-Collectives are nested [ciaunica2021, ciaunica2023b].
-Each “thinks” in the enactive sense, interacting with its environment [thompson2007, varela2016, rolla2021].
-Biological collectives can exhibit high level goal directed behaviour far beyond the capability of the parts of which they’re formed [reid2011].
-The somewhat provocative claim this paper will make is that biological self-organising systems are more intelligent [^8] In the specific efficiency-based sense defined above.
-than contemporary, predominantly fixed-stack implementations of AI [^9] This is not a claim about all possible future AI.
-One aim of the paper is to identify design principles for building more adaptable hybrid and cyberphysical systems by shifting where adaptation occurs in the stack..
-This is because biological self-organisation delegates adaptation to lower levels of abstraction [^10] For readers who want the generalisation of “levels of abstraction” beyond computing, see ESM Supplementary Note S2..
-Formally, previous work showed adaptability in embodied systems can be measured by the ‘weakness’ of the constraints they embody [bennett2025d, bennett2024a].
-Building on this, I prove a theorem (The Law of the Stack) that the ability to embody weak constraints at higher levels of abstraction depends on the ability to embody weak constraints at lower levels.
-This result shows how we can build more robust and efficient hybrid and cyberphysical systems, laying a foundation for future empirical research.
-Its applications extend beyond artificial intelligence, for example to human organisational and economic systems.
+Organoids are living systems whose behaviour is sculpted by engineered constraints.
+Human brain cells embedded in a computational interface have been trained to play Pong, achieving remarkable sample efficiency compared to a contemporary reinforcement learning agent [53].
+Similarly, Physarum slime mould is normally an indiscriminate expander.
+However it reliably computes shortest paths when placed inside a maze [62].
+These behaviours are neither programmed nor accidental.
+They arise because biological agency is coupled to carefully designed boundary conditions.
+Such systems are hybrid agents, jointly produced by bottom-up cellular dynamics and top-down environmental structure [82].
+Biological systems self-organise across scales.
+Their parts and environment interact to produce spatiotemporal order [4, 78, 76, 9, 51, 39, 92].
+Many systems proactively learn, storing information about the past to anticipate future needs [54].
+To adapt in complex environments, a system must infer cause and effect [72, 60, 7].
+Throughout, I use ‘intelligence’ in the specific sense of the sample and energy efficiency of adaptation [16, 26, 11, 14] [^1] Mainstream AI typically treats intelligence as performance across tasks.
+My conclusions concern efficiency of acquiring said performance.
+Accordingly, claims about “AI” should be read as applying to dominant contemporary ML training and deployment stacks, rather than all possible systems.
+See ESM Note S1..
+Two common explanations for biological efficiency are active inference under the Free Energy Principle (FEP) [39, 40, 71] and powerful inductive biases [26].
+These address, respectively, how agents update beliefs and act, and what structure is pre-specified.
+This paper isolates a complementary structural question, namely where in a system is adaptation permitted?
+Computers are organised into layers of abstraction.
+In dominant AI deployments, adaptation is concentrated at high software layers (model parameters and software-level state) while lower layers (instruction set, microarchitecture, hardware) are treated as fixed during deployment.
+[^2] Lower layers can carry mutable state (data, caches, firmware), and self-modifying substrates exist; the claim here is about the prevailing engineering pattern in mainstream ML systems.
+In contrast, biological multiscale competency architectures (MCA) delegate goal-directed control across scales, from cells to organs to organisms [34, 59].
+I model this difference with Stack Theory, formalising systems as stacks of abstraction layers.
+Delegation then measures how far adaptive control reaches down the stack.
+The somewhat provocative claim this paper will make is that biological self-organising systems are more intelligent (in the efficiency-of-adaptation sense) than contemporary, predominantly fixed-stack implementations of AI.
+This is not a claim about all possible future AI.
+One aim of the paper is to identify design principles for building more adaptable hybrid and cyberphysical systems by shifting where adaptation occurs in the stack.
+This is because biological self-organisation delegates adaptation to lower levels of abstraction (for the generalisation of “levels of abstraction” beyond computing, see ESM Supplementary Note S2).
+Formally, I measure adaptability by the weakness of the constraints a system embodies [20, 16].
+I then prove The Law of the Stack (Theorem 5.1), which is that utility (adaptability) achievable one layer up is bounded by the weakness of the realised policy one layer down.
+Finally, I link this to the FEP.
+Under a simple viability prior, maximising weakness is equivalent to minimising variational free energy (see ESM “Viability, lived history, and variational free energy”), which makes delegation a prerequisite for sustained free energy minimisation across scales.
+For clarity, the core contributions can be read as four claims:
+Claim 1.1 (Delegation constrains adaptability).
+In a multilayer architecture, increasing adaptability at an interface layer requires sufficiently weak policies at lower layers (formalised by Theorem 5.1).
+Claim 1.2 (Delegation is necessary for free energy minimisation).
+Under a simple viability prior, variational free energy minimisation reduces to keeping options open (maximising entropy over viable futures).
+The Law of the Stack then implies that abstract-layer free energy cannot be reduced without sufficient lower-layer slack, making delegation a structural prerequisite for sustained free energy minimisation.
+Claim 1.3 (Insufficient delegation can induce cancer-analogue fragmentation).
+Model a hybrid at some vocabulary $v$ as a distributed system with part-tasks $\alpha_{1},\dots,\alpha_{m}$ and a collective task $\omega$ whose collective-policy set is
+$\Pi_{\omega}\;:=\;\bigcap_{j=1}^{m}\Pi_{\alpha_{j}}.$
+(Thus the collective has an identity at $v$ iff $\Pi_{\omega}\neq\emptyset$.) Model top-down boundary conditions / insufficient delegation as additional constraints that restrict each part’s feasible policies to a subset $\widehat{\Pi}_{\alpha_{j}}\subseteq\Pi_{\alpha_{j}}$, inducing
+$\widehat{\Pi}_{\omega}\;:=\;\bigcap_{j=1}^{m}\widehat{\Pi}_{\alpha_{j}}.$
+If these restrictions render the system over-constrained at $v$ (i.e.
+$\widehat{\Pi}_{\omega}=\emptyset$), then no common policy exists at that level of abstraction.
+If at least one part remains locally viable (i.e.
+$\widehat{\Pi}_{\alpha_{j}}\neq\emptyset$ for some $j$), then any continued function must proceed via a proper subset of parts whose restricted policy sets still admit a non-empty intersection.
+I refer to this formal situation as “cancer-like” purely by analogy.
+Claim 1.4 (Hybrid creation as maximally weak boundary-condition design).
+Hybrid agents (e.g.
+organoids, soft robots, human–AI organisations) are best treated as boundary-condition design problems, which I formalise using Stack Theory.
+One must engineer constraints that prune low-level policy spaces just enough to realise target collective behaviour while preserving a non-empty intersection of component policy sets (collective identity).
+Formalised this way, the optimal choice of boundary conditions are the weakest sufficient constraints to achieve the desired hybrid functionality.
+The remainder of the paper develops the intuition (Section 2), states the formal framework and theorem, and applies the result to hybrid agency and failure modes.
+The discussion of potential applications spans computational, biological, human governmental and economic systems.
 2 Abstraction, Distribution and Delegation
 Conventional computing systems are organised into stacks of abstraction layers.
-More generally, I will characterise systems along three interacting dimensions: abstraction (how layers relate), distribution (how work is spread within a layer), and delegation (the level of abstraction at which adaptation or control resides).
+More generally, I characterise systems along three interacting dimensions.
+Abstraction describes how layers relate.
+Distribution describes how work is spread within a layer.
+Delegation is the level of abstraction at which adaptation or control resides.
 How efficiently a system adapts depends on the interaction between these dimensions.
-In proposing Stack Theory [^11] Initially proposed in my PhD thesis [bennett2025thesis].
-[bennett2025thesis], I merely extend this idea to systems in general.
-I’ll begin illustrating what I mean by abstraction, delegation and distribution; dimensions along which one may plot any system.
-For intuition, I will give examples of limit cases for computational, biological, human organisational and economic systems.
+In proposing Stack Theory [19], I extend this idea beyond computing.
+I illustrate abstraction, distribution, and delegation with limiting-case examples from computing, biology, and human organisations.
 2.1 Abstraction
-In a computer, a predictive model typically relies on a machine learning library, which runs on CUDA [^12] This is the software library that allows for the large scale utilisation of GPUs (a type of hardware originally developed for games), for machine learning and artificial intelligence., which runs on a GPU [^13] A graphical processing unit (GPU) is a type of hardware that allows for massively parallel vector processing.
-It was developed for the purpose of graphics processing in videogames, and later repurposed for machine learning..
+In a computer, a predictive model typically relies on a machine learning library, which runs on CUDA (a GPU programming platform), which runs on a graphics processing unit (GPU).
 GPUGPUCUDACUDAlibrarylibrarymodelmodel
-These are all examples of abstraction layers.
-A programming language is a sort of abstraction layer.
-For example Python is an abstraction layer over C, in the sense that it is interpreted by a program written in the less abstract C programming language.
-C in turn depends on a program written in Assembly language, which runs on machine code which is interpreted by hardware.
-Each layer of language is written in the language of the layer below.
-This structure allows programmers to work more efficiently, within the confines of a standardised and simplified representation of the system.
-However, these abstraction layers are incomplete, simplified pictures of the system.
-They “abstract away” unnecessary details.
-Most importantly, they are embodied in the layer below.
-Each language is a subset of the behaviour that the layer below can exhibit.
-For example, a subset of the possible behaviour a given computer can exhibit would correspond to running Python code, and a broader subset of that computer’s behaviour would necessarily correspond to running C code.
-Hence each layer in this stack is a subset of the behaviours of the layers below [^14] A higher-level abstraction is not uniquely determined by its lower-level implementation: the same abstract interface can be realised by multiple implementations.
-The claim here is that higher-level behaviour must be realised by lower-level behaviour, and that different implementations can differ in resource cost (time, energy, information), which matters for adaptability..
-This reductionist framing does not deny that useful abstractions can be robust to lower-level variation; invariants can be treated as higher-level task constraints.
-Even the physical hardware itself can be framed as an abstraction layer over the environment, being a subset of the behaviour of that environment in which the hardware exists [bennett2025thesis].
-Think of a stack of abstraction layers as a directed graph.
-Highly abstract behaviours are a consequence of ‘weaker’, more generalisable goal directed behaviour at lower levels of abstraction.
-Framed this way, everything is a stack of abstraction layers [^15] This is not an analogical or metaphorical application of computational frames to others systems, as in Bratton’s re-framing of geopolitics [bratton2016].
-Rather, this is a more general claim about how all systems work..
-For example a collective of cells can enact an organ, and a system of organs can enact an organism [levin2024] which enacts even more abstract behaviours like language or tools.
-Each of these is an ‘abstraction layer’ in a biological stack [bennett2024a].
+Abstraction layers are interfaces that hide lower-level details.
+In computation, Python is interpreted by C, which runs on assembly and machine code on hardware.
+Each higher-layer behaviour must be realised by some lower-layer behaviour, but the same abstract interface can be realised by multiple implementations.
+This many-to-one mapping is where time, energy and information costs hide, so the abstraction layer through which an agent senses and acts constrains its adaptability.
+In Stack Theory, useful invariants are treated as higher-level task constraints even when micro-implementations vary.
+Even hardware itself can be framed as an abstraction layer over the local environment [19].
+Think of a stack as a directed graph of constraints.
+Higher-level behaviours are implemented by and limited by lower-level behaviours.
+In biology, cells enact organs, organs enact organisms, and organisms enact still more abstract behaviours such as communication and tool use [59, 16].
 cellscellsorgansorgansorganismorganism
 2.2 Distribution
-In the case of biological systems lower levels of abstraction are formed of smaller or simpler parts [delgado1997], which is to say the work is distributed among parts at a level of abstraction.
-Because of this, looking at a lower level of abstraction is like ‘zooming in’ to look at a smaller scale.
-Conversely, rather than looking at low levels of abstraction like cells I can shift the level of abstraction to higher in the stack, and look at individual humans and how they form collectives at higher levels of abstraction.
-For example a population of soldiers can be a squad, and if squads co-operate they can be a platoon.
-Each layer above is the behaviour of a layer below, and the parts get smaller when I go down.
+Distribution refers to how work is spread within a layer.
+In many biological systems, lower levels of abstraction are literally smaller parts [32], so “going down” is like zooming in.
+Cells form organs and organisms form organisms.
+For human organisations, soldiers form squads and squads form platoons.
 soldierssoldierssquadssquadsplatoonplatoon
-In each layer work is distributed among members of the population.
-If level of abstraction is one dimension along which a system may be plotted, then distribution of work is another.
-In a computer I can distribute work by splitting it up into sub-tasks and allocating it to many computers.
-However, unlike biological systems a lower level of abstraction in a computer is not necessarily made up of smaller parts that form a collective.
-Looking at a lower level of abstraction is not like ‘zooming in’.
-I can run software on a solitary computer, or a cluster of computers.
-For example, Python code is run by an interpreter, which is a C program that in turn is built on Assembly, which is built on machine code that has a fixed interpretation physically hard-wired in the instruction set architecture of the CPU.
-The scale of the smallest part at each layer of abstraction is the same.
-Think of one Python program running on one interpreter, on one computer.
-This illustrates how level of abstraction can vary without affecting the scale or distribution.
+Distribution is distinct from abstraction, in that the same abstraction layer can be realised by one component or many.
+In computation, lower abstraction does not necessarily mean smaller spatial parts.
+A single Python program is interpreted by a C program, which relies on assembly and machine code executed by the same physical CPU.
+Thus one can change abstraction without changing the scale or distribution of underlying components.
 2.3 Delegated Control
-The aforementioned stack on which Python relies is not a collective.
-It is not made up of parts capable of functioning independently.
-They do not have agency.
-This means control of the system resides at another level of abstraction.
-If I am programming a computer, then control resides at the level of abstraction at which I am programming.
-If I am programming in Python, then adaptation occurs at the Python level of abstraction [^16] This is a statement about where learning/adaptation is typically implemented in conventional software stacks.
-Lower layers can of course carry mutable state (data, caches, firmware state), but are usually not designed to be autonomously goal-directed in the way MCA components are..
-Running a program in a distributed system does not always affect this.
-Distributing work is not the same thing as delegating control or agency.
-Work can be distributed among many GPUs, but if I remove a GPU from the computer it will not then pursue goals independently.
-It will sit, inert until once again integrated into a larger system.
-It requires no incentives.
-Control is imposed top-down, from higher to lower levels of abstraction.
-The parts at lower levels have no agency.
-Conversely in a biological system each cell is independently capable of goal directed behaviour [ball2023], assuming it remains alive.
-Control in a complex biological system begins bottom-up, before top-down.
-Control then propagates in both directions [^17] In the sense that members of a collective are constrained by the collective, for example members of a human society are incentivised to obey laws.
+Delegation is the level of abstraction at which adaptive control resides.
+In a typical computing stack, lower layers (instruction set, microarchitecture, hardware) implement fixed semantics rather than autonomous goal-directed behaviour.
+Adaptation therefore occurs where policy can change.
+This is usually at the software layer we program, even though lower layers can carry mutable state (data, caches, firmware).
+Distribution is not delegation.
+A workload can be spread across many GPUs without granting any component independent control.
+Remove a GPU and it does not continue pursuing the system’s objectives, or any other for that matter.
+Biological collectives differ in that cells and tissues are autonomously goal-directed at their own scale [6].
+Higher-level coherence arises when the viable policy sets of parts intersect, producing a collective policy (collective identity).
+Control can propagate bottom-up and top-down by which I mean collectives constrain their parts (e.g.
+via signals or incentives), but parts also shape collective behaviour.
+For illustrative purposes, consider how members of a collective are constrained by the collective, for example members of a human society are incentivised to obey laws.
 This is a form of top-down control.
-Bottom-up then becomes just whatever the individual humans do within those constraints, even though the system as a whole might have emerged entirely bottom-up from anarchy to begin with..
+Bottom-up then becomes just whatever the individual humans do within those constraints, even though the system as a whole might have emerged entirely bottom-up from anarchy to begin with.
 Each part of a collective acts on its own, and its actions imply a ‘policy’.
@@
 Or I could describe the human as the collective policy of cells, and see the organs as a sort of middle management layer.
-A good example to understand the limit cases for delegation is an economy.
-In an absolutely centrally planned economy, control is concentrated at the highest level of abstraction.
+Economies illustrate useful limit cases to understand delegation.
 One might think of the people in the economy as parts at lower level of abstraction.
-They do work but have no control.
-Those people could adapt, but central planner’s edicts constrain them.
-The opposite extreme would be to have control delegated entirely to the lowest level of abstraction, in anarchy unconstrained by any laws.
-The level of abstraction is often related to the granularity at which the system is represented.
-For example if the level of abstraction at which I view an economic system is humans, then I don’t worry about how work is distributed or control is delegated within those humans.
-I just concern myself with what systems can be formed of humans, like economies or armies (see fig 1).
-Of course, there are limits on what sort of physical forms can store or process information in service of a goal [sole2024], and this may limit how much I can distribute work or delegate control.
-Delegating to a lower level of abstraction often requires distributing work among parts at smaller scales [^18] For example, delegating control to the level of individual humans would mean those humans control their own work potential, but not that of others, or else we’d just be delegating to collectives within the human population..
-However as the static abstraction layers in a computer illustrate, a lower level of abstraction is not always more granular, and thus work is not always more distributed (see fig 2).
+Central planning concentrates control at the top, while extreme decentralisation delegates control to the bottom.
+In the central planning limit case, the population does work but has no control.
+In the limit those people could adapt given local information unavailable to the central planner [44], but central planner’s edicts constrain them.
+In the opposite extreme there is no coordination.
+Real systems occupy a continuum.
+Of course, there are limits on what sort of physical forms can store or process information in service of a goal [81], and this may limit how much I can distribute work or delegate control.
+Delegating control often coincides with shifting to finer granularity and greater distribution.
+However, as computers illustrate, abstraction can also change without changing the spatial scale of components (Figures 1 and 2).
 Figure 1: Control or agency and thus adaptation in a given system can take place at a particular level of abstraction.
@@
 If I program in C it doesn’t matter whether I program one computer or fifty.
-Either way my control is at the same level of abstraction (figure from [bennett2025thesis]).
+Either way my control is at the same level of abstraction (figure from [19]).
 Figure 2: Work, resources and so on can be distributed across a system.
 Distribution is distinct from delegation of control.
-For example, both food stamps and universal basic income might distribute the same amount of resources to the same people within an economic system, but the former more tightly constrains the recipients in how they use those resources (figure from [bennett2025thesis]).
+For example, both food stamps and universal basic income might distribute the same amount of resources to the same people within an economic system, but the former more tightly constrains the recipients in how they use those resources (figure from [19]).
 2.4 Reasons to Suspect Delegation Facilitates Adaptation
-I will now give a brief explanation of why I believe delegation of control to lower levels of abstraction is relevant to adaptation.
-Throughout this subsection I use examples (e.g.
-distributed drone swarms, mission command, and economies) as intuition-building thought experiments.
-They are not intended as comparative empirical evidence for the formal claim, which is established later in Theorem 5.1.
-All else being equal the lower the level of abstraction at which the system can be controlled, the more it can be specialised.
-For example a program written in C can be far more efficient than a program written in Python [oden2020].
-This is because the Python program has to be interpreted by the Python interpreter, which is a program written in C.
-By programming in C I simply cut out an abstraction layer.
-This gives the programmer more leeway to optimise the system, or make mistakes.
-Likewise another way to make computers more efficient is to use a FPGA to write a program at the hardware level, instead of compiling that program into machine code and running it on a one-size-fits-all instruction set architecture like x86 [nguyen2020].
-With each successive abstraction layer introduced, one is further removed from the ground truth to which one must adapt.
-For example software is nothing more than a state of hardware [^19] That software is a state of hardware has been addressed at length elsewhere [bennett2024a, bennett2025thesis], beyond the scope of this paper..
-It is a label I ascribe to the behaviour of hardware.
-It is a useful simplification that lets us consider what I want hardware to do in the abstract, but working at such a high level of abstraction can cause problems.
-Proposals for software ‘intelligence’ have been undermined by the fact that any claims regarding the behaviour of software are contingent upon hardware.
-The software doesn’t interact with the environment, the hardware does.
-This is because the ability to adapt is related to perceived complexity, which is entirely determined by the abstraction layer through which an agent interacts with its environment [bennett2024b].
-Consequently an ‘objective’ measure of artificial intelligence [leike2015, bennett2024a, bennett2025b] must be ‘enactive’ [thompson2007] in that it hinges not just software but hardware and the local environment as well.
-To understand the difference between biological systems and artificial intelligence, I must account for the level of abstraction at which an adaptive process takes place.
-In distributed computational systems, like a drone swarm or network, delegating control can facilitate what is called ‘graceful degradation’ [kuniavsky2010].
+Theorem 5.1 provides the formal claim.
+Here I briefly motivate why delegation should matter.
+Because perceived complexity is defined by the abstraction layer through which a system interacts with its environment [17], measures of intelligence that ignore embodiment are incomplete; an “objective” measure must be enactive [86].
+First, adaptation closer to the substrate can exploit more degrees of freedom.
+Cutting out an abstraction layer can yield large efficiency gains (C vs Python) [65]; reconfigurable substrates such as FPGAs push adaptation further down the stack [63].
+Second, delegation can improve robustness via graceful degradation, avoiding single points of failure [55].
 For example, to control a drone swarm I could single out one drone to be the ‘leader’, and have it run a program that co-ordinates the other drones.
@@
 If the leader drone were damaged, the swarm would de-cohere.
-It is a single point of failure.
-In contrast if I can get the swarm to co-ordinate through entirely decentralised control, delegating to the level of individual drones, the swarm will remain coherent regardless of how many drones are disabled [^20] This sort of decentralised swarm control can also be observed in nature, such as with flocks of starlings which fly in complex formations [cavagna2010]..
+In contrast if I can get the swarm to co-ordinate through entirely decentralised control, individual drones negotiating with those around them, then the swarm can remain coherent regardless of how many drones are disabled.
+This sort of decentralised swarm control can also be observed in nature, such as with flocks of starlings which fly in complex formations [25].
 In other words, the swarm ‘adapts’ to adverse conditions.
-The advantages of delegating control are visible in human organisations as well.
-In military doctrine, “mission command” delegates decision-making to units closest to the situation, granting latitude to deviate from detailed orders in service of higher-level intent [Shamir01102010].
-Conceptually, each layer of management is an abstraction layer: when control is delegated, feedback loops shorten because fewer layers sit between the controlling entity and local conditions.
-This supports the broader hypothesis that delegation is a key ingredient of biological self-organisation that is largely absent from mainstream contemporary artificial intelligence.
-See ESM Note S7 for a concrete historical illustration (Nordbat 2).
+Third, in human organisations delegation shortens feedback loops.
+In military doctrine, “mission command” delegates decision-making to units closest to the situation, granting latitude to deviate from detailed orders in service of higher-level intent [79].
+This trades central control for responsiveness.
+See ESM Note S7 for a concrete historical illustration (Nordbat 2 [47, 48]).
+These are all intuition-building thought experiments rather than empirical proof, but they clarify the mechanism the theorem captures.
+To adapt at high levels, a system must retain enough slack at lower levels.
 3 What is Adaptation?
@@
 causeeffect
-To make accurate inferences a system must be able to tell the difference between what it has caused, and what it has merely observed [merker2005, pearl2018, dawid2002].
+To make accurate inferences a system must be able to tell the difference between what it has caused, and what it has merely observed [60, 72, 31].
 It is often difficult to determine cause.
@@
 not the causeeffectcause
-By actively intervening in the environment, a system can experiment to establish cause [^21] Assuming such experimentation is enough to mitigate ambiguity and eliminate unknowns, and that a heuristic such as weakness is applied to ensure the system converges on variable definitions which represent cause accurately [bennett2023c]..
+By actively intervening in the environment, a system can experiment to establish cause [^3] Assuming such experimentation is enough to mitigate ambiguity and eliminate unknowns, and that a heuristic such as weakness is applied to ensure the system converges on variable definitions which represent cause accurately [14]..
 Then, it can construct a causal graph.
@@
 One option would be to assume the nodes (variables, actions and other abstract objects and properties), and then learn the causal relations between them.
-This might work well in theory [richens2024robust], but it requires foreknowledge of objects and properties between which causal relations can be established.
-The alternative is to assume the causal relations (the edges in the graph), and then learn the variables.
-This needs only attraction and repulsion from physical states [bennett2023c, bennettwelshciaunica2024, bennett2025thesis], otherwise known as valence.
+This might work well in theory [74], but it requires foreknowledge of objects and properties between which causal relations can be established.
+The alternative is to assume the causal relations (the edges in the graph), and then learn the variables [14, 19, 18].
+This needs only attraction and repulsion from physical states [12, 19], otherwise known as valence.
 I take this latter approach.
-Learning nodes to fit a graph is like figuring out which computer program [^22] Computation in this paper is about constraints, like how slime mold can be made to “compute” by imposing the right constraints on them[nakagaki2000].
-I’m not assuming Turing machines.
-was used to generate a sequence.
+While causal discovery of this sort remains difficult to engineer in open-ended settings (e.g.
+unobserved confounders and combinatorial hypothesis spaces), I treat this as a modelling stance and focus on the stack-level delegation constraint formalised in Theorem 5.1.
+Learning nodes to fit a graph is like figuring out which computer program was used to generate a sequence.
 Assume I have fed a set of inputs into a computer.
 For each input the computer responded with a particular output.
-The computer runs a ‘policy’ program which maps inputs to outputs [^23] I call such programs policies, but this should not be confused with the more specific definition policy in reinforcement learning [sutton2018]..
+The computer runs a ‘policy’ program which maps inputs to outputs [^4] I call such programs policies, but this should not be confused with the more specific definition policy in reinforcement learning [84]..
 That policy is the cause of the input-output history observed so far.
@@
 If I wish to know which policy actually caused the outputs, I need a way to whittle down the remaining possibilities.
-There are two ways to compare policies that look the same with respect to historical data, but differ with respect to possible future data [bennett2024b].
+There are two ways to compare policies that look the same with respect to historical data, but differ with respect to possible future data [17].
 One is to compare their form, and choose policies which are simpler in the sense of being more compressible.
-However, in an interactive setting simplicity is an unreliable indicator of plausibility [leike2015, bennett2024b].
+However, in an interactive setting simplicity is an unreliable indicator of plausibility [56, 17].
 The alternative is to compare policies by their function, in terms of how they behave given inputs for which I do not have historical data.
-Adaptive systems must satisfy functional constraints [sole2022].
+Adaptive systems must satisfy functional constraints [83].
 I can compare systems by these constraints.
@@
 The weaker that constraint is, the more variety of future behaviour it permits.
-Put another way, if two policies satisfy a goal, then the weaker policy satisfies it in more possible worlds [^24] To the extent that such possible worlds can be differentiated from one another within the subjective confines of an abstraction layer..
-Previous research showed that if two policies are identical with respect to historical data, then the policy most likely to have generated historical data is the weaker [^25] Formal definitions and proofs for weakness and w-maxing are in [bennett2024appendices, bennett2025thesis]; an extended intuition is in ESM Supplementary Note S3.
-of the two [bennett2023c, bennett2025d, bennett2025e].
-The policy most likely to have caused behaviour is the weakest [bennett2023c], and in experiments a system which preferred weaker policies outperformed one that preferred the simpler policies by $110-500\%$ [bennett2023b].
-This has been used to establish an upper bound on embodied ‘intelligence’ [bennett2024a].
-I will use the term ‘w-maxing’ to mean a system optimises for weak policies, and ‘simp-maxing’ to mean a system optimises for simpler policies [^26] Basically minimum description length policies [rissanen1978]., to conform to the terminology used in previous work [bennett2025b, bennett2025thesis] [^27] For formal definitions of learning and efficiency used in this framework, see [bennett2024appendices] and the thesis [bennett2025thesis]..
+Put another way, if two policies satisfy a goal, then the weaker policy satisfies it in more possible worlds [^5] To the extent that such possible worlds can be differentiated from one another within the subjective constraint of an abstraction layer..
+Previous research showed that if two policies are identical with respect to historical data, then the policy most likely to have generated historical data is the weaker [21, 19] (see ESM Supplementary Note S3 for formal details.) of the two [14, 20, 18].
+The policy most likely to have caused behaviour is the weakest [14], and in experiments a system which preferred weaker policies outperformed one that preferred the simpler policies by $110-500\%$ [15].
+This has been used to establish an upper bound on embodied ‘intelligence’ [16].
+I will use the term ‘w-maxing’ to mean a system optimises for weak policies, and ‘simp-maxing’ to mean a system optimises for simpler (minimum description length [75]) policies, to conform to the terminology used in previous work [22, 19, 21].
 Put another way, if two hypotheses appear equally good with respect to past data, the w-maxing system will use policy weakness as a heuristic to decide which policy to use, whereas a simp-maxing system will use simplicity.
+Weakness is not the same thing as simplicity.
+Simplicity is a property of a policy’s description in a chosen encoding.
+Weakness is a property of the policy’s remaining completion space inside the embodied language $L_{\mathfrak{v}}$.
+Two correct policies can therefore have very different weakness even when their descriptions are equally short, or when one is syntactically longer.
+Weakness is also not mere truth set size.
+It counts how many distinct further commitments remain available to the agent given what it can express and implement at that abstraction layer.
+In this sense, w-maxing implements a least commitment heuristic stated at the level of function, rather than at the level of syntax.
+Since the completion structure depends on the vocabulary, weakness naturally incorporates embodiment.
+Changing sensors, effectors, or internal degrees of freedom changes $\mathfrak{v}$ and changes which refinements count as distinct.
+In the uniform case where all remaining completions are treated as equally plausible, maximising weakness coincides with a maximum entropy preference over viable futures.
 An abstraction layer can be understood as a border between form and function.
-For example a program takes a form when it is written in a language, but when it is run on a computer it has a particular function or effect upon the world [^28] For example, hardware like an Apple M2 system on a chip interprets software to physical effect, based on the ARM instruction set architecture..
-Simp-maxing is a matter of form, while w-maxing is a matter of function [bennett2025b].
+For example a program takes a form when it is written in a language, but when it is run on a computer it has a particular function or effect upon the world.
+Simp-maxing is a matter of form, while w-maxing is a matter of function [22].
 When we account for interpretation between form and function it becomes clear that software is a physical thing, like a tool.
@@
 This is what “weakness” measures.
-A tool that implies a weaker constraint can complete more tasks [bennettmaruyama2022a] than another tool [^29] To be exact, a weaker policy completes a superset of tasks, rather than just more tasks.
-This is because policies are compared in the context of an ostensive definition..
-By preferring weaker policies, one avoids unnecessary commitments, meaning the weakest policies complete a ‘wider range of tasks’ than others [^30] See ESM Supplementary Note S3 for extended discussion of Bennett’s Razor, weakness, and related connections (e.g.
-maximum entropy)..
+A tool that implies a weaker constraint can complete more tasks [10] than another tool.
+By preferring weaker policies, one avoids unnecessary commitments, meaning the weakest policies complete a ‘wider range of tasks’ than others (see ESM Note S3 for discussion of weakness and related connections (e.g.
+maximum entropy).).
 To complete a ‘wider range of tasks’ in the embodied sense, a tool must be both more sample and more energy efficient, and so the weakest policy is the most efficient in both senses.
@@
 ML systems do not typically optimise for either weaker constraints on function (w-maxing), or simplicity of form (simp-maxing).
-This may explain some of their relative inefficiency [bennett2022a, bennett2023b].
+This may explain some of their relative inefficiency [13, 15].
 Even if they were w-maxing, the capacity for w-maxing is severely limited by the abstraction layer.
-An abstraction layer can be a “causal isolator” [friston2023] between a software “mind” at higher levels of abstraction, and its environment at lower levels.
+An abstraction layer can be a “causal isolator” [35] between a software “mind” at higher levels of abstraction, and its environment at lower levels.
 Embodiment by another name.
 Different abstraction layers permit different upper bounds on policy weakness, in the same way that two bodies might differ in how able they are to achieve a given goal.
-Different embodiments give different upper bounds on ‘intelligence’ [bennett2024a].
+Different embodiments give different upper bounds on ‘intelligence’ [16].
 That biological self-organising systems adapt efficiently means they are able to embody weaker policies than today’s ML systems.
-This capacity hinges on the abstraction layer [bennett2024a].
+This capacity hinges on the abstraction layer [16].
 An abstraction layer is like a window through which only part of the world can be seen.
@@
 Hardware, not just software.
-In enactivist terms, an intelligent choice of abstraction layer is one that reduces the “big world” of all information to a “small world” of relevant information [vervaeke2012, vervaeke2013a, vervaeke2013b].
+In enactivist terms, an intelligent choice of abstraction layer is one that reduces the “big world” of all information to a “small world” of relevant information [95, 94, 93].
 In a typical computer, lower levels of abstraction tend to be static.
@@
 4 Formal Framework and Definitions
-The formal definitions below are required for the core claim and for Theorem 5.1.
-The surrounding narrative provides intuition, but the theorem and its assumptions depend on these definitions.
-However this paper is based on a formal mathematical framework and it does contain a very simple proof supporting the claim.
-The formal details are also useful to connect this work here to related mathematical and experimental work [bennett2025thesis, bennett2023b, bennett2023c, bennett2024a, bennett2024b, bennettwelshciaunica2024, bennett2025e].
-Stack Theory describes an adaptive system or computation by the binary constraints or boundary conditions that define it.
-It formalises embodiment as a choice of abstraction layer: a finite vocabulary of possible constraints (programs) representing the possible configurations of a body embedded in its environment [bennett2025thesis], which are used to express “statements” constraining the body.
-Policies and tasks are then defined using statements in this embodied language.
-Delegation, distribution and abstraction are captured by the task hierarchy.
-A longer conceptual motivation (including the environment/states/programs framing and links to enactivist and pancomputational interpretations) is provided in the Electronic Supplementary Materials (ESM) Note S5 and is not required to follow the theorem below.
+The core claim and Theorem 5.1 depend upon the Stack Theory primitives defined below.
+The surrounding narrative provides intuition.
+The ESM contains additional motivation, proofs, and related technical material [19, 15, 14, 16, 17, 12, 18].
+Stack Theory describes an adaptive system in terms of the binary constraints or boundary conditions that define it.
+Embodiment is a choice of abstraction layer.
+This is a finite vocabulary of possible constraints (programs) representing the possible configurations of a body embedded in its environment [19].
+These programs are used to express “statements” constraining the body.
+Policies and tasks are defined as statements in this embodied language, and delegation is captured by how task constraints are arranged across layers.
+ESM Note S5 provides a longer conceptual motivation and links to related perspectives.
 For example, if a human raises their arm, then they are expressing a statement in their embodied language.
@@
 If the $\mathfrak{v}$-task describes a ‘fit’ organism, then the outputs it contains are ‘correct’ according to natural selection.
-A system [ciaunica2023b] that conformed to the free energy principle (FEP) would be a $\mathfrak{v}$-task for which $\langle I,O\rangle$ minimised surprisal.
-In other words, a $\mathfrak{v}$-task can describe an organism as behaviour that serves homeostatic and reproductive goals.
-For readers who wish to check the formal dependencies (see also [bennett2025thesis, bennett2024appendices]), the key definitions used in the theorem are summarised below.
+A $\mathfrak{v}$-task $\langle I,O\rangle$ can represent biological selection pressures by specifying which outputs are viable for which inputs.
+For example, an FEP-style agent can be treated as a task whose admissible futures minimise surprisal under an appropriate generative model [27, 39].
+For readers who wish to check the formal dependencies (see also [19, 21]), the key definitions used in the theorem are summarised below.
   1. 1.
-    A set $\Phi$ is assumed, whose elements are called states.
-A declarative program is a set $f$ of states, and $f$ is true relative to $\phi\in\Phi$ iff $\phi\in f$.
-$P=2^{\Phi}$ is the set of all programs.
-Embodiment is formalised as an abstraction layer, for which I single out a finite set $\mathfrak{v}\subseteq P$ called a vocabulary.
-A choice of vocabulary $\mathfrak{v}$ implies a formal language
-    $L_{\mathfrak{v}}=\{l\subseteq\mathfrak{v}:\bigcap l\neq\emptyset\}.$
+    Environment, programs, vocabularies.
+Fix a nonempty set $\Phi$ of mutually exclusive states.
+A (declarative) program is a set $p\subseteq\Phi$, and $p$ is true at $\phi\in\Phi$ iff $\phi\in p$.
+Let $P:=2^{\Phi}$ be the set of all programs.
+Embodiment is formalised as an abstraction layer, for which I single out a finite vocabulary $\mathfrak{v}\subseteq P$.
+    Intuition. A program is a set of states, and a vocabulary is the finite set of predicates the layer can express.
   2. 2.
-    Members of $L_{\mathfrak{v}}$ are called statements, and a statement $l$ is considered true relative to a state $\phi$ iff $\phi\in\bigcap l$.
-A completion of a statement $l_{1}$ is any statement $l_{2}$ s.t.
-$l_{1}\subseteq l_{2}$.
-    The extension of a statement $l$ is the set $E_{l}$ of all its completions, s.t. $x\in L_{\mathfrak{v}}$ has
-    $E_{x}=\{y\in L_{\mathfrak{v}}:x\subseteq y\}.$
-    The extension of a set $X\subseteq L_{\mathfrak{v}}$ of statements is
-    $E_{X}=\bigcup\limits_{x\in X}E_{x}.$
-    The weakness of a policy $x$ is the cardinality of its extension $\lvert E_{x}\rvert$.
+    Statements and truth sets. The language generated by $\mathfrak{v}$ is
+    $L_{\mathfrak{v}}:=\{\;l\subseteq\mathfrak{v}:T(l)\neq\emptyset\;\},\qquad T(l):=\bigcap_{p\in l}p\subseteq\Phi,$
+    where $T(l)$ is the truth set (truth conditions) of the (conjunctive) statement $l$. By convention, $T(\emptyset)=\Phi$. A statement $l$ is true at state $\phi$ iff $\phi\in T(l)$.
+    Intuition. A statement is a conjunction of vocabulary programs. Its truth set is the states where all of them hold.
   3. 3.
-    A $\mathfrak{v}$-task $\alpha$ is a pair $\langle I_{\alpha},O_{\alpha}\rangle$ where $I_{\alpha}\subset L_{\mathfrak{v}}$ are the inputs, $E_{I_{\alpha}}$ is the set of outputs available given $I_{\alpha}$, and $O_{\alpha}\subset E_{I_{\alpha}}$ are the correct outputs.
-All tasks form a hierarchy.
-If $\alpha\sqsubset\omega$ then $\omega$ is a parent of $\alpha$, meaning $I_{\alpha}\subset I_{\omega}$ and $O_{\alpha}\subseteq O_{\omega}$.
-A parent is higher level than its children.
-A parent task is delegated to child tasks, and as each task is embodied this describes control being delegated within a system.
+    Completions, extensions, weakness.
+A completion of a statement $l$ is any statement $l^{\prime}\in L_{\mathfrak{v}}$ such that $l\subseteq l^{\prime}$.
+The extension of $l$ is the set of all its completions.
+    $\mathrm{Ext}(l):=\{\;l^{\prime}\in L_{\mathfrak{v}}:l\subseteq l^{\prime}\;\}.$
+    For a set $X\subseteq L_{\mathfrak{v}}$, define $\mathrm{Ext}(X):=\bigcup_{x\in X}\mathrm{Ext}(x)$.
+(Throughout, $\mathrm{Ext}(\,\cdot\,)$ is always taken in the ambient language $L_{\mathfrak{v}}$ determined by context.) The weakness of a policy $\pi$ is the cardinality of its extension $|\mathrm{Ext}(\pi)|$.
+    Intuition.
+A completion adds further constraints while staying compatible with what has already been fixed.
+Weakness counts how many distinct completions remain available in the agent’s own embodied language.
+It is therefore a measure of optionality or autonomy at that abstraction layer.
+Weakness is not a syntactic notion like description length.
+A policy can be very short yet highly committal, or long yet weak, depending on how the vocabulary factors the world.
+Changing embodiment changes the vocabulary and therefore changes which refinements count, which is intended.
   4. 4.
-    A policy is a statement $\pi$ and a policy is correct for $\alpha$ iff
-    $\pi\in\Pi_{\alpha}=\{\pi\in L_{\mathfrak{v}}:E_{I_{\alpha}}\cap E_{\pi}=O_{\alpha}\}.$
-    Given input $i\in I_{\alpha}$ and a policy $\pi$, to infer is to choose $o\in E_{i}\cap E_{\pi}$.
+    Tasks and hierarchy.
+A $\mathfrak{v}$-task $\alpha$ is a pair $\langle I_{\alpha},O_{\alpha}\rangle$ where $I_{\alpha}\subseteq L_{\mathfrak{v}}$ is a set of admissible inputs and $O_{\alpha}\subseteq\mathrm{Ext}(I_{\alpha})$ is a set of correct outputs.
+Tasks form a hierarchy.
+If $\alpha\sqsubset\omega$ then $\omega$ is a parent of $\alpha$, meaning $I_{\alpha}\subseteq I_{\omega}$ and $O_{\alpha}\subseteq O_{\omega}$.
+    Intuition. A task specifies which outputs count as correct for each input. The hierarchy orders tasks by adding constraints.
+  5. 5.
+    Policies, correctness, inference, learning. A policy is a statement $\pi\in L_{\mathfrak{v}}$. A policy is correct for $\alpha$ iff
+    $\pi\in\Pi_{\alpha}:=\{\pi\in L_{\mathfrak{v}}:\mathrm{Ext}(I_{\alpha})\cap\mathrm{Ext}(\pi)=O_{\alpha}\}.$
+    Given input $i\in I_{\alpha}$ and a policy $\pi$, to infer is to choose $o\in\mathrm{Ext}(i)\cap\mathrm{Ext}(\pi)$.
 A task is complete if $o\in O_{\alpha}$.
-Given a task definition $\alpha$ s.t.
-$\alpha\sqsubset\omega$, to learn is to choose a policy $\pi\in\Pi_{\alpha}$, and $\omega$ has been learned from $\alpha$ if one chooses $\pi\in\Pi_{\omega}$.
-  5. 5.
+Given $\alpha\sqsubset\omega$, to learn is to choose a policy $\pi\in\Pi_{\alpha}$, and $\omega$ has been learned from $\alpha$ if one chooses $\pi\in\Pi_{\omega}$.
+    Intuition. A policy is a statement that filters which outputs are allowed. Learning chooses a policy whose completions match the observed input output pairs.
+  6. 6.
     If $\alpha_{1}\sqsubset\alpha_{2}\sqsubset\omega$ then $\alpha_{1}$ contains fewer examples of $\omega$ than $\alpha_{2}$.
@@
 For example $\mathfrak{a}$ is more intelligent than $\mathfrak{b}$ if $\mathfrak{a}$ learns $\omega$ from either $\alpha_{1}$ or $\alpha_{2}$, while $\mathfrak{b}$ learns $\omega$ from $\alpha_{2}$ but not from $\alpha_{1}$.
-  6. 6.
+    Intuition. The point of this example is that tighter constraints at one layer limit how weak a compatible higher-layer policy can be. This is why delegation matters for efficiency.
+  7. 7.
     A causal identity for $\mathfrak{b}$ is a policy classifying the interventions undertaken by $\mathfrak{b}$.
-  7. 7.
+    Intuition. Collective identity is the existence of at least one policy that is simultaneously viable for the parts and for the whole.
+  8. 8.
     $\Gamma_{P}$ contains every task in every vocabulary.
 For every $\rho\in\Gamma_{P}$ there exists an uninstantiated-task function $\lambda_{\rho}:2^{P}\rightarrow\Gamma_{P}.$ $\lambda_{\rho}$ takes a vocabulary $\mathfrak{v}^{\prime}\in 2^{P}$ and returns the highest-level child of $\rho$ that is a $\mathfrak{v}^{\prime}$-task.
-  8. 8.
+    Intuition. An uninstantiated task is a task template that can be applied to different vocabularies. This lets the same selection pressure be expressed across layers.
+  9. 9.
 […]
+    Intuition. Utility increases when the correct-output set can be realised by many policies. This captures adaptability under the embodied language.
 I set out to compare computers with biological systems. Since both are stacks of abstraction layers, I define multilayer architecture (MLA) below to describe them in comparable terms.
-  • The stack of layers of increasing abstraction is a sequence of uninstantiated tasks $\langle\lambda^{0},\lambda^{1},\dots,\lambda^{n}\rangle$ s.t.
-$\lambda^{i+1}\sqsubset\lambda^{i}$.
+  • The stack of layers of increasing abstraction is a sequence of uninstantiated tasks $\langle\lambda^{0},\lambda^{1},\dots,\lambda^{n}\rangle$ such that $\lambda^{i+1}\sqsubset\lambda^{i}$.
 Lower in the stack is a lower level of abstraction.
-  • $\mathfrak{f}:2^{P}\times L_{\mathfrak{v}}\rightarrow 2^{P}$ is a function that takes a vocabulary $\mathfrak{v}$ and a policy $\pi\in L_{\mathfrak{v}}$, and returns a vocabulary
-    $\mathfrak{v^{\prime}}=\{f\in P:\exists o\in E_{\pi}(\bigcap o=f)\}.$
-  • The state of an MLA is a sequence of policies $\langle\pi^{0},\pi^{1},\dots,\pi^{n}\rangle$ and a sequence of vocabularies $\langle\mathfrak{v}^{0},\mathfrak{v}^{1},\dots,\mathfrak{v}^{n}\rangle$ such that
+  • Given a vocabulary $\mathfrak{v}\subseteq P$ and a policy $\pi\in L_{\mathfrak{v}}$, the induced abstractor is
+    $\mathfrak{f}(\mathfrak{v},\pi):=\{\,T(o):o\in\mathrm{Ext}(\pi)\,\}\subseteq P.$
+    That is, the next-layer vocabulary consists of the truth sets of completions compatible with $\pi$.
+  • The state of an MLA is a sequence of policies $\langle\pi^{0},\pi^{1},\dots,\pi^{n}\rangle$ and a sequence of vocabularies $\langle\mathfrak{v}^{0},\mathfrak{v}^{1},\dots,\mathfrak{v}^{n}\rangle$ such that for each $i<n$,
     $\mathfrak{v}^{i+1}=\mathfrak{f}(\mathfrak{v}^{i},\pi^{i})\quad\text{and}\quad\pi^{i}\in\Pi_{\lambda^{i}(\mathfrak{v}^{i})}.$
+Intuition.
+The stack records tasks at increasing abstraction, and each layer is constrained by the layer below.
+The abstractor tells you which higher-level predicates become available once a lower-level policy has fixed which completions are possible.
+The state records the current policy and induced vocabulary at each layer.
 I say the MLA is over-constrained when there exists $i<n$ such that $\Pi_{\lambda^{i}(\mathfrak{v}^{i})}=\emptyset$, and multilayer-causal-learning (MCL) is when the MLA is not over-constrained.
+Over-constraint means some layer cannot realise any policy compatible with the higher-level task.
+The MCL is the boundary where each layer still has at least one viable policy, so collective identity is possible.
 5 Delegated Control and Collective Identity
@@
 The probability that the parts of a system embody a collective policy is higher if they embody weaker policies.
-This is because weaker policies are more likely to generalise between $\mathfrak{v}$-tasks [bennett2025e].
+This is because weaker policies are more likely to generalise between $\mathfrak{v}$-tasks [18].
 Consider the aforementioned $\alpha_{1}$ and $\alpha_{2}$.
@@
 The difference is a matter of scale.
-Just as w-maxing facilitates causal learning at the scale of an individual [bennett2023c], in the context of an MCA it will facilitate learning at multiple scales, because the system as a whole is just another $\mathfrak{v}$-task.
+Just as w-maxing facilitates causal learning at the scale of an individual [14], in the context of an MCA it will facilitate learning at multiple scales, because the system as a whole is just another $\mathfrak{v}$-task.
 Mechanistically, “negotiation” between levels is represented as constraint compatibility between policy sets.
@@
 The cells are an abstraction layer in which the collective policy of the organ is expressed, like software.
-However, software is just a state of hardware [bennett2024a].
-Put another way, ‘organ’ is a label I ascribe to the the behaviour of some cells, and ‘human’ is a label I ascribe to the behaviour of some organs.
+However, software is just a state of hardware [16].
+Put another way, ‘organ’ is a label I ascribe to the behaviour of some cells, and ‘human’ is a label I ascribe to the behaviour of some organs.
 ‘x86’ is a label I ascribe to a product of human behaviour.
-It is a static abstraction layer an engineer has embodied in silicon and then left alone; ‘fire and forget’.
+It is a static abstraction layer an engineer has embodied in silicon and then left alone.
+‘fire and forget’.
 Humans act as the lower layer from which x86 emerges, but once it is embodied in silicon we disconnect it from our stack.
@@
 For tasks at higher levels of abstraction, capacity for w-maxing depends on lower levels.
-A policy $\pi\in\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$ implies an abstraction layer in the sense that for every $o\in E_{\pi}$ there exists $f\in P$ s.t.
-$\bigcap o=f$ [^31] I can substitute $o$ for a declarative program $f$ that has the same extension, because that means $f$ and $o$ are true of the same states, having the same ‘truth conditions’ so to speak..
-Hence I can transform policies to abstraction layers, using a function $\mathfrak{f}$ that takes the extension $E_{\pi}$ of a collective’s policy $\pi$, and returns a new vocabulary $\mathfrak{v^{\prime}}=\{f\in P:\exists o\in E_{\pi}(\bigcap o=f)\}$.
+A policy $\pi\in\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$ implies an abstraction layer in the sense that for every $o\in\mathrm{Ext}(\pi)$ there exists $p\in P$ such that $T(o)=p$ (i.e.
+$p$ and $o$ are true of exactly the same states).
+Hence I can transform policies to abstraction layers, using a function $\mathfrak{f}$ that takes the extension $\mathrm{Ext}(\pi)$ of a collective’s policy $\pi$, and returns a new vocabulary $\mathfrak{v^{\prime}}=\{\,T(o):o\in\mathrm{Ext}(\pi)\,\}$.
 I can then use a sequence of uninstantiated tasks $\langle\lambda^{0},\lambda^{1}...\lambda^{n}\rangle$ s.t $\lambda^{i+1}\sqsubset\lambda^{i}$ to represent selection pressures that define ‘correctness’ at every level.
@@
 To construct layers I would start with a vocabulary $\mathfrak{v}^{i}$ at $i=0$, for example implied by a collective of cells or a lower level system.
-If this system exhibits goal directed behaviour, then it is given by the $\mathfrak{v}^{i}$-task $\lambda^{i}(\mathfrak{v}^{i})$ and a policy $\pi^{i}\in\Pi_{\lambda(\mathfrak{v}^{i})}$.
+If this system exhibits goal directed behaviour, then it is given by the $\mathfrak{v}^{i}$-task $\lambda^{i}(\mathfrak{v}^{i})$ and a policy $\pi^{i}\in\Pi_{\lambda^{i}(\mathfrak{v}^{i})}$.
 $\pi^{i}$ implies behaviour.
 I translate policy into an abstraction layer using the aforementioned $\mathfrak{f}$.
-$\mathfrak{f}$ is applied to $E_{\pi^{i}}$ to obtain a vocabulary $\mathfrak{v}^{i+1}$, and then $\lambda^{i+1}$ is applied to $\mathfrak{v}^{i+1}$ to obtain a task $\lambda^{i+1}(\mathfrak{v^{i+1}})$.
+$\mathfrak{f}$ is applied to $\mathrm{Ext}(\pi^{i})$ to obtain a vocabulary $\mathfrak{v}^{i+1}$, and then $\lambda^{i+1}$ is applied to $\mathfrak{v}^{i+1}$ to obtain a task $\lambda^{i+1}(\mathfrak{v^{i+1}})$.
 If $\Pi_{\lambda^{i+1}(\mathfrak{v^{i+1}})}\neq\emptyset$, then there exists a common purpose or “self” at this level of abstraction and I select a another policy $\pi^{i+1}\in\Pi_{\lambda^{i+1}(\mathfrak{v^{i+1}})}$.
-In purely narrative terms, the implication of all this is that adaptation at higher levels depends on adaptation at lower levels.
-Below I state The Law of the Stack with a brief proof sketch; a full proof is provided in the Electronic Supplementary Materials.
+In narrative terms, adaptation at higher levels depends on adaptation at lower levels.
+I now state The Law of the Stack and give a brief proof sketch.
+The ESM provides a full proof and additional lemmas.
 Theorem 5.1 (The Law of the Stack).
@@
 $v_{i+1}=\mathfrak{f}(v_{i},\pi_{i}),\qquad\gamma_{i+1}=\lambda_{i+1}(v_{i+1}).$
-Then the achievable utility at layer i+1i\!+\!1 is upper-bounded by the weakness of the policy at layer $i$ in the following sense:
-$\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\;\leq\;2^{|E_{\pi_{i}}|}.$
+Then the achievable utility at layer i+1i\!+\!1 is upper-bounded by the weakness of the policy at layer $i$ in the following sense.
+$\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\;\leq\;2^{|\mathrm{Ext}(\pi_{i})|}.$
 Equivalently,
-$|E_{\pi_{i}}|\;\geq\;\log_{2}\!\big(\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\big).$
-In particular, increasing $\epsilon(\lambda_{i+1}(v_{i+1}))$ requires (at least) increasing $|E_{\pi_{i}}|$, i.e. weakening the policy at the lower layer.
+$|\mathrm{Ext}(\pi_{i})|\;\geq\;\log_{2}\!\big(\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\big).$
+In particular, increasing $\epsilon(\lambda_{i+1}(v_{i+1}))$ requires (at least) increasing $|\mathrm{Ext}(\pi_{i})|$, i.e. weakening the policy at the lower layer.
+Intuition.
+The bound says that higher-layer utility cannot grow without leaving enough slack in the realised policy one layer down.
+A narrow lower-layer policy induces a small next-layer vocabulary and this caps achievable outputs.
 Proof sketch.
@@
 Because $v_{i+1}$ is finite, $|L_{v_{i+1}}|\leq 2^{|v_{i+1}|}$.
-By construction of the abstractor, $v_{i+1}$ is the image of $E_{\pi_{i}}$ under the map $o\mapsto\bigcap o$, hence $|v_{i+1}|\leq|E_{\pi_{i}}|$.
-Combining yields $\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\leq 2^{|E_{\pi_{i}}|}$.
+By construction of the abstractor, $v_{i+1}$ is the image of $\mathrm{Ext}(\pi_{i})$ under the map $o\mapsto T(o)$, hence $|v_{i+1}|\leq|\mathrm{Ext}(\pi_{i})|$.
+Combining yields $\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\leq 2^{|\mathrm{Ext}(\pi_{i})|}$.
 ∎
+Proof intuition.
+The proof sketch counts how many distinct higher-layer behaviours exist given the induced vocabulary at layer i+1i{+}1.
+The induced vocabulary is bounded by the number of completions left open by the realised policy at layer $i$.
+5.2 Delegation and variational free energy
+The Law of the Stack is a structural constraint.
+It says that degrees of freedom in higher layers is bounded by lower-layer slack.
+This aligns with the Free Energy Principle (FEP), which frames adaptive behaviour as variational free energy minimisation [39, 40, 71].
+In the ESM (“Viability, lived history, and variational free energy”) I show that under a uniform viability prior, minimising free energy is equivalent to maximising the entropy (log-count) of viable future continuations.
+In Stack Theory terms, this is exactly w-maxing.
+Combining that equivalence with Theorem 5.1 yields the corollary that reducing free energy at an abstract layer requires sufficient delegation (sufficiently weak realised policies) at lower layers.
+Intuitively, if lower layers are over-constrained, the induced higher-layer language is small, so even a maximum-entropy posterior cannot keep many viable futures open.
+Corollary 5.1 (Delegation is necessary for lowering free energy).
+Fix a planning horizon and let $\mu$ denote the viability task at layer i+1i{+}1.
+Write $O_{\mu}\subseteq L_{v_{i+1}}$ for the set of viable future continuations at that horizon.
+For any policy $\pi_{i+1}$ at layer i+1i{+}1, let $\Omega_{\pi_{i+1}}=\mathrm{Ext}(\pi_{i+1})\cap O_{\mu}$ be the set of viable futures compatible with $\pi_{i+1}$, and let $q_{\pi_{i+1}}$ be the uniform distribution on $\Omega_{\pi_{i+1}}$.
+Under the uniform viability prior assumed in the ESM, the corresponding (base-2) free energy satisfies
+$F_{2}(q_{\pi_{i+1}})=\log_{2}|O_{\mu}|-\log_{2}|\Omega_{\pi_{i+1}}|.$
+Moreover, $\Omega_{\pi_{i+1}}\subseteq L_{v_{i+1}}$, hence $|\Omega_{\pi_{i+1}}|\leq|L_{v_{i+1}}|\leq 2^{|\mathrm{Ext}(\pi_{i})|}$ as in the proof of Theorem 5.1, where $\pi_{i}$ is the realised policy at layer $i$ that induces $v_{i+1}$.
+Therefore,
+$F_{2}(q_{\pi_{i+1}})\;\geq\;\log_{2}|O_{\mu}|-|\mathrm{Ext}(\pi_{i})|.$
+In particular, free energy minimisation at layer i+1i{+}1 is bottlenecked by $|\mathrm{Ext}(\pi_{i})|$: reducing $F_{2}(q_{\pi_{i+1}})$ requires sufficient lower-layer slack, i.e.
+a weak enough realised policy at the lower layer.
+Intuition.
+In active inference terms, free energy can only be reduced by keeping enough viable futures in play (high entropy).
+The Law of the Stack proves that if lower layers are frozen or over-constrained, there is an entropy bottleneck, so abstract-layer free energy has a floor.
 Intervening in the stack limits the extent to which control can be delegated, and thus limits capacity for w-maxing or MCL.
@@
 It is static.
-It cannot adapt, which is why there is demand for field programmable gate array (FPGA) [^32] An FPGA is like a CPU that can be reconfigured for specialised applications..
+It cannot adapt, which is why there is demand for field programmable gate array (FPGA) (a reconfigurable processor substrate).
 humanshumansISAISAabstractionsabstractions
@@
 After all, an increase in utility may increase the computational cost of searching for correct policies.
-Furthermore, a spatially extended environment constrains the size of vocabularies at all levels [bennett2024b], because bounded systems can contain only finite information [bekenstein1981].
+Furthermore, a spatially extended environment constrains the size of vocabularies at all levels [17], because bounded systems can contain only finite information [8].
 I mean lower levels learn an inductive bias for higher levels, balancing utility with other factors in service of goals at every scale.
@@
 A state analogous to cancer exists for distributed systems with insufficient delegation in adverse conditions.
-Some have argued cancer is a loss of collective identity [levin2021, davies2011].
+Some have argued cancer is a loss of collective identity [57, 30].
 In a biological system, when cells are isolated from the informational structure of their collective, they lose their high level identity.
@@
 If the constraints are too tight, then no correct policies will exist and the collective’s identity will be lost.
-The set of correct policies shrinks as task difficulty increases, meaning $\mathfrak{v}$-task $\alpha$ is easier than $\omega$ if $\mid\Pi_{\alpha}\mid>\mid\Pi_{\omega}\mid$.
-If the policy at a lower level i1i-1 changes such that the task at level $i$ becomes ‘harder’333333i.e.
-by reducing $\mid\Pi^{i}\mid$ at the higher level $i$., then this could result in there being no correct policy at level $i$.
-If there are no longer any correct policies at level $i$, then the only way the system can continue to function is for part of the collective to ‘splinter’ or break away from the group, because that will loosen the constraint on correct policies for both the breakaway part and the collective it leaves behind.
-Say I have a collective of cells $\lambda^{i}(\mathfrak{v}^{i})$ that includes a cell $\alpha\sqsubset\lambda^{i}(\mathfrak{v}^{i})$ s.t.
-$\mathfrak{f}(\mathfrak{v}^{i-1}_{\mathfrak{a}},{\pi^{i-1}_{\mathfrak{a}}})=\mathfrak{v}^{i}_{\mathfrak{a}}$.
-Now assume the policy $\pi^{i-1}_{\mathfrak{a}}$ changes to $\pi^{i-1}_{\mathfrak{b}}$ s.t.
-$\Pi_{\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})}=\emptyset$.
-A subset of the collective may continue functioning as before with a shared policy, but doing so requires discarding a part of the collective which cannot adopt that policy.
-This is akin to ‘selective forgetting’ [bennett2023c] of possible false negatives or positives in order to interpolate otherwise noisy data, but instead of forgetting data inconsistent with an otherwise good policy, I would discard a child task $\alpha\sqsubset\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})$ of $\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})$ when $\Pi_{\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})}=\emptyset$, so that I obtain a task $\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})\sqsubset\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})$ s.t.
-$\alpha\not\sqsubset\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})$ and $\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})$ has correct policies, meaning $\Pi_{\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})}\neq\emptyset$.
-$\alpha$ is now isolated from the informational structure of the collective and so may have an extremely broad set of ‘correct’ policies to choose from.
+Formal characterisation of the cancer-analogue condition
+To make the above precise, let $\alpha_{1},\dots,\alpha_{m}$ be $v$-tasks representing parts of a distributed system, with policy sets $\Pi_{\alpha_{j}}\subseteq L_{v}$, and let the collective task be $\omega$ with collective-policy set $\Pi_{\omega}:=\bigcap_{j=1}^{m}\Pi_{\alpha_{j}}$ (collective identity at $v$ requires $\Pi_{\omega}\neq\emptyset$).
+Definition 6.1 (Restricted policies, over-constraint, and splinters).
+To model top-down boundary conditions, fix restricted policy sets $\widehat{\Pi}_{\alpha_{j}}\subseteq\Pi_{\alpha_{j}}$ and define the restricted collective set $\widehat{\Pi}_{\omega}:=\bigcap_{j=1}^{m}\widehat{\Pi}_{\alpha_{j}}$.
+We call the system over-constrained at $v$ (under the boundary conditions) if $\widehat{\Pi}_{\omega}=\emptyset$.
+A splinter is a non-empty index set $S\subseteq\{1,\dots,m\}$ such that $\bigcap_{j\in S}\widehat{\Pi}_{\alpha_{j}}\neq\emptyset$.
+Proposition 6.1 (Over-constraint implies loss of collective policy; any realised behaviour is splintered).
+Assume at least one part remains locally viable, i.e.
+$\widehat{\Pi}_{\alpha_{j}}\neq\emptyset$ for some $j$.
+If the system is over-constrained at $v$ (so $\widehat{\Pi}_{\omega}=\emptyset$), then there is no policy simultaneously feasible for all parts.
+Moreover, for any realised policy $\pi\in L_{v}$, the set of parts for which $\pi$ remains feasible,
+$S(\pi):=\{\,j\in\{1,\dots,m\}:\pi\in\widehat{\Pi}_{\alpha_{j}}\,\},$
+is a proper subset of $\{1,\dots,m\}$. In particular, any continued function can only proceed via a splinter (Definition 6.1).
+Proof.
+If some $\pi$ were feasible for all parts then $\pi\in\bigcap_{j=1}^{m}\widehat{\Pi}_{\alpha_{j}}=\widehat{\Pi}_{\omega}$, contradicting $\widehat{\Pi}_{\omega}=\emptyset$.
+Hence every realised policy fails at least one part’s constraints, so $S(\pi)\neq\{1,\dots,m\}$ for all $\pi\in L_{v}$.
+If $\widehat{\Pi}_{\alpha_{j}}\neq\emptyset$ then $\{j\}$ is a splinter since $\bigcap_{k\in\{j\}}\widehat{\Pi}_{\alpha_{k}}=\widehat{\Pi}_{\alpha_{j}}\neq\emptyset$.
+$\square$
+As task difficulty increases (or top-down control tightens), the set of correct collective policies shrinks.
+A task $\alpha$ is easier than $\omega$ if $|\Pi_{\alpha}|>|\Pi_{\omega}|$.
+In Stack Theory terms, a cancer-like breakdown occurs when a policy change at layer i1i{-}1 induces a new collective task at layer $i$ with no correct policies, $\Pi_{\lambda^{i}(v^{i})}=\emptyset$, while components retain the capacity for independent action.
+When $\Pi_{\lambda^{i}(v^{i})}=\emptyset$, the collective has no viable way to act as a coherent unit.
+Any continued function therefore requires dropping constraints, parent task splintering into ever narrower child tasks until those tasks are narrow enough to admit a policy.
+Part of the collective stops participating in the collective policy and reverts to a lower-level policy set with a much larger extension (more autonomy).
+This is a form of “selective forgetting” [14] applied to constraints rather than data.
 This is analogous to cancer.
 $\alpha$ is isolated and can now pursue goals which do not align with those of the collective.
-This suggests systems should be minimally constrained so as to ensure graceful degradation or ‘sloppy fitness’ [agosta2010], maintaining core functionality in as wide a variety of unanticipated circumstances as possible.
+This suggests systems should be minimally constrained so as to ensure graceful degradation or ‘sloppy fitness’ [2], maintaining core functionality in as wide a variety of unanticipated circumstances as possible.
 The more difficult the environment or the more top-down the system is controlled, the more tightly constrained behaviour becomes until the system becomes over-constrained and splinters.
@@
 Put another way, there will be less incentive for the parts of a system to splinter, as that would not greatly increase the number of correct policies available to them.
-To summarise; a state analogous to cancer occurs in a distributed system when there is an over-reliance on top-down causation to make the parts of a system conform to a collective identity they are ill suited to, under the circumstances.
-7 Hybrid Agency as Boundary-Condition Design
-Consider the aforementioned example of the organoid adapted to play the game Pong with notable sample efficiency relative to conventional reinforcement-learning baselines [khajehnejad2024].
+A state analogous to cancer occurs in a distributed system when there is an over-reliance on top-down causation to make the parts of a system conform to a collective identity they are ill suited to, under the circumstances.
+7 Hybrid Creation as Weak Boundary-Condition Design
+Consider the aforementioned example of the organoid adapted to play the game Pong with notable sample efficiency relative to conventional reinforcement-learning baselines [53].
 Hybrid systems like organoids are not fixed interpreters of software-level policy, but collectives whose lower-level parts retain bottom-up agency.
 In such systems, hybrid creation is the act of designing an interface and environment that reshapes the feasible policy space of the biological components.
-Rather than “programming” the parts directly, the designer engineers boundary conditions so that the remaining viable collective policies implement a target input–output relation at the interface.
-The designer selects boundary conditions that reshape each component’s viable policy set, hoping that the intersection admits a collective policy that realises the desired input–output mapping at the interface.
+Rather than “programming” the parts directly, the designer engineers boundary conditions so that the remaining viable collective policies implement a target input-output relation at the interface.
+Practically, hybrid creation is iterative:
+  1. 1.
+    choose boundary conditions, then
+  2. 2.
+    observe emergent collective behaviour, then
+  3. 3.
+    tighten or relax constraints to steer the intersection of viable policy sets toward the target interface behaviour and
+  4. 4.
+    repeat.
+Similar boundary-condition design issues arise in other bio-hybrid constructs (e.g.
+xenobot-like cell collectives) and in soft-robotic substrates where not just software but morphology and material properties shape the feasible low-level policy space.
+The designer selects boundary conditions that reshape each component’s viable policy set, hoping that the intersection admits a collective policy that realises the desired input-output mapping at the interface.
 According to The Law of the Stack, achieving higher utility at an interface layer requires sufficiently weak policies at lower layers.
@@
 8 Conclusions and Future Research
-In this paper I used Stack Theory to describe MCAs as stacks of abstraction layers.
-In this view policies are embodied at each layer, approximating causality in service of adaptation, and cause is best approximated by weaker policies (w-maxing) [bennett2023c].
-Formally, I prove a theorem (The Law of the Stack 5.1) showing w-maxing in higher layers requires sufficient w-maxing in lower layers.
-One reason artificial intelligence tends to be less adaptable is because it rests upon an inflexible stack of human abstractions.
-This claim is directed at contemporary, predominantly fixed-stack AI architectures.
-It is not a fundamental limitation on future computational or hybrid systems.
-My claim is that, to build more robust systems, control should be delegated to the greatest extent possible.
-Obviously this must be balanced against desired functionality.
-Policies must be aligned.
-The system must still discriminate between correct and not.
-In the earlier sections I discussed the doctrine of mission command.
-What I am suggesting is that the concept of mission command should be applied to systems in general.
-At least, to the extent that adaptability is desired.
-Put another way, any constraints which are not strictly necessary to maintain the integrity of the system should be eliminated [^34] What counts as “strictly necessary” is relative to design goals and value judgements; see ESM Supplementary Note S4..
-Humans tend to prefer ‘additive’ solutions [adams2021], by which I mean if a problem can be solved by adding a rule or subtracting a rule, we humans tend to add a rule.
-To effectively delegate control, we should subtract rules instead.
-There are also software-first approaches that approximate aspects of this delegation picture, including federated learning [nori2025federated], constructivist AGI programmes [thorisson2012, goertzel2014], and active inference [Friston2010, friston2013].
+Biological systems often adapt efficiently because they delegate adaptive control across scales.
+Using Stack Theory [19], I formalised delegation as the depth to which adaptation is permitted in a stack of abstraction layers and proved The Law of the Stack (Theorem 5.1).
+Utility one layer up is bounded by the weakness of the realised policy one layer down.
+Delegation is a structural prerequisite for high-level adaptability.
+This constraint is compatible with active inference.
+Under a simple viability prior, minimising variational free energy is equivalent to maximising entropy over viable futures (w-maxing; ESM “Viability, lived history, and variational free energy”).
+Corollary 5.1 shows that free energy minimisation inherits the same bottleneck: abstract-layer free energy cannot be reduced without sufficient lower-layer slack.
+The implication for AI is not that computation cannot be intelligent, but that contemporary AI stacks tend to concentrate adaptation at high software layers while freezing lower layers, thereby throttling achievable weakness and adaptation efficiency.
+For hybrid agents (organoids, soft robots, and human–AI organisations) the design problem becomes boundary-condition engineering: constrain parts just enough to elicit target collective behaviour while preserving a non-empty intersection of viable policies (collective identity).
+When that intersection collapses, the system can fragment into competing subagents (Section 6).
+Future work should test these predictions by varying which layers are permitted to adapt (e.g.
+hardware-in-the-loop learning, morphologically adaptive substrates, and biologically integrated controllers) and measuring resulting degrees of freedom, free energy, and robustness.
+More broadly, this paper is a stepping stone toward a unified account of delegation, weakness/max-entropy heuristics, and multiscale agency in the larger Stack Theory programme [19].
+A practical design implication is that, when adaptability is desired, constraints should be weakened as far as possible without breaking the system’s intended function.
+Constraints that are not strictly necessary relative to the design goal should be removed (What counts as “strictly necessary” is relative to design goals and value judgements.
+See ESM Supplementary Note S4.).
+Humans often prefer ‘additive’ solutions [1], meaning if a problem can be solved by adding a rule or subtracting a rule, we tend to add a rule.
+However The Law of the Stack suggests a preference for subtractive solutions might be beneficial in many circumstances.
+To speculate about future research beyond experimental verification, Stack Theory is a lens that can be applied generally.
+For example, a number of software-first approaches to AI already approximate aspects of this delegation picture, including federated learning [52], constructivist AGI programmes [90, 42], and active inference [39, 40].
 An expanded discussion (including AERA, Hyperon/ActPC-Chem, and links to scale-free active inference) is provided in the ESM (Supplementary Note S6).
-Future research might consider how adaptive hardware can be designed, to delegate control.
-Multilayer-causal-learning (MCL) provides a formal basis for ‘soft robotics’ [man2019].
-Soft robotics is a proposed class of robots made of ‘soft’ materials that adapt.
-I have formalised adaptive materials as adaptation lower in the stack, providing a formal foundation for soft robotics.
-Self-assembling nanoparticle networks [borghi2024_brainlikehw] are a step in this direction.
-They can be used to form reconfigurable logic gates [paroli2023_receptron], facilitating adaptation at very low levels of abstraction.
-Taken to its logical conclusion, what might be required is a sort of artificial life, with small parts interacting form complex systems from the small scales and low levels of abstraction up [suzuki2006, ikegami2008, ikegami2007simulating].
-As for the inductive bias mentioned in the introduction [chollet2019], it boils down to the choice of abstraction layer [bennett2024a].
-I have described a means of learning a suitable abstraction layer.
-In theory, a multilayer-causal-learning system should be sufficient to solve ARC-AGI benchmark problems in terms of both sample and energy efficiency.
-Future research should test this empirically through (i) agent-based simulations of collectives with tunable delegation and abrupt environment shifts, measuring sample/energy efficiency and robustness; (ii) comparative benchmarks in which adaptation is confined to high-level software state versus systems with reconfigurable substrates (e.g.
-adaptive firmware or FPGAs); and (iii) hybrid experiments in which boundary conditions on biological components are systematically varied to quantify trade-offs.
-A more detailed set of candidate evaluation designs is provided in ESM Note S8.
-Beyond hardware, cyberphysical systems in general can be evaluated in terms of abstraction, distribution and delegation.
-I might consider how states analogous to cancer could arise in human organisational structures, or how control can be delegated throughout an organisation to facilitate robustness and efficiency.
-For example, this could be applied to facilitate transformational bottom-up rather than incremental top-down adaptation to environmental changes.
-Free-market trade is a great example of how delegation can facilitate adaptability in a large collective.
-As Hayek pointed out [hayek1945use], not all of the information is available to a central planner and so control is best delegated to the individual.
-The accumulation of resources within my formalism is implicit, rather than explicit.
-I have been concerned only with whether a system satisfies whatever arbitrary notion of correct I choose (i.e.
-does it achieve homeostatic and reproductive goals in whatever amount of time care about?), rather than how it evolves over time as a dynamical system.
-In other words, I have concerned myself only with whether a system arrives on time at the destination I want, not the journey it took to get there.
-Future research might formalise the accumulation of resources explicitly.
-One way to describe an evolving sequence of tasks over time could be something like a stack.
-To speculate about what my results imply for delegated systems like an economy, under less than ideal conditions resources may become concentrated in the hands of a few.
-The resulting system then delegates work, but not control.
-Those parts deprived of control may then not be able to construct a policy that aligns with the collective, and so they may splinter off and become ‘cancerous’.
-To prevent this, mechanisms can be put in place to preserve decentralisation without the need for intervention from outside the system (e.g.
-government intervening to break up a monopoly).
-Each system could be located along axes of redistribution and delegation, like an alternative political compass.
-Just enough redistribution of control and the system remains delegated.
-However, too much redistribution could have a similar effect, preventing parts from adapting to local conditions.
-An example of a mechanism that can effectively preserve delegation of control is adaptive rubber banding [mi2022]; a game mechanic (e.g.
-in Mario Kart) where success accrues penalties as well as resources, keeping the game competitive.
-Another possibility is AI regulation.
-For example the UK government recently initiated “Safeguarded AI” [dalrymple2024b, dalrymple2024a], a project that aims to construct an AGI [goertzel2014] of level ASL-4 [anthropic2023] in a cyberphysical system, to safeguard against rogue AI.
-The idea of a “rogue AI” from within a cyberphysical system is analogous to cancer, adopting an “unsafe” policy at odds with that of the collective.
-AI regulations may over-constrain systems and trigger the very cancerous state they seek to prevent.
-To reiterate the earlier point about subtractive solutions [adams2021], any guardrails which are not strictly necessary must be eliminated to make room for those that are.
-In summary, I have described systems in terms of delegation and abstraction.
-I proved that adaptability at higher levels of abstraction is constrained by adaptability at lower levels (The Law of the Stack).
-I have argued that one reason biological systems are more ‘intelligent’, meaning sample and energy efficient, is because they delegate adaptation down the stack.
-This Stack Theoretic perspective lays a foundation for future empirical research, showing how we might build more robust systems of any kind.
-9 Electronic Supplementary Materials
-9.1 Mathematical Results
-The Law of the Stack
-Theorem 9.1 (The Law of the Stack).
-Assume the formal setting defined in the main text: finite vocabulary $v\subseteq P$, language $L_{v}=\{\,\ell\subseteq v:\bigcap\ell\neq\emptyset\,\}$, extensions $E_{\pi}=\{\,y\in L_{v}:\pi\subseteq y\,\}$, $v$-tasks $\gamma=\langle I_{\gamma},O_{\gamma}\rangle$, correct policies $\Pi_{\gamma}=\{\pi\in L_{v}:E_{I_{\gamma}}\cap E_{\pi}=O_{\gamma}\}$, and utility $\epsilon(\gamma)=\max_{\pi\in\Pi_{\gamma}}(|E_{\pi}|-|O_{\gamma}|)$.
-Let $\langle\lambda_{0},\dots,\lambda_{n}\rangle$ be an MLA with abstractor $\mathfrak{f}$ and state $\langle(\pi_{0},\dots,\pi_{n}),(v_{0},\dots,v_{n})\rangle$ such that $v_{i+1}=\mathfrak{f}(v_{i},\pi_{i})$ and $\pi_{i}\in\Pi_{\lambda_{i}(v_{i})}$ for all $i<n$.
-Fix any $i<n$, and define $\gamma_{i+1}=\lambda_{i+1}(v_{i+1})$ where $v_{i+1}=\mathfrak{f}(v_{i},\pi_{i})$. Then
-$\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\;\leq\;2^{|E_{\pi_{i}}|}.$
-Equivalently,
-$|E_{\pi_{i}}|\;\geq\;\log_{2}\!\big(\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\big).$
-Thus, higher utility at layer i+1i\!+\!1 requires (at least) a weaker policy at layer $i$ in the sense of larger extension cardinality.
-Proof.
-We proceed by a sequence of elementary bounds that make the layer-to-layer dependency explicit.
-Lemma 1 (Languages are bounded by the power set).
-For any finite vocabulary $v$, the language $L_{v}$ satisfies $|L_{v}|\leq 2^{|v|}$.
-Proof. By definition, $L_{v}\subseteq\mathcal{P}(v)$ because every $\ell\in L_{v}$ is a subset of $v$. Since $|\mathcal{P}(v)|=2^{|v|}$ for finite $v$, we have $|L_{v}|\leq 2^{|v|}$. $\square$
-Lemma 2 (Extensions live inside the language).
-If $\pi\in L_{v}$, then $E_{\pi}\subseteq L_{v}$, hence $|E_{\pi}|\leq|L_{v}|$.
-Proof. By definition $E_{\pi}=\{y\in L_{v}:\pi\subseteq y\}$, so $E_{\pi}\subseteq L_{v}$. Taking cardinalities yields $|E_{\pi}|\leq|L_{v}|$. $\square$
-Lemma 3 (Utility implies a lower bound on language size).
-Let $\gamma$ be a $v$-task. Then
-$\epsilon(\gamma)+|O_{\gamma}|\leq|L_{v}|.$
-Proof.
-Because $v$ is finite, $L_{v}$ is finite, hence $\Pi_{\gamma}\subseteq L_{v}$ is finite, so the maximum defining $\epsilon(\gamma)$ is attained by some $\pi^{\star}\in\Pi_{\gamma}$.
-By definition,
-$\epsilon(\gamma)=|E_{\pi^{\star}}|-|O_{\gamma}|\quad\Rightarrow\quad\epsilon(\gamma)+|O_{\gamma}|=|E_{\pi^{\star}}|.$
-By Lemma 2, $|E_{\pi^{\star}}|\leq|L_{v}|$. Combining yields the claim. $\square$
-Lemma 4 (The abstractor output vocabulary is an image of an extension).
-Let $T:L_{v_{i}}\to P$ be the map $T(o)=\bigcap o$. Under the MLA definition,
-$v_{i+1}=\mathfrak{f}(v_{i},\pi_{i})=\{\,T(o):o\in E_{\pi_{i}}\,\},$
-so $v_{i+1}$ is the image of $E_{\pi_{i}}$ under $T$. Consequently,
-$|v_{i+1}|\leq|E_{\pi_{i}}|.$
-Proof. This is just re-writing the definition of $\mathfrak{f}$ in functional form, and then applying the general fact that for finite sets $A$ and any function $g$, $|g(A)|\leq|A|$. $\square$
-Conclusion.
-Apply Lemma 3 to $\gamma_{i+1}$ in vocabulary $v_{i+1}$ to obtain
-$\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\leq|L_{v_{i+1}}|.$
-Apply Lemma 1 to bound $|L_{v_{i+1}}|\leq 2^{|v_{i+1}|}$, giving
-$\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\leq 2^{|v_{i+1}|}.$
-Finally apply Lemma 4: $|v_{i+1}|\leq|E_{\pi_{i}}|$. Since $2^{x}$ is monotone increasing,
-$2^{|v_{i+1}|}\leq 2^{|E_{\pi_{i}}|}.$
-Combining the last two displays yields
-$\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\leq 2^{|E_{\pi_{i}}|},$
-and the logarithmic form follows by applying $\log_{2}$ (also monotone). ∎
-Remark.
-In the thesis version of this result, the intuition is phrased as: larger vocabularies at layer i+1i{+}1 permit weaker (more general) policies, hence more adaptation at higher layers.
-The proof here makes the dependency explicit by bounding the achievable utility at layer i+1i{+}1 as a monotone function of the extension size $|E_{\pi_{i}}|$ at layer $i$, without requiring any additional monotonicity assumptions about $\epsilon(\lambda(\cdot))$ under set inclusion.
-Hybrid case study formalisation (Pong organoid)
-Let $C$ be a set of cells.
-For each cell $i\in C$, let $\Pi_{i}$ denote the set of viable policies for that cell under some interface/environmental conditions.
-The set of viable collective policies for the organoid is $\Pi_{C}=\bigcap_{i\in C}\Pi_{i}$.
-A conventional computer interface (electrodes, stimulus generator, game process) imposes boundary conditions $c$ that restrict which policies remain viable; write this restriction schematically as an operator $R_{c}$ such that $\Pi_{i}^{(c)}=R_{c}(\Pi_{i})$ and $\Pi_{C}^{(c)}=\bigcap_{i\in C}\Pi_{i}^{(c)}$.
-The design objective is to choose $c$ so that (i) $\Pi_{C}^{(c)}\neq\emptyset$ and (ii) the remaining collective policies implement a target output set $O_{\text{pong}}$ at the interface.
-In this view, building the hybrid is a constraint-design problem: impose the weakest (least exclusionary) boundary conditions that make $O_{\text{pong}}$ achievable while preserving non-empty intersection (collective identity).
-9.2 Supplementary Notes and Extended Discussion
+Stack Theory might also inform the design of ‘soft robots’ [58].
+Soft robots are a proposed class of robots made of ‘soft’ materials that adapt.
+I have formalised adaptive materials as adaptation lower in the stack, which can be used as a formal foundation for soft robotics.
+Self-assembling nanoparticle networks [24] are a step in this direction.
+They can be used to form reconfigurable logic gates [70], facilitating adaptation at very low levels of abstraction.
+Taken to its logical conclusion, future AI may more closely resemble artificial life [85, 45, 46].
+Finally, Stack Theory may be extended to economic and organisational systems.
+For example, free-market trade could be seen as a classic argument for delegation.
+As Hayek pointed out [44], a great deal of relevant information is spread among a population, unavailable to a central planner.
+This motivates delegation of control, to shorten feedback loops.
+However, the present formalism treats resources implicitly via task correctness, so connecting these arguments to economic dynamics would require extending Stack Theory to explicit time-series [23] and resource accumulation.
+Additional speculative extensions and applications (including organisational and economic systems, AI safety regulation, and temporal extensions of stacks) are collected in ESM Note S9.
+In summary, I have applied Stack Theory to explore how delegation of control affects adaptability.
+The Law of the Stack shows how adaptability at higher levels of abstraction is constrained by adaptability at lower levels, and delegation is structurally necessary to minimise free energy.
+This lays a foundation for a broad range of future theoretical and empirical research, showing how systems of any kind might be made more robust and adaptable.
+Acknowledgements
+First, I’d like to thank the anonymous peer reviewers for their detailed and helpful feedback.
+Second, I’d like to thank Anna Ciaunica, Peter Watts, Sean Welsh, Ricard Solé, Karl Friston, Antonio Damasio, Kingson Man, Paolo Milani, Ben Goertzel, Kristinn R.
+Thórisson, Elija Perrier and Noel Hinton for their support and feedback on this paper.
 S1. Working definition of intelligence
 There are many definitions of intelligence, and the merits of each are beyond the scope of this paper.
-Mine is based on [bennett2025thesis], which frames intelligence as the ability to complete a wide range of tasks, which is sample and energy efficiency in adaptation.
-It is similar to Wang’s definition of intelligence as adaptation with limited resources [wang2019], and Hutter’s earlier definition of intelligence as ‘the ability to satisfy goals in a wide range of environments’.
+Mine is based on [19], which frames intelligence as the ability to complete a wide range of tasks, which is sample and energy efficiency in adaptation.
+It is similar to Wang’s definition of intelligence as adaptation with limited resources [96], and Hutter’s earlier definition of intelligence as ‘the ability to satisfy goals in a wide range of environments’.
 Mainstream AI often operationalizes ‘intelligence’ as general-purpose task performance.
 By that standard, one may argue that current large language models are ‘more intelligent’ than many biological systems.
+However that definition is fundamentally flawed, as it fails to express the comparative efficiency with which humans acquire general purpose competence [16, 96, 26].
+In theory even a lookup table or rote-learning response can possess general purpose competence, given enough resources.
 This paper instead treats intelligence as the cost of adaptation (sample and energy efficiency), and the results should be read accordingly.
 S2. Levels of abstraction beyond computing
-The notion of levels of abstraction is more typically used in pure computer science settings, but can be applied to any system [bennett2025thesis].
+The notion of levels of abstraction is more typically used in pure computer science settings, but can be applied to any system [19].
 A higher level of abstraction is the behaviour of the layers below.
@@
 Weakness is formally defined in the appendix.
-Related work proved w-maxing to be optimal regardless of the abstraction layer [bennett2025e, bennett2023b, bennett2024a, bennett2025thesis].
+Related work proved w-maxing to be optimal regardless of the abstraction layer [18, 15, 16, 19].
 In narrative terms, assume two policies $x$ and $y$ both constrain the system to examples of “correct” behaviour given past inputs, but differ in weakness.
@@
 For example, if I assume the lawyer was paid at least $\mathdollar 5$, I am more weakly constrained than if I assume that lawyer was paid at least $\mathdollar 6$.
-As an epistemological razor, this means “explanations should be no more specific than necessary” or alternatively “rules should constrain no more than necessary” (see Bennett’s Razor in [bennett2023b]).
-This is a bit like the principle of maximum entropy [jaynes1957a, jaynes1957b], difference being that w-maxing also accounts for the abstraction layer, which is important because abstraction can render complexity subjective [bennett2024b].
-For a rigorous definition of learning and efficiency, see definitions in the online repository for the thesis [bennett2024appendices].
+This example also helps separate weakness from simplicity.
+The difference between $\mathdollar 5$ and $\mathdollar 6$ is not primarily about how hard the rule is to write down.
+It is about how many further commitments remain compatible with the rule, given the agent’s available distinctions.
+Weakness is defined relative to an abstraction layer.
+The layer fixes what counts as a meaningful refinement because it fixes what the system can jointly discriminate and jointly realise.
+If the embodiment changes, then the induced vocabulary changes, and so does the completion structure that weakness counts.
+Many familiar heuristics can be seen as special cases of the same idea once this completion structure is made explicit.
+If one treats all completions that remain possible at a layer as equally plausible, then choosing a correct policy that maximises $|\mathrm{Ext}(\pi)|$ is equivalent to maximising entropy over viable futures inside that layer.
+This parallels least commitment and version space intuitions, but it is stated directly in terms of the agent’s embodied language rather than in terms of an encoding chosen by an observer.
+That is what lets weakness be used across layers in Stack Theory, where a policy at one layer induces the next layer vocabulary via the abstractor map.
+In the language of agency, weakness can be read as a quantitative proxy for autonomy at a layer.
+A weaker policy leaves more degrees of freedom (more “ways of being right”) while still satisfying the same observed constraints.
+In multiscale and hybrid systems this autonomy matters because it increases the chance that component policies remain compatible with a higher-level collective policy, preserving collective identity.
+As an epistemological razor, this means “explanations should be no more specific than necessary” or alternatively “rules should constrain no more than necessary” (see Bennett’s Razor in [15]).
+This is a bit like the principle of maximum entropy [50, 49], difference being that w-maxing also accounts for the abstraction layer, which is important because abstraction can render complexity subjective [17].
+For a rigorous definition of learning and efficiency, see definitions in the online repository for the thesis [21].
 It includes proofs showing w-maxing is both necessary and sufficient to maximise the probability of ‘generalising’ from one task to another, and thus adaptability.
@@
 “Strictly necessary” is a value judgement, or an implicit choice of what ought to be.
-In the thesis [bennett2025thesis] on which this is based this ought from which other descriptions of ought follow is derived from the state transitions of the environment.
+In the thesis [19] on which this is based this ought from which other descriptions of ought follow is derived from the state transitions of the environment.
 These transitions preserve some aspects of that environment, while destroying others.
@@
 S5. Conceptual motivation for the formal setting
-Stack Theory frames the description of an objective reality in terms of choices of subjective lens (an abstraction layer) [bennett2025thesis].
-The full justification for these definitions is out of scope and addressed at length elsewhere [bennett2024a, bennett2025thesis, bennett2025e].
+Stack Theory frames the description of an objective reality in terms of choices of subjective lens (an abstraction layer) [19].
+The full justification for these definitions is out of scope and addressed at length elsewhere [16, 19, 18].
 However I will provide a brief overview.
@@
 Whatever the environment is, there must be change or difference for there to be anything at higher levels of abstraction.
-Hence there must be a set of states where each state is one point of difference [^35] Nothing is assumed about what these states might ontologically be or epistemologically contain..
+Hence there must be a set of states where each state is one point of difference (Nothing is assumed about what these states might ontologically be or epistemologically contain.).
 Because these states must be mutually exclusive, this frames time as difference.
@@
 Everything that exists is a set of facts, and each timeline preserves different sets of facts.
-A timeline is like an opinion; as time progresses some aspects of the environment are preserved and others are destroyed.
-Each timeline preserves different aspects, expressing a different “ought” which related work uses to explain the origins of goal directed behaviour [bennett2025thesis].
+A timeline is like an opinion.
+As time progresses some aspects of the environment are preserved and others are destroyed.
+Each timeline preserves different aspects, expressing a different “ought” which related work uses to explain the origins of goal directed behaviour [19].
 If the set of states is infinite then there are infinite possible worlds.
@@
 I do not mean like the words humans use to communicate.
-I mean a language of embodiment, where every statement in that language constrains rather than describes the system, like the aforementioned boundary conditions on slime mold [nakagaki2000] [^36] Using this frame cognition can be formalised in terms that align with both enactivist [thompson2007] and pancomputationalist [piccinini2021] perspectives.
-Hence, cognition within Stack Theory is also called Pancomputational Enactivism[bennett2024a].
+I mean a language of embodiment, where every statement in that language constrains rather than describes the system, like the aforementioned boundary conditions on slime mold [62] (Using this frame cognition can be formalised in terms that align with both enactivist [86] and pancomputationalist [73] perspectives.
+Hence, cognition within Stack Theory is also called Pancomputational Enactivism[16].)
 S6. Extended related work on delegated software architectures and Active Inference/FEP
@@
 Federated machine learning delegates control to some extent (e.g.
-shifting data ownership and parts of optimisation to edge devices), which can improve privacy and efficiency [nori2025federated].
-Constructivist theories of artificial general intelligence emphasise multilayered self-organisation at the level of software while acknowledging at least some of the constraints imposed by hardware [thorisson2012, goertzel2014].
-The Auto-catalytic Endogenous Reflective Architecture (AERA) is a goal-driven ampliative reasoning system [nivel2013] based on constructivist principles that relies on causal knowledge representation to achieve autonomy [thorisson2018cumlearncause], “growing” knowledge in a collective meta-model through cumulative continuous learning [thorisson2019cumlearning], and explicitly targeting adaptability [thorisson2014autonomousNatCom, thorisson2020, eberding2024, sheikhlar2024] [^37] See also http://www.openaera.org..
-ActPC-Chem [goertzel2024] is a recent theory behind the distributed Hyperon architecture [goertzel2023] that was influenced by theoretical and methodological insights from AERA.
+shifting data ownership and parts of optimisation to edge devices), which can improve privacy and efficiency [52].
+Constructivist theories of artificial general intelligence emphasise multilayered self-organisation at the level of software while acknowledging at least some of the constraints imposed by hardware [90, 42].
+The Auto-catalytic Endogenous Reflective Architecture (AERA) is a goal-driven ampliative reasoning system [64] based on constructivist principles that relies on causal knowledge representation to achieve autonomy [89], “growing” knowledge in a collective meta-model through cumulative continuous learning [88], and explicitly targeting adaptability [87, 91, 33, 80] (See also http://www.openaera.org.).
+ActPC-Chem [43] is a recent theory behind the distributed Hyperon architecture [41] that was influenced by theoretical and methodological insights from AERA.
 Active Inference/FEP in relation to MCL.
-Active Inference (and the Free Energy Principle, FEP) provides a normative account of perception and action in which an agent maintains a generative model and selects actions to minimise variational free energy (and, by extension, expected free energy) [Friston2010, friston2013, parr2024].
-The present paper addresses a different axis: it states a structural constraint on stacks of abstraction layers, namely that what can be achieved at layer i+1i{+}1 (utility $\epsilon(\gamma_{i+1})$) is bounded by the extension (weakness) of the realised policy at layer $i$ (Theorem 5.1).
+Active Inference (and the Free Energy Principle, FEP) provides a normative account of perception and action in which an agent maintains a generative model and selects actions to minimise variational free energy (and, by extension, expected free energy) [39, 40, 71].
+The present paper addresses a different axis.
+It states a structural constraint on stacks of abstraction layers, namely that what can be achieved at layer i+1i{+}1 (utility $\epsilon(\gamma_{i+1})$) is bounded by the extension (weakness) of the realised policy at layer $i$ (Theorem 5.1).
 In this sense, MCL is about where adaptation is implemented (and how much latitude lower layers retain), rather than about the particular inferential objective used within a layer.
-An Active Inference agent implemented on a predominantly fixed computational substrate can still be bottlenecked by static lower layers; conversely, a delegated multiscale substrate can in principle host a range of learning rules, including Active Inference.
+An Active Inference agent implemented on a predominantly fixed computational substrate can still be bottlenecked by static lower layers.
+Conversely, a delegated multiscale substrate can in principle host a range of learning rules, including Active Inference.
 Potential complementarity and open questions.
 Several lines of work suggest points of contact, without implying formal equivalence.
-ActPC-Chem frames delegation in explicitly Active Inference terms [goertzel2024, Friston2010, friston2013].
-“Scale-free” active inference proposes that similar dynamical principles repeat across scales [friston2023b, friston2024a], which is conceptually close to the MCA picture and to the idea that policy alignment must be maintained across layers to preserve collective identity.
-Delegated [ororbia2022, ororbia2023, ororbia2023b] or federated [friston2024b] Active Inference approaches may therefore be promising candidates for building efficient hybrid systems at scale.
+ActPC-Chem frames delegation in explicitly Active Inference terms [43, 39, 40].
+“Scale-free” active inference proposes that similar dynamical principles repeat across scales [36, 37], which is conceptually close to the MCA picture and to the idea that policy alignment must be maintained across layers to preserve collective identity.
+Delegated [68, 69, 67] or federated [38] Active Inference approaches may therefore be promising candidates for building efficient hybrid systems at scale.
 Separately, work on “mortal computation” seeks to approach thermodynamic limits (e.g.
-the Landauer bound) by reducing memory–processor separation and associated bottlenecks [ororbia2024mortalcomputationfoundationbiomimetic, aurell2012, zou2021], which resonates with the present paper’s emphasis on efficiency and with the intuition that pushing adaptation downward can reduce overhead.
+the Landauer bound) by reducing memory and processor separation and associated bottlenecks [66, 5, 97], which resonates with the present paper’s emphasis on efficiency and with the intuition that pushing adaptation downward can reduce overhead.
 However, the theories still differ in their primitives (e.g.
-tasks/weakness versus generative models/free energy), and it remains an open question whether weakness (w-maxing) can be recovered as, or justified by, a principle within the Active Inference apparatus.
-Establishing such a mapping is an important direction for future work.
+tasks and weakness versus generative models and free energy).
+Even so, a simple mapping is available when viability is treated as a hard constraint and preferences are uniform across viable futures, see the Mathematical Results subsection “Viability, lived history, and variational free energy”.
+A deeper identification in the general non-uniform case remains open.
 S7. Mission command example (Nordbat 2)
 A well documented example of the effectiveness of mission command in facilitating adaptation is Nordbat 2, a joint peacekeeping force deployed to Bosnia in the early 90s.
-Mission command allowed individual units to deviate from convention, adapt to local conditions and adopt a more ‘trigger-happy’ disposition that proved to be substantially more effective than the more rigid, rule based approach taken by the Dutch forces operating in the same region [ingesson2016].
-S8. From analogy to evidence: candidate empirical and simulation tests
+Mission command allowed individual units to deviate from convention, adapt to local conditions and adopt a more ‘trigger-happy’ disposition that proved to be substantially more effective than the more rigid, rule based approach taken by the Dutch forces operating in the same region [47, 48].
+Strategic-level intent and a small set of constraints are communicated downward, while tactical-level decision-makers fill in the remaining degrees of freedom on the ground.
+In this reading, the bridge between leadership and soldiers is not a detailed plan but a constraint system that keeps local improvisation aligned with higher-level purpose.
+The purpose of this note is not to treat Nordbat 2 as empirical evidence for Theorem 5.1, but to provide an intuition-building picture of what Stack Theory means by delegation [47, 48].
+Mapped to the formalism, the strategic layer specifies a parent task $\omega$ (intent, mandate, and coordination requirements), while the tactical layer faces a family of child tasks $\alpha\sqsubset\omega$ that vary with local information and contingencies.
+Mission command enlarges the set of admissible completions by avoiding over-specification of the realised policy.
+It preserves a larger extension $|\mathrm{Ext}(\pi_{i})|$ at the lower layer, enabling more distinct higher-layer behaviours to be expressed one layer up (Theorem 5.1).
+For hybrid and cyberphysical design, the analogy is that effective “control” can often be achieved by engineering boundary conditions (interfaces, reward signals, resource limits, safety constraints) rather than by prescribing micro-actions.
+Over-prescriptive constraints risk collapsing the intersection of viable component policies (loss of collective identity), whereas intent-plus-boundaries can preserve slack for local adaptation while maintaining coherence.
+S8. Candidate empirical and simulation tests
 The main text uses analogies (mission command, economic delegation, and distributed swarms) as intuition-building thought experiments rather than as evidence.
-To turn the delegation hypothesis into a research programme, one can state testable predictions and compare systems under controlled variation of delegation depth.
+It is primarily conceptual.
+This note lists candidate tests that target the core claim of Theorem 5.1.
+The goal is to compare systems under controlled variation of where adaptation is permitted in the stack.
 1. Agent-based simulations with tunable delegation.
-Construct an agent-based model of a collective where (i) individuals have local perception and local policy spaces, (ii) a collective objective is defined at a higher level, and (iii) a “delegation parameter” controls how much decision-making is centralised versus left to individuals (e.g.
-by constraining local policy spaces or routing decisions through a leader).
-Expose the collective to non-stationary environments (distribution shifts, component failures) and measure: sample efficiency (data required to regain acceptable performance), robustness (graceful degradation under failures), and identity preservation (whether the collective policy remains coherent rather than splintering into incompatible sub-collectives).
-Theorem 5.1 predicts that as delegation is pushed upward (lower layers become static or over-constrained), achievable utility at higher layers should become bounded by the reduced weakness of the lower-layer realised policies.
+Construct an agent-based model of a collective.
+Include a delegation parameter that controls whether decisions are taken locally or routed through higher-level coordination.
+Expose the collective to abrupt environmental shifts.
+Measure sample efficiency, energy use, robustness, and identity preservation.
+Test whether deeper delegation improves efficiency in the sense predicted by reduced weakness of lower-layer realised policies.
 2. Comparative benchmarking across stack depth.
 Compare learning systems that differ only in where adaptation is permitted.
-For example, keep a fixed high-level learning rule but vary whether lower layers are static (fixed hardware / fixed microcode) versus adaptive (e.g.
-reconfigurable hardware, adaptive firmware, or reconfigurable substrates).
-Measure sample and energy efficiency on tasks with distribution shift and compositional generalisation (e.g.
-ARC-style tasks).
-This isolates the “where in the stack” hypothesis from the choice of algorithm at the top.
+Hold the top-layer objective fixed and vary whether lower layers are static or reconfigurable.
+For example, compare fixed-substrate deployment against deployment that allows adaptation of firmware, routing, or morphology.
+Measure adaptation speed and compute the effective induced vocabulary size at each layer when possible.
+This isolates the stack-level delegation hypothesis from algorithm choice at the top.
 3. Hybrid boundary-condition sweeps.
-For hybrid agents such as simulated systems of human and AI agents, where the goal is to “align” the AI agents with goal goals, systematically vary the boundary conditions (stimulation regimes, feedback mappings, resource constraints) and quantify the trade-off between task performance and preservation of collective identity (non-empty intersections of viable policy sets, as in the organoid formalisation above).
-The design expectation is that weaker (less exclusionary) boundary conditions should preserve a larger viable policy space while still enabling the desired interface-level behaviour.
+For hybrid agents such as simulated mixtures of human and AI agents or engineered organoids, treat alignment as hybrid agent design.
+Sweep boundary conditions that prune low-level policy spaces.
+Quantify the trade-off between task performance at the interface and preservation of non-empty collective policy intersection.
+Look for phase transitions where the intersection collapses, which corresponds to loss of collective identity.
 4. Interaction with Active Inference/FEP.
-Implement Active Inference-style control at one or more layers, but run the same comparison as above: (i) active inference on a fixed substrate versus (ii) active inference on a substrate that permits deeper delegation (adaptation in lower layers).
-This helps disentangle improvements due to the inferential objective from improvements due to stack-level delegation.
-These designs are intended as starting points rather than definitive protocols; the unifying aim is to replace analogy with controlled variation of delegation, and to measure adaptability in the efficiency-of-adaptation sense used throughout this paper.
+Implement Active Inference style control at one or more layers.
+Run the same comparison in an architecture that differs only in whether lower layers can adapt.
+This disentangles gains due to the inferential objective from gains due to deeper delegation.
+S9. Extended applications and speculative directions
+This note collects material that is peripheral to the main theorem but may be useful for future-work framing.
+These applications are intended as design hypotheses rather than as evidence for Theorem 5.1.
+Organisational and economic systems.
+Free-market trade is a great example of how delegation can facilitate adaptability in a large collective.
+As Hayek pointed out [44], not all of the information is available to a central planner and so control is best delegated to the individual.
+However, the accumulation of resources within my formalism is implicit, rather than explicit.
+I have been concerned only with whether a system satisfies whatever arbitrary notion of correct I choose (i.e.
+does it achieve homeostatic and reproductive goals in whatever amount of time I care about?), rather than how it evolves over time as a dynamical system.
+In other words, I have concerned myself only with whether a system arrives on time at the destination I want, not the journey it took to get there.
+Future research might extend Stack Theory to formalise time-series, or the accumulation of various resources.
+For example, a stack could represent a sequence of tasks over time where higher layers are later in time.
+Earlier I mentioned that task correctness can denote more holistic notions of emergent features robust to lower-level variation.
+A time-series model could describe the conditions under which such invariants emerge.
+It could be used to explore how the delegation of control, distribution of work and the accrual of resources change over time.
+For example under some conditions a free market may degrade into a maladaptive, top-down system.
+Those parts deprived of control or resources may then not be able to construct a policy that aligns with the larger whole, and so splinter off.
+To preserve the integrity of the system, policies which might support the desired invariants could be modelled and evaluated.
+Policy could be located along axes of distribution and delegation, like an alternative political compass.
+In a simple zero-sum game like a race, one example of such a policy is adaptive rubber banding [61].
+A game mechanic (e.g.
+in Mario Kart) where success accrues risk as well as reward, keeping the game competitive.
+Regulation and over-constraint risks.
+The cancer analogy suggests a general caution.
+Guardrails that over-constrain a system may precipitate the failure mode they aim to prevent by collapsing the intersection of viable policies.
+For example, survey of 3,306 Germans found that restrictions promoting carbon-neutral lifestyles would trigger strong negative responses because they ‘restrict freedom’ [77].
+This may be most relevant to proposals for safeguarded AGI and related governance approaches [29, 28, 3, 42], given the formal results pertain specifically to a generalisation-optimal model of intelligence [18].
+These remarks are speculative and are intended as design hypotheses rather than claims about current regulatory regimes.

### v7

@@
-Are Biological Systems More Intelligent Than Artificial Intelligence?
+Hybrid Agency, Artificial Intelligence, Cyberphysical Systems
 Abstract
@@
 Systems are modelled as stacks of abstraction layers (Stack Theory) and compared by how effectively they delegate agentic control down their stacks.
-I illustrate this using computational, biological, military, governmental, and economic systems.
+I illustrate this using computational, biological, military, and economic systems.
 Contemporary AI typically relies on static, human-engineered stacks whose lower layers are fixed during deployment.
@@
 Biological systems are more intelligent because they delegate adaptation.
-Formally, I prove a theorem (The Law of the Stack) showing that adaptability at higher layers is bottlenecked by adaptability at lower layers.
-I further show that, under standard viability assumptions, maximising adaptability is equivalent to minimising variational free energy, implying that delegation is necessary for free-energy minimisation.
-Generalising bioelectric accounts of cancer as isolation from collective informational structures, I analyse cancer-like failure modes in non-biological systems when delegation is inadequate.
-This yields design principles for building robust systems via delegated control, and reframes hybrid agents (e.g.
+Formally, I prove three results.
+The first theorem is The Law of the Stack, which shows adaptability at higher layers is bottlenecked by adaptability at lower layers.
+Second, I show maximising adaptability is equivalent to minimising variational free energy under viability assumptions, proving delegation is necessary for free-energy minimisation.
+Finally, I generalise bioelectric accounts of cancer to show cancer-like failures can occur in non-biological systems when delegation is inadequate.
+These yield design principles for building robust systems via delegated control, re-framing hybrid agents (e.g.
 organoids or human–AI systems) as weak boundary-condition design problems in which constraints shape low-level policy spaces while preserving collective identity.
+keywords:
+Stack Theory, delegation, w-maxing, adaptability, hybrid agency
 1 Introduction
 Organoids are living systems whose behaviour is sculpted by engineered constraints.
-Human brain cells embedded in a computational interface have been trained to play Pong, achieving remarkable sample efficiency compared to a contemporary reinforcement learning agent [53].
-Similarly, Physarum slime mould is normally an indiscriminate expander.
-However it reliably computes shortest paths when placed inside a maze [62].
+Human brain cells embedded in a computational interface have been trained to play Pong, achieving remarkable sample efficiency compared to a contemporary reinforcement learning agent [khajehnejad2024].
+Similarly, Physarum slime mould is normally an indiscriminate expander, but it reliably computes shortest paths when placed inside a maze [nakagaki2000].
 These behaviours are neither programmed nor accidental.
 They arise because biological agency is coupled to carefully designed boundary conditions.
-Such systems are hybrid agents, jointly produced by bottom-up cellular dynamics and top-down environmental structure [82].
+Such systems are hybrid agents, jointly produced by bottom-up cellular dynamics and top-down environmental structure [sole2019].
 Biological systems self-organise across scales.
-Their parts and environment interact to produce spatiotemporal order [4, 78, 76, 9, 51, 39, 92].
-Many systems proactively learn, storing information about the past to anticipate future needs [54].
-To adapt in complex environments, a system must infer cause and effect [72, 60, 7].
-Throughout, I use ‘intelligence’ in the specific sense of the sample and energy efficiency of adaptation [16, 26, 11, 14] [^1] Mainstream AI typically treats intelligence as performance across tasks.
+Their parts and environment interact to produce spatiotemporal order [ashby1947, seeley2002, rosas2018, bell2007, kelso1997, Friston2010, tognoli2014].
+Many systems proactively learn, storing information about the past to anticipate future needs [kukushkin2024].
+To adapt in complex environments, a system must infer cause and effect [pearl2018, merker2005, barron2016].
+Throughout, I use ‘intelligence’ in the specific sense of the sample and energy efficiency of adaptation [bennett2024a, chollet2019, bennettmaruyama2022b, bennett2023c] [^1] Mainstream AI typically treats intelligence as performance across tasks.
 My conclusions concern efficiency of acquiring said performance.
@@
 See ESM Note S1..
-Two common explanations for biological efficiency are active inference under the Free Energy Principle (FEP) [39, 40, 71] and powerful inductive biases [26].
+Two common explanations for biological efficiency are active inference under the Free Energy Principle (FEP) [Friston2010, friston2013, parr2024] and powerful inductive biases [chollet2019].
 These address, respectively, how agents update beliefs and act, and what structure is pre-specified.
@@
 In dominant AI deployments, adaptation is concentrated at high software layers (model parameters and software-level state) while lower layers (instruction set, microarchitecture, hardware) are treated as fixed during deployment.
-[^2] Lower layers can carry mutable state (data, caches, firmware), and self-modifying substrates exist; the claim here is about the prevailing engineering pattern in mainstream ML systems.
-In contrast, biological multiscale competency architectures (MCA) delegate goal-directed control across scales, from cells to organs to organisms [34, 59].
+[^2] Lower layers can carry mutable state (data, caches, firmware), and self-modifying substrates exist.
+The claim only pertains to the prevailing engineering pattern in mainstream ML systems.
+In contrast, biological multiscale competency architectures (MCA) delegate goal-directed control across scales, from cells to organs to organisms [fields2020, levin2024].
 I model this difference with Stack Theory, formalising systems as stacks of abstraction layers.
@@
 This is not a claim about all possible future AI.
+Delegation is also not a binary distinction between “static AI” and “dynamic biology”.
+It is a continuum.
+Many deployed AI systems concentrate adaptation into an offline training phase, whereas biological systems distribute plasticity and regulation across morphology, metabolism and development.
 One aim of the paper is to identify design principles for building more adaptable hybrid and cyberphysical systems by shifting where adaptation occurs in the stack.
 This is because biological self-organisation delegates adaptation to lower levels of abstraction (for the generalisation of “levels of abstraction” beyond computing, see ESM Supplementary Note S2).
-Formally, I measure adaptability by the weakness of the constraints a system embodies [20, 16].
+Formally, I measure adaptability by the weakness of the constraints a system embodies [bennett2025d, bennett2024a].
 I then prove The Law of the Stack (Theorem 5.1), which is that utility (adaptability) achievable one layer up is bounded by the weakness of the realised policy one layer down.
@@
 $\widehat{\Pi}_{\alpha_{j}}\neq\emptyset$ for some $j$), then any continued function must proceed via a proper subset of parts whose restricted policy sets still admit a non-empty intersection.
-I refer to this formal situation as “cancer-like” purely by analogy.
+I refer to this formal situation as “cancer-like” purely by analogy (see Definition 6.1 and Proposition 6.1).
 Claim 1.4 (Hybrid creation as maximally weak boundary-condition design).
 Hybrid agents (e.g.
-organoids, soft robots, human–AI organisations) are best treated as boundary-condition design problems, which I formalise using Stack Theory.
+organoids, soft robots, human–AI organisations) are best treated as boundary-condition design problems, which I formalise using Stack Theory [^3] This paper’s use of “Stack Theory” is a formal model of layered control and delegation, distinct from and complementary to Bratton’s geopolitical use of “The Stack” [bratton2016]..
 One must engineer constraints that prune low-level policy spaces just enough to realise target collective behaviour while preserving a non-empty intersection of component policy sets (collective identity).
-Formalised this way, the optimal choice of boundary conditions are the weakest sufficient constraints to achieve the desired hybrid functionality.
+Formalised this way, the optimal choice of boundary conditions is the weakest set of sufficient constraints to achieve the desired hybrid functionality.
 The remainder of the paper develops the intuition (Section 2), states the formal framework and theorem, and applies the result to hybrid agency and failure modes.
@@
 How efficiently a system adapts depends on the interaction between these dimensions.
-In proposing Stack Theory [19], I extend this idea beyond computing.
+In proposing Stack Theory [bennett2025thesis], I extend this idea beyond computing.
 I illustrate abstraction, distribution, and delegation with limiting-case examples from computing, biology, and human organisations.
@@
 In Stack Theory, useful invariants are treated as higher-level task constraints even when micro-implementations vary.
-Even hardware itself can be framed as an abstraction layer over the local environment [19].
+Even hardware itself can be framed as an abstraction layer over the local environment [bennett2025thesis].
 Think of a stack as a directed graph of constraints.
 Higher-level behaviours are implemented by and limited by lower-level behaviours.
-In biology, cells enact organs, organs enact organisms, and organisms enact still more abstract behaviours such as communication and tool use [59, 16].
+In biology, cells enact organs, organs enact organisms, and organisms enact still more abstract behaviours such as communication and tool use [levin2024, bennett2024a].
 cellscellsorgansorgansorganismorganism
@@
 Distribution refers to how work is spread within a layer.
-In many biological systems, lower levels of abstraction are literally smaller parts [32], so “going down” is like zooming in.
+In many biological systems, lower levels of abstraction are literally smaller parts [delgado1997], so “going down” is like zooming in.
 Cells form organs and organisms form organisms.
@@
 Remove a GPU and it does not continue pursuing the system’s objectives, or any other for that matter.
-Biological collectives differ in that cells and tissues are autonomously goal-directed at their own scale [6].
+Biological collectives differ in that cells and tissues are autonomously goal-directed at their own scale [ball2023].
 Higher-level coherence arises when the viable policy sets of parts intersect, producing a collective policy (collective identity).
@@
 This is a form of top-down control.
-Bottom-up then becomes just whatever the individual humans do within those constraints, even though the system as a whole might have emerged entirely bottom-up from anarchy to begin with.
+Bottom-up then becomes just whatever the individual humans do within those constraints.
 Each part of a collective acts on its own, and its actions imply a ‘policy’.
@@
 If collective and individual policies align in service of homeostasis, then the system can persist.
-It has a collective identity or goal.
-That identity can constrain parts of the collective.
-For example, I might say the level of abstraction is cells, and then describe an organ as a collective policy of those cells.
-Likewise, I might look at the level of organs, and describe a human as a collective policy of those organs.
-Or I could describe the human as the collective policy of cells, and see the organs as a sort of middle management layer.
 Economies illustrate useful limit cases to understand delegation.
@@
 In the central planning limit case, the population does work but has no control.
-In the limit those people could adapt given local information unavailable to the central planner [44], but central planner’s edicts constrain them.
+In the limit those people could adapt given local information unavailable to the central planner [hayek1945use], but central planner’s edicts constrain them.
 In the opposite extreme there is no coordination.
 Real systems occupy a continuum.
-Of course, there are limits on what sort of physical forms can store or process information in service of a goal [81], and this may limit how much I can distribute work or delegate control.
+Of course, there are limits on what sort of physical forms can store or process information in service of a goal [sole2024], and this may limit how much I can distribute work or delegate control.
 Delegating control often coincides with shifting to finer granularity and greater distribution.
@@
 If I program in C it doesn’t matter whether I program one computer or fifty.
-Either way my control is at the same level of abstraction (figure from [19]).
+Either way my control is at the same level of abstraction (figure from [bennett2025thesis]).
 Figure 2: Work, resources and so on can be distributed across a system.
 Distribution is distinct from delegation of control.
-For example, both food stamps and universal basic income might distribute the same amount of resources to the same people within an economic system, but the former more tightly constrains the recipients in how they use those resources (figure from [19]).
+For example, both food stamps and universal basic income might distribute the same amount of resources to the same people within an economic system, but the former more tightly constrains the recipients in how they use those resources (figure from [bennett2025thesis]).
 2.4 Reasons to Suspect Delegation Facilitates Adaptation
@@
 Here I briefly motivate why delegation should matter.
-Because perceived complexity is defined by the abstraction layer through which a system interacts with its environment [17], measures of intelligence that ignore embodiment are incomplete; an “objective” measure must be enactive [86].
-First, adaptation closer to the substrate can exploit more degrees of freedom.
-Cutting out an abstraction layer can yield large efficiency gains (C vs Python) [65]; reconfigurable substrates such as FPGAs push adaptation further down the stack [63].
-Second, delegation can improve robustness via graceful degradation, avoiding single points of failure [55].
-For example, to control a drone swarm I could single out one drone to be the ‘leader’, and have it run a program that co-ordinates the other drones.
-However that would make the swarm vulnerable.
-If the leader drone were damaged, the swarm would de-cohere.
-In contrast if I can get the swarm to co-ordinate through entirely decentralised control, individual drones negotiating with those around them, then the swarm can remain coherent regardless of how many drones are disabled.
-This sort of decentralised swarm control can also be observed in nature, such as with flocks of starlings which fly in complex formations [25].
-In other words, the swarm ‘adapts’ to adverse conditions.
+Because perceived complexity is defined by the abstraction layer through which a system interacts with its environment [bennett2024b], measures of intelligence that ignore embodiment are incomplete.
+An “objective” measure must be enactive [thompson2007].
+First, adaptation closer to the substrate can exploit more degrees of freedom and specialisation.
+Programming at a lower level of abstraction can yield large efficiency gains (e.g.
+C vs Python) [oden2020].
+Reconfigurable substrates such as FPGAs can push adaptation further down the stack, enabling even more specialisation and efficiency [nguyen2020].
+Second, delegation can improve robustness via graceful degradation, avoiding single points of failure [kuniavsky2010].
+For example, to co-ordinate a drone swarm I could program one drone to plan the movements of the others.
+However, if that leader drone were damaged, the swarm would fail to co-ordinate.
+A more resilient approach is to program the swarm in a decentralised manner where each drone negotiates solutions with its neighbours, so the swarm has no single point of failure.
+Decentralised swarm control also emerges naturally, such as with flocks of starlings flying in complex formations [cavagna2010].
 Third, in human organisations delegation shortens feedback loops.
-In military doctrine, “mission command” delegates decision-making to units closest to the situation, granting latitude to deviate from detailed orders in service of higher-level intent [79].
+In military doctrine, “mission command” delegates decision-making to units closest to the situation, granting latitude to deviate from detailed orders in service of higher-level intent [Shamir01102010].
 This trades central control for responsiveness.
-See ESM Note S7 for a concrete historical illustration (Nordbat 2 [47, 48]).
+See ESM Note S7 for a concrete historical illustration (Nordbat 2 [ingesson2016, ingesson2017]).
 These are all intuition-building thought experiments rather than empirical proof, but they clarify the mechanism the theorem captures.
@@
 causeeffect
-To make accurate inferences a system must be able to tell the difference between what it has caused, and what it has merely observed [60, 72, 31].
+To make accurate inferences a system must be able to tell the difference between what it has caused, and what it has merely observed [merker2005, pearl2018, dawid2002].
 It is often difficult to determine cause.
@@
 not the causeeffectcause
-By actively intervening in the environment, a system can experiment to establish cause [^3] Assuming such experimentation is enough to mitigate ambiguity and eliminate unknowns, and that a heuristic such as weakness is applied to ensure the system converges on variable definitions which represent cause accurately [14]..
+By actively intervening in the environment, a system can experiment to establish cause [^4] Assuming such experimentation is enough to mitigate ambiguity and eliminate unknowns, and that a heuristic such as weakness is applied to ensure the system converges on variable definitions which represent cause accurately [bennett2023c]..
 Then, it can construct a causal graph.
@@
 One option would be to assume the nodes (variables, actions and other abstract objects and properties), and then learn the causal relations between them.
-This might work well in theory [74], but it requires foreknowledge of objects and properties between which causal relations can be established.
-The alternative is to assume the causal relations (the edges in the graph), and then learn the variables [14, 19, 18].
-This needs only attraction and repulsion from physical states [12, 19], otherwise known as valence.
+This might work well in theory [richens2024robust], but it requires foreknowledge of objects and properties between which causal relations can be established.
+The alternative is to assume the causal relations (the edges in the graph), and then learn the variables [bennett2023c, bennett2025thesis, bennett2025e].
+This needs only attraction and repulsion from physical states [bennettwelshciaunica2024, bennett2025thesis], otherwise known as valence.
 I take this latter approach.
@@
 For each input the computer responded with a particular output.
-The computer runs a ‘policy’ program which maps inputs to outputs [^4] I call such programs policies, but this should not be confused with the more specific definition policy in reinforcement learning [84]..
+The computer runs a ‘policy’ program which maps inputs to outputs [^5] I call such programs policies, but this should not be confused with the more specific definition policy in reinforcement learning [sutton2018]..
 That policy is the cause of the input-output history observed so far.
@@
 If I wish to know which policy actually caused the outputs, I need a way to whittle down the remaining possibilities.
-There are two ways to compare policies that look the same with respect to historical data, but differ with respect to possible future data [17].
+There are two ways to compare policies that look the same with respect to historical data, but differ with respect to possible future data [bennett2024b].
 One is to compare their form, and choose policies which are simpler in the sense of being more compressible.
-However, in an interactive setting simplicity is an unreliable indicator of plausibility [56, 17].
+However, in an interactive setting simplicity is an unreliable indicator of plausibility [leike2015, bennett2024b].
 The alternative is to compare policies by their function, in terms of how they behave given inputs for which I do not have historical data.
-Adaptive systems must satisfy functional constraints [83].
+Adaptive systems must satisfy functional constraints [sole2022].
 I can compare systems by these constraints.
@@
 Put another way, if two policies satisfy a goal, then the weaker policy satisfies it in more possible worlds [^6] To the extent that such possible worlds can be differentiated from one another within the subjective constraint of an abstraction layer..
-Previous research showed that if two policies are identical with respect to historical data, then the policy most likely to have generated historical data is the weaker [21, 19] (see ESM Supplementary Note S3 for formal details.) of the two [14, 20, 18].
-The policy most likely to have caused behaviour is the weakest [14], and in experiments a system which preferred weaker policies outperformed one that preferred the simpler policies by $110-500\%$ [15].
-This has been used to establish an upper bound on embodied ‘intelligence’ [16].
-I will use the term ‘w-maxing’ to mean a system optimises for weak policies, and ‘simp-maxing’ to mean a system optimises for simpler (minimum description length [75]) policies, to conform to the terminology used in previous work [22, 19, 21].
+Previous research showed that if two policies are identical with respect to historical data, then the policy most likely to have generated historical data is the weaker [bennett2024appendices, bennett2025thesis] (see ESM Supplementary Note S3 for formal details.) of the two [bennett2023c, bennett2025d, bennett2025e].
+The policy most likely to have caused behaviour is the weakest [bennett2023c], and in experiments a system which preferred weaker policies outperformed one that preferred the simpler policies by $110-500\%$ [bennett2023b].
+This has been used to establish an upper bound on embodied ‘intelligence’ [bennett2024a].
+I will use the term ‘w-maxing’ to mean a system optimises for weak policies, and ‘simp-maxing’ to mean a system optimises for simpler (minimum description length [rissanen1978]) policies, to conform to the terminology used in previous work [bennett2025b, bennett2025thesis, bennett2024appendices].
 Put another way, if two hypotheses appear equally good with respect to past data, the w-maxing system will use policy weakness as a heuristic to decide which policy to use, whereas a simp-maxing system will use simplicity.
@@
 Simplicity is a property of a policy’s description in a chosen encoding.
-Weakness is a property of the policy’s remaining completion space inside the embodied language $L_{\mathfrak{v}}$.
+The weakness of a policy is the degrees of freedom remaining to a system that embodies that policy.
 Two correct policies can therefore have very different weakness even when their descriptions are equally short, or when one is syntactically longer.
@@
 For example a program takes a form when it is written in a language, but when it is run on a computer it has a particular function or effect upon the world.
-Simp-maxing is a matter of form, while w-maxing is a matter of function [22].
+Simp-maxing is a matter of form, while w-maxing is a matter of function [bennett2025b].
 When we account for interpretation between form and function it becomes clear that software is a physical thing, like a tool.
@@
 This is what “weakness” measures.
-A tool that implies a weaker constraint can complete more tasks [10] than another tool.
-By preferring weaker policies, one avoids unnecessary commitments, meaning the weakest policies complete a ‘wider range of tasks’ than others (see ESM Note S3 for discussion of weakness and related connections (e.g.
-maximum entropy).).
+A tool that implies a weaker constraint can complete more tasks [bennettmaruyama2022a] than another tool.
+By preferring weaker policies, one avoids unnecessary commitments, meaning the weakest policies complete a ‘wider range of tasks’ than others (see ESM Note S3 for discussion of weakness and related concepts, such as maximum entropy).
 To complete a ‘wider range of tasks’ in the embodied sense, a tool must be both more sample and more energy efficient, and so the weakest policy is the most efficient in both senses.
@@
 ML systems do not typically optimise for either weaker constraints on function (w-maxing), or simplicity of form (simp-maxing).
-This may explain some of their relative inefficiency [13, 15].
+This may explain some of their relative inefficiency [bennett2022a, bennett2023b].
 Even if they were w-maxing, the capacity for w-maxing is severely limited by the abstraction layer.
-An abstraction layer can be a “causal isolator” [35] between a software “mind” at higher levels of abstraction, and its environment at lower levels.
+An abstraction layer can be a “causal isolator” [friston2023] between a software “mind” at higher levels of abstraction, and its environment at lower levels.
 Embodiment by another name.
 Different abstraction layers permit different upper bounds on policy weakness, in the same way that two bodies might differ in how able they are to achieve a given goal.
-Different embodiments give different upper bounds on ‘intelligence’ [16].
+Different embodiments give different upper bounds on ‘intelligence’ [bennett2024a].
 That biological self-organising systems adapt efficiently means they are able to embody weaker policies than today’s ML systems.
-This capacity hinges on the abstraction layer [16].
+This capacity hinges on the abstraction layer [bennett2024a].
 An abstraction layer is like a window through which only part of the world can be seen.
@@
 Hardware, not just software.
-In enactivist terms, an intelligent choice of abstraction layer is one that reduces the “big world” of all information to a “small world” of relevant information [95, 94, 93].
+In enactivist terms, an intelligent choice of abstraction layer is one that reduces the “big world” of all information to a “small world” of relevant information [vervaeke2012, vervaeke2013a, vervaeke2013b].
 In a typical computer, lower levels of abstraction tend to be static.
@@
 The surrounding narrative provides intuition.
-The ESM contains additional motivation, proofs, and related technical material [19, 15, 14, 16, 17, 12, 18].
+The ESM contains additional motivation, proofs, and related technical material [bennett2025thesis, bennett2023b, bennett2023c, bennett2024a, bennett2024b, bennettwelshciaunica2024, bennett2025e].
 Stack Theory describes an adaptive system in terms of the binary constraints or boundary conditions that define it.
 Embodiment is a choice of abstraction layer.
-This is a finite vocabulary of possible constraints (programs) representing the possible configurations of a body embedded in its environment [19].
+This is a finite vocabulary of possible constraints (programs) representing the possible configurations of a body embedded in its environment [bennett2025thesis].
 These programs are used to express “statements” constraining the body.
@@
 A $\mathfrak{v}$-task $\langle I,O\rangle$ can represent biological selection pressures by specifying which outputs are viable for which inputs.
-For example, an FEP-style agent can be treated as a task whose admissible futures minimise surprisal under an appropriate generative model [27, 39].
-For readers who wish to check the formal dependencies (see also [19, 21]), the key definitions used in the theorem are summarised below.
+For example, an FEP-style agent can be treated as a task whose admissible futures minimise surprisal under an appropriate generative model [ciaunica2023b, Friston2010].
+For readers who wish to check the formal dependencies (see also [bennett2025thesis, bennett2024appendices]), the key definitions used in the theorem are summarised below.
   1. 1.
@@
     Intuition. Utility increases when the correct-output set can be realised by many policies. This captures adaptability under the embodied language.
-I set out to compare computers with biological systems. Since both are stacks of abstraction layers, I define multilayer architecture (MLA) below to describe them in comparable terms.
+I set out to compare computers with biological systems.
+Both are stacks of abstraction layers.
+Hence I define a stack (meaning a multilayer architecture or MLA) below.
+I do this to describe biological and contemporary computer systems in comparable terms.
   • The stack of layers of increasing abstraction is a sequence of uninstantiated tasks $\langle\lambda^{0},\lambda^{1},\dots,\lambda^{n}\rangle$ such that $\lambda^{i+1}\sqsubset\lambda^{i}$.
@@
     That is, the next-layer vocabulary consists of the truth sets of completions compatible with $\pi$.
-  • The state of an MLA is a sequence of policies $\langle\pi^{0},\pi^{1},\dots,\pi^{n}\rangle$ and a sequence of vocabularies $\langle\mathfrak{v}^{0},\mathfrak{v}^{1},\dots,\mathfrak{v}^{n}\rangle$ such that for each $i<n$,
+  • The state of the stack (MLA) is a sequence of policies $\langle\pi^{0},\pi^{1},\dots,\pi^{n}\rangle$ and a sequence of vocabularies $\langle\mathfrak{v}^{0},\mathfrak{v}^{1},\dots,\mathfrak{v}^{n}\rangle$ such that for each $i<n$,
     $\mathfrak{v}^{i+1}=\mathfrak{f}(\mathfrak{v}^{i},\pi^{i})\quad\text{and}\quad\pi^{i}\in\Pi_{\lambda^{i}(\mathfrak{v}^{i})}.$
@@
 The state records the current policy and induced vocabulary at each layer.
-I say the MLA is over-constrained when there exists $i<n$ such that $\Pi_{\lambda^{i}(\mathfrak{v}^{i})}=\emptyset$, and multilayer-causal-learning (MCL) is when the MLA is not over-constrained.
+I say the stack (MLA) is over-constrained when there exists $i<n$ such that $\Pi_{\lambda^{i}(\mathfrak{v}^{i})}=\emptyset$, and multilayer-causal-learning (MCL) is when the stack is not over-constrained.
 Over-constraint means some layer cannot realise any policy compatible with the higher-level task.
@@
 The probability that the parts of a system embody a collective policy is higher if they embody weaker policies.
-This is because weaker policies are more likely to generalise between $\mathfrak{v}$-tasks [18].
+Classical cybernetics characterises regulation in terms of variety.
+Ashby’s law of requisite variety says that a regulator must possess at least as much variety as the disturbances it must absorb.
+Delegated control can be read as a constructive way to generate such requisite variety by distributing adaptive degrees of freedom across layers rather than concentrating them in a single controller [ashby1957, wiener1948cybernetics].
+This also connects to cybernetic accounts of hierarchical viability and self-regulation, such as Beer’s Viable System Model [beer1995brain].
+This is because weaker policies are more likely to generalise between $\mathfrak{v}$-tasks [bennett2025e].
 Consider the aforementioned $\alpha_{1}$ and $\alpha_{2}$.
@@
 The difference is a matter of scale.
-Just as w-maxing facilitates causal learning at the scale of an individual [14], in the context of an MCA it will facilitate learning at multiple scales, because the system as a whole is just another $\mathfrak{v}$-task.
+Just as w-maxing facilitates causal learning at the scale of an individual [bennett2023c], in the context of an MCA it will facilitate learning at multiple scales, because the system as a whole is just another $\mathfrak{v}$-task.
 Mechanistically, “negotiation” between levels is represented as constraint compatibility between policy sets.
@@
 The cells are an abstraction layer in which the collective policy of the organ is expressed, like software.
-However, software is just a state of hardware [16].
+However, software is just a state of hardware [bennett2024a].
 Put another way, ‘organ’ is a label I ascribe to the behaviour of some cells, and ‘human’ is a label I ascribe to the behaviour of some organs.
@@
 To construct layers I would start with a vocabulary $\mathfrak{v}^{i}$ at $i=0$, for example implied by a collective of cells or a lower level system.
-If this system exhibits goal directed behaviour, then it is given by the $\mathfrak{v}^{i}$-task $\lambda^{i}(\mathfrak{v}^{i})$ and a policy $\pi^{i}\in\Pi_{\lambda^{i}(\mathfrak{v}^{i})}$.
+If this system exhibits goal-directed behaviour, then it is given by the $\mathfrak{v}^{i}$-task $\lambda^{i}(\mathfrak{v}^{i})$ and a policy $\pi^{i}\in\Pi_{\lambda^{i}(\mathfrak{v}^{i})}$.
 $\pi^{i}$ implies behaviour.
@@
 $\mathfrak{f}$ is applied to $\mathrm{Ext}(\pi^{i})$ to obtain a vocabulary $\mathfrak{v}^{i+1}$, and then $\lambda^{i+1}$ is applied to $\mathfrak{v}^{i+1}$ to obtain a task $\lambda^{i+1}(\mathfrak{v^{i+1}})$.
-If $\Pi_{\lambda^{i+1}(\mathfrak{v^{i+1}})}\neq\emptyset$, then there exists a common purpose or “self” at this level of abstraction and I select a another policy $\pi^{i+1}\in\Pi_{\lambda^{i+1}(\mathfrak{v^{i+1}})}$.
+If $\Pi_{\lambda^{i+1}(\mathfrak{v^{i+1}})}\neq\emptyset$, then there exists a common purpose or “self” at this level of abstraction and I select another policy $\pi^{i+1}\in\Pi_{\lambda^{i+1}(\mathfrak{v^{i+1}})}$.
 In narrative terms, adaptation at higher levels depends on adaptation at lower levels.
-I now state The Law of the Stack and give a brief proof sketch.
-The ESM provides a full proof and additional lemmas.
+I now state The Law of the Stack and give a brief proof sketch (the ESM provides a full proof and additional lemmas).
+Formally the next theorem quantifies a delegation bottleneck, that utility at layer i+1i{+}1 is upper-bounded by the weakness (slack) of the realised policy at layer $i$.
 Theorem 5.1 (The Law of the Stack).
-Let $\langle\lambda_{0},\lambda_{1},\dots,\lambda_{n}\rangle$ be an MLA with abstractor $\mathfrak{f}$ and state $\langle(\pi_{0},\dots,\pi_{n}),(v_{0},\dots,v_{n})\rangle$ such that $v_{i+1}=\mathfrak{f}(v_{i},\pi_{i})$ and $\pi_{i}\in\Pi_{\lambda_{i}(v_{i})}$ for all $i<n$.
+Let $\langle\lambda_{0},\lambda_{1},\dots,\lambda_{n}\rangle$ be a stack (MLA) with abstractor $\mathfrak{f}$ and state $\langle(\pi_{0},\dots,\pi_{n}),(v_{0},\dots,v_{n})\rangle$ such that $v_{i+1}=\mathfrak{f}(v_{i},\pi_{i})$ and $\pi_{i}\in\Pi_{\lambda_{i}(v_{i})}$ for all $i<n$.
 Fix any $i<n$, and write
 $v_{i+1}=\mathfrak{f}(v_{i},\pi_{i}),\qquad\gamma_{i+1}=\lambda_{i+1}(v_{i+1}).$
+Assume $\Pi_{\gamma_{i+1}}\neq\emptyset$ so that $\epsilon(\gamma_{i+1})$ is well-defined.
 Then the achievable utility at layer i+1i\!+\!1 is upper-bounded by the weakness of the policy at layer $i$ in the following sense.
@@
 Proof sketch.
-Utility $\epsilon(\gamma_{i+1})$ is attained by some correct policy at layer i+1i\!+\!1, whose extension is a subset of $L_{v_{i+1}}$, so $\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\leq|L_{v_{i+1}}|$.
+Since $\Pi_{\gamma_{i+1}}\neq\emptyset$, utility $\epsilon(\gamma_{i+1})$ is attained by some correct policy at layer i+1i\!+\!1, whose extension is a subset of $L_{v_{i+1}}$, so $\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\leq|L_{v_{i+1}}|$.
 Because $v_{i+1}$ is finite, $|L_{v_{i+1}}|\leq 2^{|v_{i+1}|}$.
@@
 It says that degrees of freedom in higher layers is bounded by lower-layer slack.
-This aligns with the Free Energy Principle (FEP), which frames adaptive behaviour as variational free energy minimisation [39, 40, 71].
+This aligns with the Free Energy Principle (FEP), which frames adaptive behaviour as variational free energy minimisation [Friston2010, friston2013, parr2024].
 In the ESM (“Viability, lived history, and variational free energy”) I show that under a uniform viability prior, minimising free energy is equivalent to maximising the entropy (log-count) of viable future continuations.
@@
 Intuitively, if lower layers are over-constrained, the induced higher-layer language is small, so even a maximum-entropy posterior cannot keep many viable futures open.
+The next corollary makes the delegation dependence formal. Under the uniform viability prior, abstract-layer free energy cannot be reduced below a floor set by lower-layer slack.
 Corollary 5.1 (Delegation is necessary for lowering free energy).
@@
 Write $O_{\mu}\subseteq L_{v_{i+1}}$ for the set of viable future continuations at that horizon.
-For any policy $\pi_{i+1}$ at layer i+1i{+}1, let $\Omega_{\pi_{i+1}}=\mathrm{Ext}(\pi_{i+1})\cap O_{\mu}$ be the set of viable futures compatible with $\pi_{i+1}$, and let $q_{\pi_{i+1}}$ be the uniform distribution on $\Omega_{\pi_{i+1}}$.
+For any policy $\pi_{i+1}$ at layer i+1i{+}1 with $\Omega_{\pi_{i+1}}\neq\emptyset$, let $\Omega_{\pi_{i+1}}=\mathrm{Ext}(\pi_{i+1})\cap O_{\mu}$ be the set of viable futures compatible with $\pi_{i+1}$, and let $q_{\pi_{i+1}}$ be the uniform distribution on $\Omega_{\pi_{i+1}}$.
 Under the uniform viability prior assumed in the ESM, the corresponding (base-2) free energy satisfies
@@
 $F_{2}(q_{\pi_{i+1}})\;\geq\;\log_{2}|O_{\mu}|-|\mathrm{Ext}(\pi_{i})|.$
-In particular, free energy minimisation at layer i+1i{+}1 is bottlenecked by $|\mathrm{Ext}(\pi_{i})|$: reducing $F_{2}(q_{\pi_{i+1}})$ requires sufficient lower-layer slack, i.e.
+In particular, free energy minimisation at layer i+1i{+}1 is bottlenecked by $|\mathrm{Ext}(\pi_{i})|$, so reducing $F_{2}(q_{\pi_{i+1}})$ requires sufficient lower-layer slack, i.e.
 a weak enough realised policy at the lower layer.
@@
 The Law of the Stack proves that if lower layers are frozen or over-constrained, there is an entropy bottleneck, so abstract-layer free energy has a floor.
+Remark (accuracy and complexity).
+In general, variational free energy admits a decomposition into an accuracy term and a complexity term.
+The uniform-viability model used in Corollary 5.1 folds accuracy into hard constraints, namely the viability constraint $\omega\in\Omega$, and the history-consistency restriction $\pi\in\Pi_{\mathfrak{h}}$ within the confines of an embodied vocabulary $\mathfrak{v}$.
+Utility and viability constrain what counts as accurate behaviour, and delegation affects residual flexibility within that accuracy envelope.
 Intervening in the stack limits the extent to which control can be delegated, and thus limits capacity for w-maxing or MCL.
@@
 After all, an increase in utility may increase the computational cost of searching for correct policies.
-Furthermore, a spatially extended environment constrains the size of vocabularies at all levels [17], because bounded systems can contain only finite information [8].
+Furthermore, a spatially extended environment constrains the size of vocabularies at all levels [bennett2024b], because bounded systems can contain only finite information [bekenstein1981].
 I mean lower levels learn an inductive bias for higher levels, balancing utility with other factors in service of goals at every scale.
@@
 A state analogous to cancer exists for distributed systems with insufficient delegation in adverse conditions.
-Some have argued cancer is a loss of collective identity [57, 30].
+Some have argued cancer is a loss of collective identity [levin2021, davies2011].
 In a biological system, when cells are isolated from the informational structure of their collective, they lose their high level identity.
@@
 Formal characterisation of the cancer-analogue condition
+Formally, the next definition and proposition show that if boundary conditions over-constrain a distributed system so that no collective policy exists, then any realised behaviour must be splintered, because it can only satisfy a proper subset of parts.
+This is a formal analogue rather than a biological claim.
+It captures the failure mode where the collective-policy intersection disappears while local viability persists, so continued function is only possible via coherent subcoalitions.
+In Stack Theory terms, this is the zero-slack limit of the delegation bottleneck quantified by Theorem 5.1.
+Tightening lower-layer constraints can shrink the induced higher-layer feasible set all the way to emptiness.
 To make the above precise, let $\alpha_{1},\dots,\alpha_{m}$ be $v$-tasks representing parts of a distributed system, with policy sets $\Pi_{\alpha_{j}}\subseteq L_{v}$, and let the collective task be $\omega$ with collective-policy set $\Pi_{\omega}:=\bigcap_{j=1}^{m}\Pi_{\alpha_{j}}$ (collective identity at $v$ requires $\Pi_{\omega}\neq\emptyset$).
@@
 Hence every realised policy fails at least one part’s constraints, so $S(\pi)\neq\{1,\dots,m\}$ for all $\pi\in L_{v}$.
+If $S(\pi)\neq\emptyset$ then $\pi\in\bigcap_{j\in S(\pi)}\widehat{\Pi}_{\alpha_{j}}$, so $S(\pi)$ is a splinter.
 If $\widehat{\Pi}_{\alpha_{j}}\neq\emptyset$ then $\{j\}$ is a splinter since $\bigcap_{k\in\{j\}}\widehat{\Pi}_{\alpha_{k}}=\widehat{\Pi}_{\alpha_{j}}\neq\emptyset$.
 $\square$
+Remark (integration and $\Phi_{\mathrm{IIT}}$).
+Proposition 6.1 characterises collective identity as the existence of a non-empty intersection of feasible policy-sets.
+This is a structural notion of integration.
+In the Stack Theoretic depiction of consciousness, a loss of collective identity would necessarily imply loss of selves (e.g.
+reafference) and the disintegration of phenomenal content [bennett2025thesis, bennettwelshciaunica2024, bennett2026a].
+Beyond Stack Theory, Integrated Information Theory formalises integration via the quantity $\Phi_{\mathrm{IIT}}$, which decreases when a system decomposes into weakly coupled parts [tononi2004, oizumi2014].
+Under that lens, splintering can be read as a loss of integration, rather than merely an over-constraint at one layer.
+Remark (temporal splintering and negotiation).
+The set-intersection definition in Proposition 6.1 is static.
+Hybrid agents often face transient conflicts where $\widehat{\Pi}_{\omega}(t)=\emptyset$ for short intervals.
+A natural dynamic refinement is to treat each part’s feasible set as time-indexed and to model resolution as a higher-level regulatory process that reallocates variety across layers to restore a non-empty viable intersection.
+This is closely aligned with cybernetic treatments of adaptive organisation, including Beer’s Viable System Model and Ashby’s requisite variety principle [beer1995brain, ashby1957].
+Future work may further extend this framing of cancer using the time semantics module associated with Stack Theory [bennett2026a].
 As task difficulty increases (or top-down control tightens), the set of correct collective policies shrinks.
@@
 Part of the collective stops participating in the collective policy and reverts to a lower-level policy set with a much larger extension (more autonomy).
-This is a form of “selective forgetting” [14] applied to constraints rather than data.
+This is a form of “selective forgetting” [bennett2023c] applied to constraints rather than data.
 This is analogous to cancer.
 $\alpha$ is isolated and can now pursue goals which do not align with those of the collective.
-This suggests systems should be minimally constrained so as to ensure graceful degradation or ‘sloppy fitness’ [2], maintaining core functionality in as wide a variety of unanticipated circumstances as possible.
+This suggests systems should be minimally constrained so as to ensure graceful degradation or ‘sloppy fitness’ [agosta2010], maintaining core functionality in as wide a variety of unanticipated circumstances as possible.
 The more difficult the environment or the more top-down the system is controlled, the more tightly constrained behaviour becomes until the system becomes over-constrained and splinters.
@@
 7 Hybrid Creation as Weak Boundary-Condition Design
-Consider the aforementioned example of the organoid adapted to play the game Pong with notable sample efficiency relative to conventional reinforcement-learning baselines [53].
+Consider the aforementioned example of the organoid adapted to play the game Pong with notable sample efficiency relative to conventional reinforcement-learning baselines [khajehnejad2024].
 Hybrid systems like organoids are not fixed interpreters of software-level policy, but collectives whose lower-level parts retain bottom-up agency.
@@
 Biological systems often adapt efficiently because they delegate adaptive control across scales.
-Using Stack Theory [19], I formalised delegation as the depth to which adaptation is permitted in a stack of abstraction layers and proved The Law of the Stack (Theorem 5.1).
+Using Stack Theory [bennett2025thesis], I formalised delegation as the depth to which adaptation is permitted in a stack of abstraction layers and proved The Law of the Stack (Theorem 5.1).
 Utility one layer up is bounded by the weakness of the realised policy one layer down.
@@
 This constraint is compatible with active inference.
-Under a simple viability prior, minimising variational free energy is equivalent to maximising entropy over viable futures (w-maxing; ESM “Viability, lived history, and variational free energy”).
-Corollary 5.1 shows that free energy minimisation inherits the same bottleneck: abstract-layer free energy cannot be reduced without sufficient lower-layer slack.
+Under a simple viability prior, minimising variational free energy is equivalent to maximising entropy over viable futures (w-maxing; see ESM “Viability, lived history, and variational free energy”).
+Corollary 5.1 shows that free energy minimisation inherits the same bottleneck, which is that abstract-layer free energy cannot be reduced without sufficient lower-layer slack.
 The implication for AI is not that computation cannot be intelligent, but that contemporary AI stacks tend to concentrate adaptation at high software layers while freezing lower layers, thereby throttling achievable weakness and adaptation efficiency.
-For hybrid agents (organoids, soft robots, and human–AI organisations) the design problem becomes boundary-condition engineering: constrain parts just enough to elicit target collective behaviour while preserving a non-empty intersection of viable policies (collective identity).
+For hybrid agents (organoids, soft robots, and human–AI organisations) the design problem is to impose boundary conditions, constraining parts just enough to elicit target collective behaviour while preserving a non-empty intersection of viable policies (collective identity).
 When that intersection collapses, the system can fragment into competing subagents (Section 6).
@@
 hardware-in-the-loop learning, morphologically adaptive substrates, and biologically integrated controllers) and measuring resulting degrees of freedom, free energy, and robustness.
-More broadly, this paper is a stepping stone toward a unified account of delegation, weakness/max-entropy heuristics, and multiscale agency in the larger Stack Theory programme [19].
-A practical design implication is that, when adaptability is desired, constraints should be weakened as far as possible without breaking the system’s intended function.
-Constraints that are not strictly necessary relative to the design goal should be removed (What counts as “strictly necessary” is relative to design goals and value judgements.
-See ESM Supplementary Note S4.).
-Humans often prefer ‘additive’ solutions [1], meaning if a problem can be solved by adding a rule or subtracting a rule, we tend to add a rule.
-However The Law of the Stack suggests a preference for subtractive solutions might be beneficial in many circumstances.
+More broadly, this paper is a stepping stone toward a unified account of delegation, weakness/max-entropy heuristics, and multiscale agency in the larger Stack Theory programme [bennett2025thesis].
 To speculate about future research beyond experimental verification, Stack Theory is a lens that can be applied generally.
-For example, a number of software-first approaches to AI already approximate aspects of this delegation picture, including federated learning [52], constructivist AGI programmes [90, 42], and active inference [39, 40].
+For example, a number of software-first approaches to AI already approximate aspects of this delegation picture, including federated learning [nori2025federated], constructivist AGI programmes [thorisson2012, goertzel2014], and active inference [Friston2010, friston2013].
 An expanded discussion (including AERA, Hyperon/ActPC-Chem, and links to scale-free active inference) is provided in the ESM (Supplementary Note S6).
-Stack Theory might also inform the design of ‘soft robots’ [58].
+Stack Theory might also inform the design of ‘soft robots’ [man2019].
 Soft robots are a proposed class of robots made of ‘soft’ materials that adapt.
 I have formalised adaptive materials as adaptation lower in the stack, which can be used as a formal foundation for soft robotics.
-Self-assembling nanoparticle networks [24] are a step in this direction.
-They can be used to form reconfigurable logic gates [70], facilitating adaptation at very low levels of abstraction.
-Taken to its logical conclusion, future AI may more closely resemble artificial life [85, 45, 46].
+Self-assembling nanoparticle networks [borghi2024_brainlikehw] are a step in this direction.
+They can be used to form reconfigurable logic gates [paroli2023_receptron], facilitating adaptation at very low levels of abstraction.
+Taken to its logical conclusion, future AI may more closely resemble artificial life [suzuki2006, ikegami2008, ikegami2007simulating].
 Finally, Stack Theory may be extended to economic and organisational systems.
-For example, free-market trade could be seen as a classic argument for delegation.
-As Hayek pointed out [44], a great deal of relevant information is spread among a population, unavailable to a central planner.
+For example, free-market trade could also be seen as a classic argument for delegation.
+As Hayek pointed out [hayek1945use], a great deal of relevant information is spread among a population, unavailable to a central planner.
 This motivates delegation of control, to shorten feedback loops.
-However, the present formalism treats resources implicitly via task correctness, so connecting these arguments to economic dynamics would require extending Stack Theory to explicit time-series [23] and resource accumulation.
+However, the present formalism treats resources implicitly via task correctness, so connecting these arguments to economic dynamics would require extending Stack Theory to explicit time-series [bennett2026a] and resource accumulation.
 Additional speculative extensions and applications (including organisational and economic systems, AI safety regulation, and temporal extensions of stacks) are collected in ESM Note S9.
@@
 The Law of the Stack shows how adaptability at higher levels of abstraction is constrained by adaptability at lower levels, and delegation is structurally necessary to minimise free energy.
-This lays a foundation for a broad range of future theoretical and empirical research, showing how systems of any kind might be made more robust and adaptable.
+A practical design implication is that, when adaptability is desired, constraints should be weakened as far as possible without breaking the system’s intended function.
+Constraints that are not strictly necessary relative to the design goal should be removed, though what counts as “strictly necessary” is relative to design goals and value judgements (See ESM Supplementary Note S4.).
+Humans often prefer ‘additive’ solutions [adams2021], meaning if a problem can be solved by adding a rule or subtracting a rule, we tend to add a rule.
+However The Law of the Stack suggests a preference for subtractive solutions might be beneficial in many circumstances.
+The Law of the Stack might also be viewed as a formal justification of the classical principle of subsidiarity, often traced back to Aristotle, namely that higher levels of organization should coordinate and enable lower levels rather than replace functions they can adequately perform [aristotle_politics].
+It points out a principle common to many systems, formalised in terms of Stack Theory and structurally tied to the FEP.
+Hence this paper lays a foundation for a broad range of future theoretical and empirical research, showing how systems of any kind might be made more robust and adaptable.
 Acknowledgements
-First, I’d like to thank the anonymous peer reviewers for their detailed and helpful feedback.
-Second, I’d like to thank Anna Ciaunica, Peter Watts, Sean Welsh, Ricard Solé, Karl Friston, Antonio Damasio, Kingson Man, Paolo Milani, Ben Goertzel, Kristinn R.
+I thank the anonymous peer reviewers for their detailed and helpful feedback.
+It made a substantial difference.
+I also thank Anna Ciaunica, Peter Watts, Sean Welsh, Ricard Solé, Karl Friston, Antonio Damasio, Kingson Man, Paolo Milani, Ben Goertzel, Kristinn R.
 Thórisson, Elija Perrier and Noel Hinton for their support and feedback on this paper.
-S1. Working definition of intelligence
-There are many definitions of intelligence, and the merits of each are beyond the scope of this paper.
-Mine is based on [19], which frames intelligence as the ability to complete a wide range of tasks, which is sample and energy efficiency in adaptation.
-It is similar to Wang’s definition of intelligence as adaptation with limited resources [96], and Hutter’s earlier definition of intelligence as ‘the ability to satisfy goals in a wide range of environments’.
-Mainstream AI often operationalizes ‘intelligence’ as general-purpose task performance.
-By that standard, one may argue that current large language models are ‘more intelligent’ than many biological systems.
-However that definition is fundamentally flawed, as it fails to express the comparative efficiency with which humans acquire general purpose competence [16, 96, 26].
-In theory even a lookup table or rote-learning response can possess general purpose competence, given enough resources.
-This paper instead treats intelligence as the cost of adaptation (sample and energy efficiency), and the results should be read accordingly.
-S2. Levels of abstraction beyond computing
-The notion of levels of abstraction is more typically used in pure computer science settings, but can be applied to any system [19].
-A higher level of abstraction is the behaviour of the layers below.
-For example, an organ like a heart is a behaviour of cells, human is a behaviour of some organs, and language is a behaviour of humans.
-S3. Weakness and w-maxing (extended intuition)
-Weakness is formally defined in the appendix.
-Related work proved w-maxing to be optimal regardless of the abstraction layer [18, 15, 16, 19].
-In narrative terms, assume two policies $x$ and $y$ both constrain the system to examples of “correct” behaviour given past inputs, but differ in weakness.
-If $x$ is weaker than $y$, that means it commits to less, but achieves the same end.
-For example, if I assume the lawyer was paid at least $\mathdollar 5$, I am more weakly constrained than if I assume that lawyer was paid at least $\mathdollar 6$.
-This example also helps separate weakness from simplicity.
-The difference between $\mathdollar 5$ and $\mathdollar 6$ is not primarily about how hard the rule is to write down.
-It is about how many further commitments remain compatible with the rule, given the agent’s available distinctions.
-Weakness is defined relative to an abstraction layer.
-The layer fixes what counts as a meaningful refinement because it fixes what the system can jointly discriminate and jointly realise.
-If the embodiment changes, then the induced vocabulary changes, and so does the completion structure that weakness counts.
-Many familiar heuristics can be seen as special cases of the same idea once this completion structure is made explicit.
-If one treats all completions that remain possible at a layer as equally plausible, then choosing a correct policy that maximises $|\mathrm{Ext}(\pi)|$ is equivalent to maximising entropy over viable futures inside that layer.
-This parallels least commitment and version space intuitions, but it is stated directly in terms of the agent’s embodied language rather than in terms of an encoding chosen by an observer.
-That is what lets weakness be used across layers in Stack Theory, where a policy at one layer induces the next layer vocabulary via the abstractor map.
-In the language of agency, weakness can be read as a quantitative proxy for autonomy at a layer.
-A weaker policy leaves more degrees of freedom (more “ways of being right”) while still satisfying the same observed constraints.
-In multiscale and hybrid systems this autonomy matters because it increases the chance that component policies remain compatible with a higher-level collective policy, preserving collective identity.
-As an epistemological razor, this means “explanations should be no more specific than necessary” or alternatively “rules should constrain no more than necessary” (see Bennett’s Razor in [15]).
-This is a bit like the principle of maximum entropy [50, 49], difference being that w-maxing also accounts for the abstraction layer, which is important because abstraction can render complexity subjective [17].
-For a rigorous definition of learning and efficiency, see definitions in the online repository for the thesis [21].
-It includes proofs showing w-maxing is both necessary and sufficient to maximise the probability of ‘generalising’ from one task to another, and thus adaptability.
-S4. Normativity and “strict necessity” in design
-As Hume argued, a description of what ought to be cannot be derived from only descriptions of what is.
-“Strictly necessary” is a value judgement, or an implicit choice of what ought to be.
-In the thesis [19] on which this is based this ought from which other descriptions of ought follow is derived from the state transitions of the environment.
-These transitions preserve some aspects of that environment, while destroying others.
-Here however I am more concerned with the design of systems which serve a human purpose.
-Strictly necessary is a matter of desired functionality.
-S5. Conceptual motivation for the formal setting
-Stack Theory frames the description of an objective reality in terms of choices of subjective lens (an abstraction layer) [19].
-The full justification for these definitions is out of scope and addressed at length elsewhere [16, 19, 18].
-However I will provide a brief overview.
-To make claims that hold regardless of the abstraction layer, we must rely only on what holds for every abstraction layer.
-From our subjective perspective we can’t know what is at the bottom of our stack (it may be infinite).
-I call this bottom layer the environment.
-Whatever the environment is, there must be change or difference for there to be anything at higher levels of abstraction.
-Hence there must be a set of states where each state is one point of difference (Nothing is assumed about what these states might ontologically be or epistemologically contain.).
-Because these states must be mutually exclusive, this frames time as difference.
-Each possible sequence of states is a timeline or possible world.
-Truth is defined ontologically, relative to states.
-A set of states is effectively a set of truth conditions, so a possible fact is called a declarative program, and a program is true given a state $\phi$ if it contains that state.
-Every aspect of the environment is just a set of programs.
-Everything that exists is a set of facts, and each timeline preserves different sets of facts.
-A timeline is like an opinion.
-As time progresses some aspects of the environment are preserved and others are destroyed.
-Each timeline preserves different aspects, expressing a different “ought” which related work uses to explain the origins of goal directed behaviour [19].
-If the set of states is infinite then there are infinite possible worlds.
-Assume I have a robotic system.
-I get an abstraction layer for the system by choosing a finite aspect $\mathfrak{v}$ of programs called a vocabulary.
-Everything the abstraction layer can express is a subset of its vocabulary.
-For example if a system has an arm it can raise or lower, then there are programs in $\mathfrak{v}$ which correspond to actuation of the arm in some form, though it need not be neat symbolic representations like ’raise arm’ or ’lower arm’.
-Since states are mutually exclusive, $\mathfrak{v}$ gives us an embodied formal language.
-I do not mean like the words humans use to communicate.
-I mean a language of embodiment, where every statement in that language constrains rather than describes the system, like the aforementioned boundary conditions on slime mold [62] (Using this frame cognition can be formalised in terms that align with both enactivist [86] and pancomputationalist [73] perspectives.
-Hence, cognition within Stack Theory is also called Pancomputational Enactivism[16].)
-S6. Extended related work on delegated software architectures and Active Inference/FEP
-There are examples of systems that take something like the present approach, albeit within the confines of software abstraction layers.
-Federated machine learning delegates control to some extent (e.g.
-shifting data ownership and parts of optimisation to edge devices), which can improve privacy and efficiency [52].
-Constructivist theories of artificial general intelligence emphasise multilayered self-organisation at the level of software while acknowledging at least some of the constraints imposed by hardware [90, 42].
-The Auto-catalytic Endogenous Reflective Architecture (AERA) is a goal-driven ampliative reasoning system [64] based on constructivist principles that relies on causal knowledge representation to achieve autonomy [89], “growing” knowledge in a collective meta-model through cumulative continuous learning [88], and explicitly targeting adaptability [87, 91, 33, 80] (See also http://www.openaera.org.).
-ActPC-Chem [43] is a recent theory behind the distributed Hyperon architecture [41] that was influenced by theoretical and methodological insights from AERA.
-Active Inference/FEP in relation to MCL.
-Active Inference (and the Free Energy Principle, FEP) provides a normative account of perception and action in which an agent maintains a generative model and selects actions to minimise variational free energy (and, by extension, expected free energy) [39, 40, 71].
-The present paper addresses a different axis.
-It states a structural constraint on stacks of abstraction layers, namely that what can be achieved at layer i+1i{+}1 (utility $\epsilon(\gamma_{i+1})$) is bounded by the extension (weakness) of the realised policy at layer $i$ (Theorem 5.1).
-In this sense, MCL is about where adaptation is implemented (and how much latitude lower layers retain), rather than about the particular inferential objective used within a layer.
-An Active Inference agent implemented on a predominantly fixed computational substrate can still be bottlenecked by static lower layers.
-Conversely, a delegated multiscale substrate can in principle host a range of learning rules, including Active Inference.
-Potential complementarity and open questions.
-Several lines of work suggest points of contact, without implying formal equivalence.
-ActPC-Chem frames delegation in explicitly Active Inference terms [43, 39, 40].
-“Scale-free” active inference proposes that similar dynamical principles repeat across scales [36, 37], which is conceptually close to the MCA picture and to the idea that policy alignment must be maintained across layers to preserve collective identity.
-Delegated [68, 69, 67] or federated [38] Active Inference approaches may therefore be promising candidates for building efficient hybrid systems at scale.
-Separately, work on “mortal computation” seeks to approach thermodynamic limits (e.g.
-the Landauer bound) by reducing memory and processor separation and associated bottlenecks [66, 5, 97], which resonates with the present paper’s emphasis on efficiency and with the intuition that pushing adaptation downward can reduce overhead.
-However, the theories still differ in their primitives (e.g.
-tasks and weakness versus generative models and free energy).
-Even so, a simple mapping is available when viability is treated as a hard constraint and preferences are uniform across viable futures, see the Mathematical Results subsection “Viability, lived history, and variational free energy”.
-A deeper identification in the general non-uniform case remains open.
-S7. Mission command example (Nordbat 2)
-A well documented example of the effectiveness of mission command in facilitating adaptation is Nordbat 2, a joint peacekeeping force deployed to Bosnia in the early 90s.
-Mission command allowed individual units to deviate from convention, adapt to local conditions and adopt a more ‘trigger-happy’ disposition that proved to be substantially more effective than the more rigid, rule based approach taken by the Dutch forces operating in the same region [47, 48].
-Strategic-level intent and a small set of constraints are communicated downward, while tactical-level decision-makers fill in the remaining degrees of freedom on the ground.
-In this reading, the bridge between leadership and soldiers is not a detailed plan but a constraint system that keeps local improvisation aligned with higher-level purpose.
-The purpose of this note is not to treat Nordbat 2 as empirical evidence for Theorem 5.1, but to provide an intuition-building picture of what Stack Theory means by delegation [47, 48].
-Mapped to the formalism, the strategic layer specifies a parent task $\omega$ (intent, mandate, and coordination requirements), while the tactical layer faces a family of child tasks $\alpha\sqsubset\omega$ that vary with local information and contingencies.
-Mission command enlarges the set of admissible completions by avoiding over-specification of the realised policy.
-It preserves a larger extension $|\mathrm{Ext}(\pi_{i})|$ at the lower layer, enabling more distinct higher-layer behaviours to be expressed one layer up (Theorem 5.1).
-For hybrid and cyberphysical design, the analogy is that effective “control” can often be achieved by engineering boundary conditions (interfaces, reward signals, resource limits, safety constraints) rather than by prescribing micro-actions.
-Over-prescriptive constraints risk collapsing the intersection of viable component policies (loss of collective identity), whereas intent-plus-boundaries can preserve slack for local adaptation while maintaining coherence.
-S8. Candidate empirical and simulation tests
-The main text uses analogies (mission command, economic delegation, and distributed swarms) as intuition-building thought experiments rather than as evidence.
-It is primarily conceptual.
-This note lists candidate tests that target the core claim of Theorem 5.1.
-The goal is to compare systems under controlled variation of where adaptation is permitted in the stack.
-1. Agent-based simulations with tunable delegation.
-Construct an agent-based model of a collective.
-Include a delegation parameter that controls whether decisions are taken locally or routed through higher-level coordination.
-Expose the collective to abrupt environmental shifts.
-Measure sample efficiency, energy use, robustness, and identity preservation.
-Test whether deeper delegation improves efficiency in the sense predicted by reduced weakness of lower-layer realised policies.
-2. Comparative benchmarking across stack depth.
-Compare learning systems that differ only in where adaptation is permitted.
-Hold the top-layer objective fixed and vary whether lower layers are static or reconfigurable.
-For example, compare fixed-substrate deployment against deployment that allows adaptation of firmware, routing, or morphology.
-Measure adaptation speed and compute the effective induced vocabulary size at each layer when possible.
-This isolates the stack-level delegation hypothesis from algorithm choice at the top.
-3. Hybrid boundary-condition sweeps.
-For hybrid agents such as simulated mixtures of human and AI agents or engineered organoids, treat alignment as hybrid agent design.
-Sweep boundary conditions that prune low-level policy spaces.
-Quantify the trade-off between task performance at the interface and preservation of non-empty collective policy intersection.
-Look for phase transitions where the intersection collapses, which corresponds to loss of collective identity.
-4. Interaction with Active Inference/FEP.
-Implement Active Inference style control at one or more layers.
-Run the same comparison in an architecture that differs only in whether lower layers can adapt.
-This disentangles gains due to the inferential objective from gains due to deeper delegation.
-S9. Extended applications and speculative directions
-This note collects material that is peripheral to the main theorem but may be useful for future-work framing.
-These applications are intended as design hypotheses rather than as evidence for Theorem 5.1.
-Organisational and economic systems.
-Free-market trade is a great example of how delegation can facilitate adaptability in a large collective.
-As Hayek pointed out [44], not all of the information is available to a central planner and so control is best delegated to the individual.
-However, the accumulation of resources within my formalism is implicit, rather than explicit.
-I have been concerned only with whether a system satisfies whatever arbitrary notion of correct I choose (i.e.
-does it achieve homeostatic and reproductive goals in whatever amount of time I care about?), rather than how it evolves over time as a dynamical system.
-In other words, I have concerned myself only with whether a system arrives on time at the destination I want, not the journey it took to get there.
-Future research might extend Stack Theory to formalise time-series, or the accumulation of various resources.
-For example, a stack could represent a sequence of tasks over time where higher layers are later in time.
-Earlier I mentioned that task correctness can denote more holistic notions of emergent features robust to lower-level variation.
-A time-series model could describe the conditions under which such invariants emerge.
-It could be used to explore how the delegation of control, distribution of work and the accrual of resources change over time.
-For example under some conditions a free market may degrade into a maladaptive, top-down system.
-Those parts deprived of control or resources may then not be able to construct a policy that aligns with the larger whole, and so splinter off.
-To preserve the integrity of the system, policies which might support the desired invariants could be modelled and evaluated.
-Policy could be located along axes of distribution and delegation, like an alternative political compass.
-In a simple zero-sum game like a race, one example of such a policy is adaptive rubber banding [61].
-A game mechanic (e.g.
-in Mario Kart) where success accrues risk as well as reward, keeping the game competitive.
-Regulation and over-constraint risks.
-The cancer analogy suggests a general caution.
-Guardrails that over-constrain a system may precipitate the failure mode they aim to prevent by collapsing the intersection of viable policies.
-For example, survey of 3,306 Germans found that restrictions promoting carbon-neutral lifestyles would trigger strong negative responses because they ‘restrict freedom’ [77].
-This may be most relevant to proposals for safeguarded AGI and related governance approaches [29, 28, 3, 42], given the formal results pertain specifically to a generalisation-optimal model of intelligence [18].
-These remarks are speculative and are intended as design hypotheses rather than claims about current regulatory regimes.